FLOSSII: Logbook

FLOSSII: Chronological Logbook Entries


Return to Logbook Top Page
Entry Date Title Site Author #Graphics
229 Tue 22-Apr-2003pyg IC now okayoncley
228 Tue 22-Apr-2003pyg intercomparisonall stnsknudson
227 Thu 10-Apr-2003TRH intercomparisonall stnsknudson
226 Mon 07-Apr-2003TearDown Sensor IDsall stnsmilitzer
225 Tue 01-Apr-2003daily status March 31all stnsbeierle
224 Fri 28-Mar-2003Tcam studysagedelany
223 Fri 28-Mar-2003tower site visittowerdelany
222 Fri 28-Mar-2003power diedtoweroncley
221 Fri 28-Mar-2003bad T/RH fan at 20mtoweroncley
220 Thu 27-Mar-2003daily status March 27all stnsbeierle
219 Wed 26-Mar-2003daily status March 26all stnsbeierle
218 Tue 25-Mar-2003daily status March 25all stnsbeierle
217 Mon 24-Mar-2003daily status March 24all stnsbeierle
216 Sat 22-Mar-2003Set CO2 lag to 1 sectowermaclean
215 Fri 21-Mar-2003daily status March 21all stnsbeierle
214 Fri 21-Mar-2003daily status March 20all stnsbeierle
213 Fri 21-Mar-2003status March 17all stnsbeierle
212 Mon 17-Mar-2003status March 17all stnsbeierle
211 Fri 14-Mar-2003status March 14all stnsbeierle
210 Thu 13-Mar-2003status March 13all stnsbeierle
209 Wed 12-Mar-2003status March 12all stnsbeierle
208 Tue 11-Mar-2003status March 11all stnsbeierle
207 Mon 10-Mar-2003status March 7all stnsbeierle
206 Mon 10-Mar-2003status March 6all stnsbeierle
205 Mon 10-Mar-2003status March 5all stnsbeierle
204 Mon 10-Mar-2003status March 4all stnsbeierle
203 Thu 06-Mar-2003snowingoncley
202 Thu 06-Mar-2003cleaned kryptonlakeoncley
201 Thu 06-Mar-2003clean radslakeoncley
200 Thu 06-Mar-2003CO2 cal gas amountstoweroncley
199 Thu 06-Mar-2003krypton cleanedsageoncley
198 Thu 06-Mar-2003rads cleanedsageoncley
197 Thu 06-Mar-2003cleaned krypton at 2mtoweroncley
196 Thu 06-Mar-2003rad cleanedtoweroncley
195 Thu 06-Mar-2003still windy with blowing snowoncley
194 Wed 05-Mar-2003blowing snowoncley
193 Wed 05-Mar-2003KingAir flew this AMoncley
192 Wed 05-Mar-2003some clouds this AMall stnsoncley1
191 Tue 04-Mar-2003swapped 0.5-1.0 TRHtoweroncley
190 Tue 04-Mar-2003Lake site surveylakeoncley
189 Tue 04-Mar-2003Gravimetric soil samplesall stnsoncley
188 Tue 04-Mar-2003T/RH changetoweroncley
187 Tue 04-Mar-2003snowingall stnsoncley
186 Sun 02-Mar-2003Daily status, feb 28beierle
185 Sun 02-Mar-2003Daily status, Feb 27beierle
184 Sun 02-Mar-2003Daily status, Feb 26beierle
183 Sun 02-Mar-2003king air flew todayoncley
182 Sun 02-Mar-2003fog/frostall stnsoncley
181 Sat 01-Mar-2003analysis of T.5m fantoweroncley1
180 Sat 01-Mar-2003fog this AMall stnsoncley
179 Fri 28-Feb-2003analysis thoughtsoncley
178 Fri 28-Feb-2003GPS testingsageoncley
177 Fri 28-Feb-2003weird 1m sonic spectratoweroncley1
176 Fri 28-Feb-2003low clouds just breaking uponcley
175 Thu 27-Feb-2003Intermittent GPS outages at sagesageoncley
174 Thu 27-Feb-2003USFS station further Southtoweroncley
173 Thu 27-Feb-2003brushed snow caps from 15-30mtoweroncley
172 Thu 27-Feb-2003CO2 cal flows and supplimentary calstowerdelany
171 Thu 27-Feb-2003Boom trolleytowerdelany
170 Thu 27-Feb-2003station visitsageoncley
169 Wed 26-Feb-2003T/RH's on towertowerdelany
168 Wed 26-Feb-2003Removing radiometer snow doughnutsall stnsdelany
167 Tue 25-Feb-2003Daily status, Feb 25all stnsbeierle
166 Tue 25-Feb-2003Site inspection after snowall stnsdelany
165 Mon 24-Feb-2003Daily status, Feb 24all stnsbeierle
164 Mon 24-Feb-2003Deep soil moisture at Lakelakedelany
163 Fri 21-Feb-2003Daily status, Feb 21all stnsbeierle
162 Wed 20-Feb-2003Daily status, Feb 20all stnsbeierle
161 Wed 19-Feb-2003Daily status, Feb 19all stnsbeierle
160 Sat 22-Feb-2003power changestoweroncley
159 Sat 22-Feb-200310V regulators addedall stnsoncley
158 Sat 22-Feb-2003TRH work - finding why 10m is a maxtoweroncley
157 Fri 21-Feb-2003Tasks completedtowerdelany
156 Fri 21-Feb-2003Cosmos ADAM fan replacedtowerdelany
155 Fri 21-Feb-2003Weather conditionsall stnsmahrt
154 Thu 20-Feb-2003Frost detectors installedtowerdelany
153 Thu 20-Feb-2003Supplimentary CO2 calibrationstowerdelany
152 Thu 20-Feb-2003Deep soil moisture at Sagesagedelany
151 Wed 19-Feb-2003Cleaned boom pigstowerdelany
150 Wed 19-Feb-2003New sonic boom angles at towertowersemmer
149 Wed 19-Feb-2003change disk drivesbasesemmer
148 Wed 19-Feb-2003COSMOS rebootedtowersemmer
147 Wed 19-Feb-2003thermocouples runningsagesemmer
146 Wed 19-Feb-2003CO2 flow rates adjustedtowerdelany
145 Wed 19-Feb-2003 Kr cleanedall stnsdelany
144 Tue 18-Feb-2003Thermocouples at Sagesagesemmer
143 Tue 18-Feb-2003cosmos down at 1:30 pm localtowersemmer
142 Tue 18-Feb-2003Guy tensions checkedtowerdelany
141 Tue 18-Feb-2003Tower beacontowerdelany
140 Tue 18-Feb-2003Radiometer cleaningall stnsdelany
139 Tue 18-Feb-2003CO2 calibration gasestowerdelany
138 Thu 06-Feb-2003Daily status, Feb 5all stnsbeierle
137 Wed 05-Feb-2003Daily status, Feb 5all stnsbeierle
136 Wed 05-Feb-2003CO2 cal gas statustowerdelany
135 Tue 04-Feb-2003Daily status, Feb 4all stnsbeierle
134 Mon 03-Feb-2003Daily status, Feb 3all stnsbeierle
133 Sat 01-Feb-2003Licor CO2 back to lifetowermaclean
132 Fri 31-Jan-2003Daily status, Jan 31all stnsbeierle
131 Thu 30-Jan-2003Daily status, Jan 30all stnsbeierle
130 Thu 30-Jan-2003Daily status, Jan 29all stnsbeierle
129 Thu 30-Jan-2003Daily status, Jan 28all stnsbeierle
128 Tue 28-Jan-2003Sage / Lake soil cable problem checkslakemilitzer
127 Tue 28-Jan-2003Sonic Level offsets: lake, 15mlakemilitzer
126 Tue 28-Jan-2003Cosmos Cooler / overheating fixedtowermilitzer
125 Tue 28-Jan-2003Radiometers Cleanedall stnsmilitzer
124 Tue 28-Jan-2003Kryptons cleanedall stnsmilitzer
123 Tue 28-Jan-2003Grass TP01 soil properties fixedtowermilitzer
122 Mon 27-Jan-2003Jan 27 work: 25m rad. installed/fixed, t.10m, calgas note, guysall stnsmilitzer
121 Mon 27-Jan-2003Daily status Jan 27all stnsbeierle
120 Fri 24-Jan-2003Daily status Jan 24all stnsbeierle
119 Thu 23-Jan-2003Daily status Jan 23all stnsbeierle
118 Wed 22-Jan-2003Soil sensor configurationall stnsdelany1
117 Wed 22-Jan-2003Daily status Jan 22all stnsbeierle
116 Wed 22-Jan-2003TP01 wiringall stnssemmer1
115 Tue 21-Jan-2003Daily status Jan 21all stnsbeierle
114 Fri 17-Jan-2003Daily status Jan 17all stnsbeierle
113 Thu 16-Jan-2003Daily status Jan 16all stnsbeierle
112 Wed 15-Jan-2003Daily status Jan 15all stnsbeierle
111 Tue 14-Jan-2003Daily status Jan 14all stnsbeierle
110 Mon 13-Jan-2003krypton at 20 meterstowermaclean
109 Fri 10-Jan-2003Daily status Jan 10all stnsbeierle
108 Thu 09-Jan-2003Daily status Jan 9all stnsbeierle
107 Thu 09-Jan-2003riming eventstoweroncley
106 Wed 08-Jan-2003Daily status Jan 8all stnsbeierle
105 Wed 08-Jan-2003Shortened CO2 cal periodstowermaclean
104 Wed 08-Jan-2003yesterday's tasksall stnsoncley
103 Wed 08-Jan-2003Daily status Jan 7all stnsbeierle
102 Tue 07-Jan-2003Daily status Jan 6all stnsbeierle
101 Tue 07-Jan-2003testing of thermal properties probesall stnsmaclean
100 Tue 07-Jan-2003other needsbaseoncley
99 Tue 07-Jan-2003print needs color cartridgebaseoncley
98 Tue 07-Jan-2003sunriseall stnsoncley
97 Tue 07-Jan-2003fogall stnsoncley
96 Mon 06-Jan-2003visit taskstoweroncley
95 Fri 03-Jan-2003Daily Status Jan 3all stnsbeierle
94 Fri 03-Jan-2003Calibration gas regulator installationtowerdelany
93 Thu 02-Jan-2003Daily Status Jan 2all stnsbeierle
92 Tue 31-Dec-2002logger filesall stnsoncley
91 Sun 29-Dec-2002sonic tiltsall stnsoncley9
90 Fri 27-Dec-2002no rain tips from Nov 20 to Dec 16towermaclean
89 Fri 27-Dec-2002tower stufftoweroncley
88 Fri 27-Dec-2002Daily Status Dec 27all stnsoncley
87 Tue 24-Dec-2002stopped krypton_heat scripttowermaclean
86 Sun 22-Dec-2002krypton heater at 2m on towertowermaclean
85 Thu 19-Dec-2002Daily Status Dec 19all stnsbeierle
84 Thu 19-Dec-2002Corrected therm prop. timing at lakelakemaclean
83 Thu 19-Dec-2002Checked logger coefs at 25mtowermaclean
82 Thu 19-Dec-2002Daily Status Dec 18all stnsbeierle
81 Thu 19-Dec-2002Daily status Dec 17all stnsbeierle
80 Wed 18-Dec-2002Soil gravimetric/bulk densityall stnsdelany
79 Wed 18-Dec-2002sage visitsageoncley
78 Wed 18-Dec-2002Lake visitlakeoncley
77 Wed 18-Dec-2002Boom PYG mount Modificationtowerrucker
76 Tue 17-Dec-2002snowing todayall stnsoncley
75 Tue 17-Dec-2002EVE Q1&Q2oncley
74 Tue 17-Dec-2002electric fence oktowerrucker
73 Tue 17-Dec-20024th solar panel installedsageoncley
72 Tue 17-Dec-20024m rads uptoweroncley
71 Tue 17-Dec-2002Installation of boom pyrgeometerstowerdelany
70 Tue 17-Dec-2002Gravimetric soil samplesall stnsdelany
69 Mon 16-Dec-2002T/RH swapped backtoweroncley
68 Mon 16-Dec-2002CO2 setuptowerdelany
67 Mon 16-Dec-2002Gravimetric soil samplesall stnsdelany
66 Mon 16-Dec-2002stufflakeoncley
65 Mon 16-Dec-2002Wind generator swappedlakerucker
64 Mon 16-Dec-2002krypton heater addedtoweroncley
63 Mon 16-Dec-2002aster rebootedbaseoncley
62 Mon 16-Dec-2002EVE rebootedlakeoncley
61 Mon 16-Dec-2002TP01 installedlakeoncley
60 Mon 16-Dec-2002TP01 installedtoweroncley
59 Mon 16-Dec-2002Daily status, Dec 16all stnsbeierle
58 Mon 16-Dec-2002krypton heater scripttowermaclean
57 Mon 16-Dec-2002Power glitch caused adam crashtowermaclean
56 Fri 13-Dec-2002Daily status, Dec 13all stnsbeierle
55 Thu 12-Dec-2002Added dat function to correct Rlw.in at lakelakemaclean
54 Thu 12-Dec-2002Changed Tsoil sampling at sage,lakelakemaclean
53 Thu 12-Dec-2002Daily status, Dec 12all stnsbeierle
52 Wed 11-Dec-2002Daily status, Dec 11all stnsbeierle
51 Wed 11-Dec-2002Modifed grass logger Tsoil samplingtowermaclean
50 Tue 10-Dec-2002Daily status, Dec 10all stnsbeierle
49 Mon 09-Dec-2002Daily status, Dec 9all stnsbeierle
48 Mon 09-Dec-2002replaced disk drivebaseknudson
47 Mon 09-Dec-2002removed uplooking pir sensorlakeknudson
46 Mon 09-Dec-2002soil profile array cable installedtowerknudson
45 Mon 09-Dec-20022m TRHtowerknudson
44 Fri 06-Dec-2002Pressure errortoweroncley1
43 Fri 06-Dec-2002Daily status, Dec 6all stnsbeierle
42 Thu 05-Dec-2002Daily status, Dec 5all stnsbeierle
41 Thu 05-Dec-2002Daily status, Dec 4all stnsbeierle
40 Tue 03-Dec-2002Daily status, Dec 3all stnsbeierle
39 Mon 02-Dec-2002Daily status, Dec 2all stnsbeierle
38 Sun 01-Dec-2002Tcase&Tdome.in coefs at lakelakemaclean
37 Wed 27-Nov-2002New settings for cameratowermaclean
36 Wed 27-Nov-2002Camera white balance documentationtowermaclean
35 Tue 26-Nov-2002Daily status, Nov 26all stnsbeierle
34 Tue 26-Nov-2002color balance on axis cameratowermaclean
33 Mon 25-Nov-2002Daily status, Nov 25all stnsmaclean
32 Fri 22-Nov-2002Fix lake coefficients for pyginlakesemmer
31 Thu 21-Nov-2002Swapping USB data disksbasemaclean
30 Thu 21-Nov-2002Snow depth gaugetowerknudson
29 Thu 21-Nov-2002Rnet at lakelakemaclean
28 Thu 21-Nov-2002PSPs swapped at sagesagemaclean
27 Thu 21-Nov-2002Krypton Placementall stnsmilitzer
26 Thu 21-Nov-2002Stn2 / Stn3 Rad Fans / Icingall stnsmilitzer
25 Thu 21-Nov-2002Sonic boom anglestowermaclean
24 Thu 21-Nov-2002PAM Stn2 Reconfigurationsagemilitzer
23 Thu 21-Nov-2002PAM Stn3 Reconfigurationlakemilitzer
22 Wed 20-Nov-2002rad stand uptoweroncley
21 Wed 20-Nov-2002TRH heights on towertowermaclean
20 Mon 18-Nov-2002November setupall stnsmaclean
19 Fri 15-Nov-2002Greasewood 1.5m radiation CR10X codetowersemmer
18 Fri 15-Nov-20024m radiation CR10X codetowersemmer
17 Fri 15-Nov-200225 meter radiation CR10X codetowersemmer
16 Fri 15-Nov-2002Grass, Lake, and Sage CR10X codeall stnssemmer
15 Fri 15-Nov-2002PAM sage CR10X coffssagesemmer
14 Fri 15-Nov-2002PAM Lake CR10X coffslakesemmer
13 Fri 15-Nov-2002tower-grass CR10X coffstowersemmer
12 Tue 12-Nov-2002Axis camera settingstowermaclean
11 Mon 11-Nov-2002Krypton Calsall stnssemmer
10 Fri 08-Nov-2002Cosmostowermaclean
9 Fri 08-Nov-2002Daisytowermaclean
8 Mon 04-Nov-2002PHONE #'sbasemilitzer
7 Mon 14-Oct-2002FLOSS02 soil sensorsall stnsdelany
6 Fri 11-Oct-2002Lake Site Visit : + Boom Angleslakemilitzer
5 Fri 11-Oct-2002Sage Site Visit : + Boom Anglessagemilitzer
4 Fri 11-Oct-2002Hft Probes 1,2 upside downlakemilitzer
3 Fri 11-Oct-2002Soil Echo Probes Rotated to Verticalall stnsmilitzer
2 Fri 11-Oct-2002GUY WIRE TENSIONStowerknudson
1 Tue 08-Oct-2002Installation of soil sensorsall stnsdelany


229: Radiometers, Site , Tue 22-Apr-2003 14:39:34 MDT, pyg IC now okay
We have just fixed a few cabling problems: in.4m.bare and grease and now think
that all pygs are working properly.

228: Radiometers, Site all stns, Tue 22-Apr-2003 10:05:37 MDT, pyg intercomparison

the pyg intercomparison was set up Monday on the penthouse roof.
all radiometers were mounted looking up - one psp (K&Z 970377, Lake)
 and 13 pygs, as follows:

	Grass	 	P26416 pin
			P29260 pout
	Bare (4M)	31981 pin1
			31980 pout1
			31976 p1.5m
	Grease		31978 pout
	25M		31975 pin1
			31974 pout1
			31979 pin2
			31977 pout2
	Sage		P26417 pin
			P27907 pout
	Lake		P29137 - not used - went bad during the project
			P29136 pout - note on radiometer - comp circ removed

	4 t/rh's were also set up to complete their intercomparison

			TRH 003 baseline
			TRH 206 baseline
			TRH 502 PamIII
			TRH 101 PamIII


227: Temp_RH, Site all stns, Thu 10-Apr-2003 15:40:56 MDT, TRH intercomparison
  
  4/4/03
  
  the trh's were set up on the penthouse roof on a single sawhorse
to do a sensor intercomparison.
The channels are as follows:
	Main Tower:
		TRH
ch 00	.5m	001	has broken bell
ch 01	1m	206	has broken bell
ch 02	2m	104	has broken bell
ch 03	5m	701 	has broken bell
ch 04	10m	702
ch 05	15m	007 
ch 06	20m	704	replaced fan - no good since before teardown?
ch 07	25m	008 
ch 08	30m	003

ch 09	Sage:	101	disconnected initialy due to lack of inlet tubes 

ch 10	Lake:	502	disconnected initialy due to lack of inlet tubes

ch 11  cal lab  703
ch 12  cal lab	004	has broken bell
ch 13  cal lab  204




226: Status, Site all stns, Mon 07-Apr-2003 10:07:20 MDT, TearDown Sensor IDs
Main Tower:
	CSAT	KH2o	TRH
.5m	674		001
1m			206
2m	539	1389	104
5m	673		701 (my notes show this trh at 6m?)
10m	672	1391	702
15m	671		007 (this trh boom folded up against North leg :-(
20m	677	1390	704
25m	??	??	008 (csat/kh2o packed before id'd)
30m			003

Sage:	536	1395	101	soil logger profile #3
				Rnet   94196
				RswIn  940187
				RswOut 970378
				RlwIn  26417
				RlwOut 27907

Lake:	537	1397	502	soil logger profile #6
				Rnet   94197
				RswIn  970377
				RswOut    186
				RlwOut 29136

Bare:				RlwOut 31976 (individual sensor)
Greasewood:			RlwOut 31978 (individual sensor)

We did not record the other individual radiometer components
since they are identified by site/box.



225: Status, Site all stns, Tue 01-Apr-2003 16:00:26 MST, daily status March 31
Weather Plots:  OK

Rain Gauge: OK

Vbatt: OK

Sonics 2m: OK

Sonics profile: OK

Kryptons: OK, there was a large diurnal variation.

Sonic variance (2m): OK

H2O-kh2o: OK  

Tv and P:  OK

Scalar variance: OK

2m Sonic QC: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2:  OK.

TRHdiff 2m:  OK

TRH_diff_profiler: OK

Pressure, Elect Temp: OK

Thermocouples, Ice Detectors:  OK

Radiation stns 1-3: OK

Long Wave Rad:. OK

|Tdome-Tcase| : OK

Tsoil: OK

Qsoil: Qsoil.b.10cm.tower is having problems. It experiances large increases from ~20%vol to between 100-120% vol during the day.

Gsoil:Gsoil.b.10cm.sage is bad and b.10cm.lake still isn't working properly.  

Soil Thermal Properties: Some strange glitches still appear, especially in the third vpile plot near zero, but they do not last longer than a few hours at a time?

T Profiles: OK

RH profiles:  OK

u,v spectra, stn 1-3: OK. Some white noise from boht v and u.2m at frequencies greater than 10. 

w, tc spectra: OK

u:w, w:tc cospectra: OK

u,v spectra tower: At lower frequencies the traces do not follow the 2/3 line very well. There is also a small amount of white noise about 10 Hz.

w, tc spectra tower: OK 

kryton spectra: 0000-0030 plot doesn't look as clean there is more variaton between the sensors.  kh2o.2m.lake is measuring a bit higher than the other sensors, and above 1.00 Hz kh20.2m kind of trails off on its own.

w,kh2o cospectra: OK

CO2 cals and spectra: CO2 variance level is below the noise level of the sensor. 


















224: Log, Site sage, Fri 28-Mar-2003 09:50:18 MST, Tcam study
2003, Mar 26, Wed
Tony Delany and Steve Oncley visited FLOSS02 to undertake an experiment
to view the instrumented sagebush using the thermal imaging camera.
Arrived FLOSS site at 15:10 
No snow on ground at Sage site.

Set up TCam on tripod 2.5m from sagebrush, approx 1m above surface.
Adjusted range and spread to view surface, stems and foliage.
Did the Non Uniformity Correction.
16:21:44 GPS time started 100 images at 10s intervals.
Trange 12.1 - 4.2 degeeC.
Allowed to run for ~ 70 images and then carried Tcamera to 3m using ladder.
Took 8 - 10 overlapping images looking down on bush and surrounds.
Took 8 -10 overlapping images looking under Darkhorse.
Matched the thermal images with digital camera shots of the same areas.

Set TCam back on tripod
16:46:32 GPS time started 100 images at 1 min intervals
Trange 15.0 - -1.3 degreeC
Ran until 18:25 

18:30;30 started 210 images at 1 min intervals
Snow started to fall
Left site for dinner, Snow began to fall heavilly
Returned at 20:30, ~ 5 cm of snow on ground, snow encrusted on side but non
on lens. Decided to allow camera to complete the run.
At 22:00 snow now cold and powdery.

Started night time run 22:06:00 210 images at 3 min intervals
TRange 4.2 - -6

Mar 27, Thurs
Arrived at Sage at 06:00, TCam had no snow on it.
Viewfinder image of bush was "fuzzy" although image of a hand was sharp.
Stopped at 161/210 at ~ 06:05.
Did Non Uniformity Correction.

Started at 06:12:20 210 images at 30 s interval
TRange 1.3 - > -10 degreeC
Stopped at 07:59

Reset TRange to 4.2 - <-10 degreeC
Start 08:02:30, 210 images at 30 s interval
Finished by 09:59

Retained Trange 4.2 - >-10 degreeC
Start 10:04:20 210 images at 30s intervals

 IR Camera runs for PM Wed, 26 Mar 03 through AM Thu, 27 Mar 03
 -------------------------------------------------------------------
 Run  Start       Stop     # images   Interval Trange               Object  
 1    16:21                ~ 70        10s      12.1 - 4.2           Sagebush
                  16:44    ~ 20        10s      12.1 - 4.2           Top ofSagebush/
		                                                       Below Rad stand
 2    16:46       18:25     210        1min     15 - -1.3            Sagebush
 3    18:30       22:00     210        1min     15 - -1.3            Sagebush
 4    22:06 27Mar 06:05     210        3min     4.2 - -6             Sagebush
 5    06:12       07:59     210        30s     -1.3 - >-10 [-11.5]   Sagebrush
 6    08:02       09:59     210        30s      4.2 - >-10 [-13.9]   Sagebush
 7    10:04      ~11:45     210        30s      4.2 - >-10 [-13.9]   Sagebrush

223: Log, Site tower, Fri 28-Mar-2003 09:49:17 MST, tower site visit
2003, Mar 26 Wed
TD and SO
Arrived at FLOSS for thermal camera study (see next logbook entry)
While TCam operating walked to tower to check out the state of access.
Upper stretch muddy but passable. The ditch was only slightly muddy.
Three areas after the bend where the lake drainage flowed were bad, bad.
SO has digital images.

Sage site: no snow prior to evening ppt
Grease site: only a portion of the drift remaining, water standing on surface.
Grass: no snow on surface prior to evening snow. Ground thawed to 10cm and frozen below.

CO2
S 420psi  960psi N
intake 7 lpm

222: Power, Site tower, Fri 28-Mar-2003 09:46:15 MST, power died
[the next 3 entries were written on 26&27 March, but couldn't be entered until
now.]

Apparently, there were 3 short power hits about 19:30 on 26 March which caused 
the adams to reboot twice.  Lake & Sage (of course) stayed up, and the UPS kept
the base up, so only a bit of data were lost and only from the tower.

Before it is erased, we should: "fgrep ups /var/log/messages > ups_log_entry"
and include this in logbook to record the occurance of all of these events.

221: Temp_RH, Site tower, Fri 28-Mar-2003 09:44:16 MST, bad T/RH fan at 20m
About 17 March, the fan at 20m started dying.  Even though we were there last 
Wed-Thursday, we didn't realize this until the tower was too icy to climb.
Thus, this level will remain bad through the rest of the program.  Sorry...

220: Status, Site all stns, Thu 27-Mar-2003 15:44:08 MST, daily status March 27
Weather Plots:  OK

Rain Gauge: OK

Vbatt: OK

Sonics 2m: Lake 2m anemometer went a little crazy over night. There is a lot of variability between 1800 MST 3/36 and 1100 MST 2/27, and then it calms back down?

Sonics profile: OK

Kryptons: OK, there was a large diurnal variation.

Sonic variance (2m): for w'w', v'v', and u'u' lake hit it's threshold between 1800-1200 MST.  

H2O-kh2o: OK  

Tv and P:  OK

Scalar variance: Very high measurements here also.  For tc'tc' lake hit it's threshold and remained there between 1800 MST on 3/26 and 1200 on 3.27.  For kh2o'kh2o' all three sensors maxed out from time to time during the same time period. 
(it was the same for the scalar variance profile) 
2m Sonic QC: sage had some pretty prominent flags between 1000 and 1300 MST

Fluxes at 2m: Lake hit the threshold throughout the night. 

Flux profiles: OK

CO2:  Dropped down to 200 at 1900 MST 3/26 but recovered at 2000 MST.

TRHdiff 2m:  There is a small gap in the data from ~1930 to 2000 MST on 3/26.

TRH_diff_profiler: OK

Pressure, Elect Temp: OK

Thermocouples, Ice Detectors:  OK

Radiation stns 1-3: Rsum.tower(1) and Rsum.lake(3) fall out of line with the other sensors at ~1300 MST. 

Long Wave Rad:. OK

|Tdome-Tcase| : OK

Tsoil: Temps of b.2.5.sage, b.5cm.sage, b.10cm.sage all increases fairly abruptly at 1200 and are slowly decreasing. 


Qsoil: OK

Gsoil:Gsoil.b.10cm.sage is bad and neither a.10cm.lake nor b.10cm.lake are working properly. 


Soil Thermal Properties: Some strange glitches still appear from time to time, especially in the third vpile plot near zero, but they do not last for long.

T Profiles: OK

RH profiles:  OK

u,v spectra, stn 1-3: 2m.lake trace off quite a bit from the others. 2m and lake are measuring ~.5 while lakes range is between 1-100. 

w, tc spectra: same as u,v spectra

u:w, w:tc cospectra: lake trace again has a lot of variability and far off other sensors.

u,v spectra tower: u.2m and v.2m begin to deviate from the other traces just before .100 Hz but only between 0000-0030. White noise at frequencies greater than 10 Hz

w, tc spectra tower: same as u,v,specta tower 

kryton spectra: 0000-0030 plot is missing a chunck of data just about the .001 frequency.  During 1200-1230 neither kh2o.2m or kh2o.sage look very good.

w,kh2o cospectra: lake cospectra is show a lot of variabilty compared with other sites. Range is between -1 and 3.5. 

CO2 cals and spectra: CO2 variance level is below the noise level of the sensor. 















219: Status, Site all stns, Wed 26-Mar-2003 15:59:09 MST, daily status March 26
Weather Plots:  OK

Rain Gauge: No rain, or snow 

Vbatt: OK

Sonics 2m: OK.  

Sonics profile: OK

Kryptons: OK, no diurnal variation.

H2O-kh2o: OK  

Tv and P:  OK

Scalar variance: OK

2m Sonic QC: sage had some pretty prominent flags between 1000 and 1300 MST

Fluxes at 2m: OK

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  OK.

TRH_diff_profiler: OK

Pressure, Elect Temp: OK

Thermocouples, Ice Detectors:  OK

Long Wave Rad:. Rlw.out.grease stn0 is/was again measuring less outgoing radiation than the other sensors during the day.

|Tdome-Tcase| : OK

Tsoil: OK. 


Qsoil: OK

Gsoil:Gsoil.b.10cm.sage is bad and neither a.10cm.lake nor b.10cm.lake are working properly. 


Soil Thermal Properties: Some strange glitches still appear from time to time, especially in the third vpile plot near zero, but they do not last for long.

T Profiles: OK

RH profiles:  OK

u,v spectra, stn 1-3: 2m sensors experiance some white noise above 10 Hz.

w, tc spectra: OK

u:w, w:tc cospectra: OK

u,v spectra tower: OK.  White noise at frequencies greater than 10 Hz

w, tc spectra tower: Good, some white noise from tc spectra above 10 Hz. 

kryton spectra: Quite a bit of white noise at the higher frequencies. 

CO2 cals and spectra: CO2 variance level is below the noise level of the sensor. 














218: Status, Site all stns, Tue 25-Mar-2003 14:09:57 MST, daily status March 25
Weather Plots:  OK

Rain Gauge: OK

Vbatt: OK

Sonics 2m: OK.  

Sonics profile: OK

Kryptons: OK, no abrupt diurnal variation.

H2O-kh2o: OK  

Tv and P:  OK

Scalar variance: OK

2m Sonic QC: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  OK.

TRH_diff_profiler: OK

Pressure, Elect Temp: OK

Thermocouples, Ice Detectors:  OK

Long Wave Rad:. Rlw.out.grease stn0 seems to be OK today.

|Tdome-Tcase| : OK

Tsoil: the temp of some of the sage sensors (b.2.5cm, b.5cm, c.5cm and b.10cm) increased rapidly at just after 1000 MST (the b.2.5cm sensor climbed from ~0 deg to almost 10 deg) while other sensors stayed the same temp. 


Qsoil: OK

Gsoil:Gsoil.b.10cm.sage is bad and b.10cm.lake is not working properly, there is definately something wrong. 


Soil Thermal Properties: OK 

T Profiles: OK

RH profiles:  OK

u,v spectra, stn 1-3: 2m sensor experiances some white noise above 10 Hz.

w, tc spectra: OK

u:w, w:tc cospectra: OK

u,v spectra tower: Look great.  Are all in good agreement. White noise at frequencies greater than 10 Hz

w, tc spectra tower: Good, some white noise from tc spectra above 10 Hz. 

kryton spectra:OK

CO2 cals and spectra: CO2 variance level is below the noise level of the sensor. 












217: Status, Site all stns, Mon 24-Mar-2003 14:55:29 MST, daily status March 24
Weather Plots:  OK

Rain Gauge: OK

Vbatt: OK

Sonics 2m: OK.  

Sonics profile: OK

Kryptons: Look pretty bad at night.

H2O-kh2o: OK  

Tv and P:  OK

Scalar variance: kh2o'kh2o'.2m, all sites, have some pretty large spikes throughout the night from 2000 MST on 3/23 to 0600 MST on 3/24. 

2m Sonic QC: Flag of lake data between 2100 and 2300 on 3/23

Fluxes at 2m: OK

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  OK.

TRH_diff_profiler: 20m sensor has a temperature spike (~3-4 deg)and humidity spike (~15-20% betweeen 1100 and 1200 MST. I can't really decide if the 30m sensor trace looks ok.  I think sensor may be acting up a bit.

Pressure, Elect Temp: OK

Thermocouples, Ice Detectors:  OK

Long Wave Rad:. Rlw.out.grease stn0 is measuring less outgoing radiation than the other sensors during the day.

|Tdome-Tcase| : out.bar.stn0 had a glitch/spike to 1.5 at 2100 MST 3/23.

Tsoil: OK


Qsoil: OK

Gsoil:Gsoil.b.10cm.sage is bad and both b.10cm.lake and a.10cm.lake are still having problems. They were off on Sunday 3/23 but came back into line with other sensors around 2100 MST that night. 


Soil Thermal Properties: Still quite a bit of oscillation near zero in some of the vpile plots. Vpile plots from lake have negative values again. 

T Profiles: OK

RH profiles:  OK

u,v spectra, stn 1-3: 2m sensor experiances some white noise above 10 Hz.

w, tc spectra: OK

u:w, w:tc cospectra: OK

u,v spectra tower: Look great.  Are all in good agreement.

w, tc spectra tower: Good, some white noise from tc spectra above 10 Hz. 

kryton spectra:kh20.2m specta from 0000-0030 is off, compared with the other traces. 

CO2 cals and spectra: CO2 variance level is below the noise level of the sensor. 










216: CO2, Site tower, Sat 22-Mar-2003 10:38:29 MST, Set CO2 lag to 1 sec

The w,CO2 correlation plots have shown an approximate lag of 1 second
in the CO2 data, so I added a l=1.0 parameter in prep.config
for co2mr. It had the desired effect in today's correlation
vs lag plot - it shifted the peak back almost to 0.
215: Status, Site all stns, Fri 21-Mar-2003 15:55:28 MST, daily status March 21
Weather Plots:  OK

Rain Gauge: OK

Vbatt: OK

Sonics 2m: OK.  

Sonics profile: OK

Kryptons: OK, there was diurnal variation.

H2O-kh2o: OK  

Tv and P:  OK

Scalar variance: kh2o'kh2o'.2m, both lake and tower shot up to 3.5 or higher at times, between 1700 MST on 3/20 and 0300 MST on 3/21. 

Fluxes at 2m: tower showed a great deal of variation between 2000 MST on 3/20 and 0300 MST on 3/21?

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  OK.

TRH_diff_profiler: 20m sensor is reading slightly warmer (~3-4 deg)and more more moist _(~15-20%)than other sensors betweeen 0900 and 1500 MST.

Pressure, Elect Temp: OK

Thermocouples, Ice Detectors:  OK

Long Wave Rad:. OK

|Tdome-Tcase| : .OK

Tsoil: OK


Qsoil: Qsoil.b.10cm.tower(1) had two 100% glitches. Data recovered shortly after.

Gsoil:Gsoil.b.10cm.sage is bad and both b.10cm.lake and a.10cm.lake are acting strangely. I will keep my eye on them. 


Soil Thermal Properties: Still quite a bit of oscillation near zero in some of the vpile plots. 

T Profiles: 20m sensor is reading warmer than other sensors on the tower through out the day.  Greatest difference is about 6 deg at 1300 MST. 

RH profiles:  OK

u,v spectra, stn 1-3: OK 

w, tc spectra: tc spectra between 0000-0030 has some white noise at frequencies
greater than one.

u:w, w:tc cospectra: OK

u,v spectra tower: Look great.  Are all in good agreement.

w, tc spectra tower: Good, some white noise at 0000-0030. 

kryton spectra: Looking at 0000-0030 plot kh20.2m specta is off on it's own, compared with the other traces. The w,kh2o cospectra plot at 0000-0030 shows that w.2m.kh2o.2m has a lot more variation than the other sensors. 

CO2 cals and spectra: CO2 variance level is below the noise level of the sensor. 









214: Status, Site all stns, Fri 21-Mar-2003 14:26:22 MST, daily status March 20
Weather Plots:  OK

Rain Gauge: OK

Vbatt: OK

Sonics 2m: OK.  

Sonics profile: OK

Kryptons: OK, there was diurnal variation.

H2O-kh2o: OK  

Tv and P:  OK

Scalar variance: kh2o'kh2o'.2m, both lake and tower shot up to 3.5 or higher at times, between 1700 and 2000 MST 

Fluxes at 2m: tower showed a great deal of variation between 2000 MST on 3/20 and 0300 MST on 3/21

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  OK.

TRH_diff_profiler:  OK

Pressure, Elect Temp: oscillation of tbox data is result of the coolers. It is now warm enough at the site for them to kick on .

Thermocouples, Ice Detectors:  OK

Long Wave Rad:. OK

|Tdome-Tcase| : .OK

Tsoil: OK


Qsoil: OK  

Gsoil:Gsoil.b.10cm.sage is bad and both b.10cm.lake and a.10cm.lake are acting strangely. I will keep my eye on them. 


Soil Thermal Properties: OK 

T Profiles: OK

RH profiles:  OK

u,v spectra, stn 1-3: OK 

w, tc spectra: OK

u:w, w:tc cospectra: OK

u,v spectra tower: Look great.  Are all in good agreement.

w, tc spectra tower: Good, some white noise at 0000-0030. 

kryton spectra: OK

CO2 cals and spectra: OK 






213: Status, Site all stns, Fri 21-Mar-2003 09:52:40 MST, status March 17
Weather Plots:  OK

Rain Gauge: OK

Vbatt: OK

Sonics 2m: OK.  

Sonics profile: OK

Kryptons: There is a lot of variation. There are many spikes and dips. Looks like the instruments are having trouble. 

H2O-kh2o: OK  

Virtual Temp: Missing data between ~0500 MST and ~1500 MST. 

uvw variance 2m::  u'u'.2m and v'v'.2m have data spikes as large as 9 m/s^2 between 0500 and 1300 MST.

uvw profile variance: Have similar data spikes, same magnitude, during same time period.

Scalar variance (2m and profile): Large values, and a lot of variation from 3/16 to 3/17 at 1500 MST.  Things died down after that. 

Sonic QC 2m: Must have either been very gusty or there are data errors.  There are many flags between 0400 and 1500 MST.  

Fluxes at 2m: OK

Flux profiles: OK

CO2:  Seemed to be misbehaving between 3/16 1700 MST and 0700 on 3/17.

TRHdiff 2m and Profile:  OK.

TRH_diff_profiler:  30m sensor may be malfunctioning, it doesn't follow along with the other sensors, and at times reads more moist then the 1m sensor. Need to keep an eye on. 

Pressure, Elect Temp: oscillation of tbox data is result of the coolers. It is now warm enough at the site for them to kick on .

Thermocouples, Ice Detectors:  OK

Long Wave Rad:. OK

|Tdome-Tcase| : .OK

Tsoil: Tsoila.10cm sage looks to have crashed.  Tsoil.b.10cm lake may have crashed also.  Neither of the sensors seem to be recording good data.


Qsoil: Strange glitch from Qsoil.c.10cm tower at 1300 MST.   

Gsoil:Gsoil.b.10cm.sage is still not measuring properly. Gsoil.b.10cm. lake is having problems also. Will keep and eye on.

Soil Thermal Properties: Instruments are still behaving strangely. No large glitches like seen in the previous two days, but I will still keep and eye on them. 

T Profiles: OK

RH profiles:  OK

u,v spectra, stn 1-3:  u spectra at 12:00-12:30 doesn't look very good. The datafrom each site vary from one another and lake and 2m don't follow the line particularly well.  

w, tc spectra: neither spectra at 1200-1230 look very good. Same problem as with uv specta. Lake and 2m specta vary from sage, and they don't follow the 2/3 line.

u:w, w:tc cospectra: Plots from 1200-1230 ook really bad..not sure what's going on?

u,v spectra tower: some white noise at higher frequencies

w, tc spectra tower: some white noise at higher frequencies

kryton spectra: all sensors, except kh2o.2m.sage have white noise greater than .100 Hz between 0000-0030.  

CO2 cals and spectra:  OK








212: Status, Site all stns, Mon 17-Mar-2003 16:03:10 MST, status March 17
Phone lines are down.  Last data we have is from 0900 MST.  Phone company apparently looking into problem.
211: Status, Site all stns, Fri 14-Mar-2003 14:06:42 MST, status March 14
Weather Plots:  OK

Rain Gauge: OK

Vbatt: OK, only got down to ~12.5

Sonics 2m: OK.  

Sonics profile: OK

Kryptons: OK, no diurnal variation.

H2O-kh2o: OK  

Tv and P:  OK

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  OK.

TRH_diff_profiler:  OK

Pressure, Elect Temp: oscillation of tbox data is result of the coolers. It is now warm enough at the site for them to kick on .

Thermocouples, Ice Detectors:  OK

Long Wave Rad:. Rlw.out.grease stn0 is measuring between 60-80 W/m^2 less outgoing radiation during the day than other sensors.

|Tdome-Tcase| : .OK

Tsoil: Increases in temperature measurements are a result of spring coming. Air temp is increasing and the ground is thawing.


Qsoil: Increased measurements are a result of the ground thawing out with increased temperature.  

Gsoil:Gsoil.b.10cm.sage has gone bad.  Data has not looked right for days and todays data looks the worst. Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.


Soil Thermal Properties: Instruments are still behaving strangely. No large glitches like seen in the previous two days, but I will still keep and eye on them. 

T Profiles: OK

RH profiles:  OK

u,v spectra, stn 1-3:   

w, tc spectra: 

u:w, w:tc cospectra: 

u,v spectra tower: 

w, tc spectra tower: 

kryton spectra:  

CO2 cals and spectra:  





210: Status, Site all stns, Thu 13-Mar-2003 14:53:07 MST, status March 13
Weather Plots:  OK

Rain Gauge: No sign of any data.

Vbatt: OK, only got down to ~12.5

Sonics 2m: OK.  

Sonics profile: OK

Kryptons: OK, no diurnal variation. Again, kh2o.20m.stn0 is measuring more moisture than other sensors (~.5-1 g/m^3 difference)

H2O-kh2o: Something broke the code which affected the H2O-kh2o plot, causing duplicate sensor colors and names.  Gordons is working on the problem.  

Tv and P:  OK

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  OK.

TRH_diff_profiler:  OK

Pressure, Elect Temp: oscillation of tbox data is result of the coolers. It is now warm enough at the site for them to kick on .

Thermocouples, Ice Detectors:  OK

Long Wave Rad:. Rlw.out.grease stn0 is measuring less outgoing radiation during the day than other sensors.

|Tdome-Tcase| : .OK

Tsoil: Increases in temperature measurements are a result of spring coming. Air temp is increasing and the ground is thawing.


Qsoil: Increased measurements are a result of the ground thawing out with increased temperature.  

Gsoil:  OK. Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
Gsoil.b.10cm.sage may be experiencing some problems also. Data from that sensor looks wacky.

Soil Thermal Properties: Behaving strangely. Need to keep and eye one. 
Vpile at both sage and lake suddenly dropped causing glitches of -.08 (sage on 3/12) and -.20 (lake on 3/13)

T Profiles: OK

RH profiles:  OK

u,v spectra, stn 1-3:   OK.  Both spectra at 2m have a little white noise at
frequencies
greater than 10.

w, tc spectra: OK

u:w, w:tc cospectra: wtc cospectra 12:00-12:30 plot shows that w.2m.lake.tc.2m.lake is measuring lower (by ~.03 f*S(m/s degC) than other sensors at frequencies > .05 Hz. 

u,v spectra tower: all in good agreement. Some white noise at f greater than 10 Hz

w, tc spectra tower: OK.  

kryton spectra:  OK, some white noise shown in 0000-0030 plot at frequencies greater than 1 Hz. From 0000-0030 plot kh20.2m is not in line with the other readings.

CO2 cals and spectra:  CO2 variance below the noise level of the sensor. 




209: Status, Site all stns, Wed 12-Mar-2003 15:09:27 MST, status March 12
Weather Plots:  OK

Rain Gauge: No sign of any data.

Vbatt: OK only got down to ~12.5

Sonics 2m: OK.  

Sonics profile: OK

Kryptons: OK, no diurnal variation. Again, kh2o.20m.stn0 is measuring more moisture than other sensors (~1 g/m^3 difference)

H2O-kh2o: OK

Tv and P:  OK

Scalar variance: single data spike from sage at 0200 MST that reached almost .4 (g/m^3)^2

Fluxes at 2m: OK

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  OK.

TRH_diff_profiler:  OK

Pressure, Elect Temp: tbox for both cosmos and daisy begin oscillating around 1000 MST +/-2-3 deg.

Thermocouples, Ice Detectors:  OK

Long Wave Rad:. Rlw.out.grease stn0 is measuring less outgoing radiation during the day than other sensors.

|Tdome-Tcase| : .OK

Tsoil:  Something went wrong with b.2.5cm sage, c.5cm sage, and b.5cm sage.  At 1000 MST mesurements began to increase (as much as 8 deg C in 3 hours) and as of 1400 MST they have not recoved. 

Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil: Sensors are acting strangely. At sage, a.5cm, b.5cm and c.5cm all increase from  ~5 % to ~25 % between 1600 MST 3/11/02 and 1400 MST (which is as much of the plot as I can see at this time). Qsoil.c.5cm.lake is behaving the same although it increased from ~13 % to 35% and looks like it is most probablely still increasing. 

Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil:  OK. Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
May be experiancing some problems also.  Gsoil.a.10cm tower trace is different than what other sensors are measuring.  The rest of the sensors were in good agreement until ~1100 MST when they began to vary.  Gsoil.b.10cm sage began to plummet a little after 1100 the is continuing to drop (at 1400 it was down to -10)

Soil Thermal Properties: OK
Steve has been keeping a
close eye on the data and the sensors appear to working properly, and
are giving us some good data.

T Profiles: OK

RH profiles:  OK

u,v spectra, stn 1-3:   OK.  Both spectra have some white noise at
frequencies
greater than 10.

kryton spectra:  OK, some white noise shown in 0000-0030 plot at frequencies greater than 1 Hz

CO2 cals and spectra:  CO2 variance below the noise level of the sensor. 



208: Status, Site all stns, Tue 11-Mar-2003 16:09:05 MST, status March 11
Weather Plots:  OK

Rain Gauge: OK

Vbatt: OK only got down to ~12.5

Sonics 2m: OK.  

Sonics profile: OK

Kryptons: OK, no diurnal variation. kh2o.20m.stn0 is measuring more moisture than other sensors (between .5-1 g/m^3 difference)

H2O-kh2o: OK

Tv and P:  OK

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  OK.

TRH_diff_profiler:  OK

Thermocouples, Ice Detectors:  OK

Long Wave Rad:. Rlw.out.grease stn0 is measuring less outgoing radiation during the day than other sensors. 

|Tdome-Tcase| : .OK

Tsoil:  . Slight variation between the data, but differences aren't 
too great.  Temperature range is between 0 and -2.0 degC)

Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil:  OK. Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
These all look a little strange to me, however that is probably because
Gsoil.c.10cmsage is not messing up the scale. Fans, I assume, are what
cause the rippling of the
traces.

Soil Thermal Properties: OK
Steve has been keeping a
close eye on the data and the sensors appear to working properly, and
are giving us some good data.

T Profiles: OK

RH profiles:  OK

u,v spectra, stn 1-3:   OK.  Both spectra have some white noise at
frequencies
greater than 10.

w,tc spectra, stn 1-3:  tc.2m 1200-1230 spectra from the sage site
varies from the other specta traces at the higher frequencies.

u:v, w:tc cospectra: Might want to look at lake(1200-1230).
Measurements are lower than what are coming out of sage and tower


u,v spectra tower: Are all in good agreement. Some white noise  at
frequencies greater than 10 Hz.

w,tc spectra tower:  for 1200-1230 plot signals vary quite a bit from one another.  There is also some white noise at frequencies greater than 10 Hz especially for tc30m, tc20m, tc15m, and tc10m. 

w,tc cospectra: 1200-1230 plot, there is some variation between the traces. w.1mtc.1m and w.2m.tc.2m are measuring higher.

kryton spectra:  OK, some white noise from kh2o.2m.

CO2 cals and spectra:  CO2 variance below the noise level of the sensor. 


207: Status, Site all stns, Mon 10-Mar-2003 10:24:45 MST, status March 7
Weather Plots:  OK

Rain Gauge: Looks different.  The green spikes don't reach 2.0mm like
usual, they are much smaller ((~.75)

Vbatt: OK. Only got down to ~12.8

Sonics 2m: OK.  w.15m stn0  measures  less that other sensors by ~.2-.6
off and on.


Sonics profile: OK

Kryptons: No diurnal variation but both kh20.2m.lake and kh20.2m.sage
had data spikes between 4-5 g/m3.

H2O-kh2o: OK

Tv and P:  OK

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  OK.

TRH_diff_profiler:  OK

Thermocouples, Ice Detectors:  Ttc.h stn0  varies a lot from other
sensors. May be exposed to the sun???
Ice detector data has a spike from 0 to .7 at 1400 MST 3/6..
Ttc.ref.stn0 is a reference temperature and will vary from the other
Ttc's.

Long Wave Rad:. OK

|Tdome-Tcase| : .OK

Tsoil:  . Slight variation between the data, but differences don't seem
too great.

Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil:  Gsoil.a.10cm lake still looks like it is going bad.  It's trace
looks nothing like what the other sensors are measuring.

Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
These all look a little strange to me, however that is probably because
Gsoil.c.10cmsage is not messing up the scale. Fans, I assume, are what
cause the rippling of the
traces.

Soil Thermal Properties: OK
Steve has been keeping a
close eye on the data and the sensors appear to working properly, and
are giving us some good data.

T Profiles: OK

RH profiles:  OK

u,v spectra, stn 1-3:   OK.  Both spectra have some white noise at
frequencies
greater than 10.

w,tc spectra, stn 1-3:  tc.2m 1200-1230 spectra from the sage site
varies from the other data collected (why between 1200-1230 and not
0000-0030?).

u:v, w:tc cospectra: Might want to look at sage instrument (1200-1230).
Sage measurements are different that what's coming out of lake and tower
(which are in good agreement).

u,v spectra tower: Are all in good agreement. Some white noise  at
frequencies greater than 10 Hz.

w,tc spectra tower:  Are all in good agreement.  Quite a bit of white
noise  from tc spectra at frequencies greater than 1.0 Hz at 1200-1230.

kryton spectra:  OK, some white noise from kh2o.2m.

CO2 cals and spectra:  Spectra doesn't look very good.  It's doesn't
follow the line at all.

206: Status, Site all stns, Mon 10-Mar-2003 10:24:10 MST, status March 6
Weather Plots:  OK

Rain Gauge: OK

Vbatt: OK. Only got down to ~12.8

Sonics 2m: OK.  w.15m stn0  measures  less that other sensors by ~.2-.6
off and on.


Sonics profile: OK

Kryptons: No diurnal variation but both kh20.2m.tower and kh20.2m.sage
had data spikes between 4-5 g/m3.

H2O-kh2o: OK

Tv and P:  OK

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: w'tc'.1m stn0 had two data spikes (~.10 m/s degC) while
other fluxes were around -.10.

CO2:  OK

TRHdiff 2m:  OK.

TRH_diff_profiler:  OK

Thermocouples, Ice Detectors:  Ttc.h stn0  varies a lot from other
sensors. May be exposed to the sun???
Ice detector data has a spike from 0 to 1.5 at 1000 MST.
Ttc.ref.stn0 is a reference temperature and will vary from the other
Ttc's.

Long Wave Rad:. Rlw.out.bare stn0, between 2000 MST 3/5 and 1000 MST
3/6, is way out of sync with other sensors  with differences as great as
40 W/m2.

|Tdome-Tcase| : .OK

Tsoil:  . Tsoil.a.5cm.sage and Tsoil.a.10cm sage are measuring slightly
colder soil temperatures (~4.5-5 decC) than other sensors (3-3.5 devC).

Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil:  Gsoil.a.10cm lake looks like it is going bad.  It's trace looks
nothing like what the other sensors are measuring.

Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
These all look a little strange to me, however that is probably because
Gsoil.c.10cmsage is not messing up the scale. Fans, I assume, are what
cause the rippling of the
traces.

Soil Thermal Properties: Data from Vpile.5cm should be looked at more
closely.  Data in third plot doesn't look right.
Steve has been keeping a
close eye on the data and the sensors appear to working properly, and
are giving us some good data.

T Profiles: OK

RH profiles:  OK

u,v spectra, stn 1-3:   More variation in the data at lower frequencies
than at high. Also, both spectra have some white noise at frequencies
greater than 10.

w,tc spectra, stn 1-3:  tc.2m 1200-1230 spectra from the sage site
varies from the other data collected (why between 1200-1230 and not
000-0030?).

u:v, w:tc cospectra: Might want to look at sage instrument. Sage
measurements are different that what's coming out of lake and tower
(which are in good agreement).

u,v spectra tower: Are all in good agreement. Some white noise  at
frequencies greater than 10 Hz.

w,tc spectra tower:  Are all in good agreement.  Quite a bit of white
noise  from tc spectra at frequencies greater than 1.0 Hz

kryton spectra:  OK, some white noise from kh2o.2m.

CO2 cals and spectra:  Spectra doesn't look very good.  It's doesn't
follow the line at all.

205: Status, Site all stns, Mon 10-Mar-2003 10:23:38 MST, status March 5
Weather Plots:  OK

Rain Gauge: OK

Vbatt: OK. Got down to ~12.5

Sonics 2m: w.15m stn0 is measuring less that other sensors by ~.2-.6
between after 2200 MST 3/4.

Sonics profile: OK

Kryptons: Fog at night results in diurnal variation again.

H2O-kh2o: OK

Tv and P:  OK

Scalar variance: OK

Fluxes at 2m: OK. ( Sage flux is always just a bit higher than fluxes at
other sites, what is the reason for this?)

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  Spike in data at both sage and lake sites of ~2.5 deg at
1000 MST.  RH also spiked at that time from 2% to 15%.

TRH_diff_profiler:  Spike of temp data noticeable here also.

Thermocouples, Ice Detectors:  Ttc? (light blue line...can't tell what
stn it is) varies from other sensors by ~5deg.   Ttc.ref.stn0 is a
reference
temperature and will vary from the other Ttc's.

Long Wave Rad:. OK

|Tdome-Tcase| : .OK

Tsoil:  . Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil:  Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
These all look a little strange to me, however that is probably because
Gsoil.c.10cmsage is not messing up the scale. Fans, I assume, are what
cause the rippling of the traces.

Soil Thermal Properties:  Are looking better.  Steve has been keeping a
close eye on the data and the sensors appear to working properly, and
are giving us some good data.

T Profiles: OK, second T sensor (1 meter) seems to be in line now.

RH profiles:  OK

u,v spectra, stn 1-3:   More variation in the data at lower frequencies
than at high. Also, v spectra has some white noise at frequencies
greater than 10.

w,tc spectra, stn 1-3:  tc.2m 1200-1230 spectra from the sage site
varies from the other data collected. Trace doesn't match data from lake
and tower at frequencies greater than .010 Hz.

u:v, w:tc cospectra: Might want to look at sage instruments. Sage
measurements are different that what's coming out of lake and tower
(which are in good agreement).

u,v spectra tower: Some white noise  at frequencies greater than 10 Hz.

w,tc spectra tower:  Quite a bit of white noise  from tc spectra
(1200-1230) at frequencies greater than 1.0 Hz

kryton spectra:  OK, some white noise from kh2o.2m.

CO2 cals and spectra:  Spectra doesn't look very good.  It's doesn't
follow the line at all.

204: Status, Site all stns, Mon 10-Mar-2003 10:22:22 MST, status March 4
Weather Plots:  OK

Rain Gauge: OK

Vbatt: OK. Got down to ~12.5

Sonics 2m: w.15m stn0 is measuring less that other sensors by ~.2-.6
between 1100 MST on 3/2 and 0800 MST on 3/4.

Sonics profile: OK

Kryptons: Fog at night results in diurnal variation again.

H2O-kh2o: OK

Tv and P:  Sharp drop in T at ~0800 MSt from -5 to -7 deg for all
sensors.

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  Spike in data at both sage and lake sites of ~2.5 deg at
1000 MST.  RH also spiked at that time from 2% to 15%.

TRH_diff_profiler:  Spike of temp data noticeable here also.

Thermocouples, Ice Detectors:  OK. Spike in ice detector data was a
result of Steve's hand over the sensor.  Ttc.ref.stn0 is a reference
temperature and will vary from the other Ttc's.

Long Wave Rad:. OK

|Tdome-Tcase| : .OK

Tsoil:  . Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil:  Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
Cannot see Gsoilc.10cm.lakem either it's in line with other sensors, or
not making measurements.

Soil Thermal Properties:  Are looking better.  Steve has been keeping a
close eye on the data and the sensors appear to working properly, and
are giving us some good data.

T Profiles: Second sensor up from surface is reading too warm.  Steve
checked for ice build up, or evidence of a problem but found none.  It
is assumed that temperature difference is the result of a bias.

RH profiles:  Second sensor up is drier than others by approx. 20%
around 1000,  looks fine otherwise.

u,v spectra, stn 1-3:   v spectra has some white noise at frequencies
greater than 10.

w,tc spectra, stn 1-3:  tc.2m at 0000-0030 had a data spike just above
10 Hz.
tc.2m 1200-1230 spectra vary from one another for all sites at
frequencies greater than .010 Hz.

u,v spectra tower:  at 1200-1230 there is a spike in the v2m data, and
some white noise at frequencies greater than 10 Hz.

w,tc spectra tower:  Some white noise at frequencies greater than 10 Hz

kryton spectra:  OK, some white noise (0000-0030) at frequencies greater
than
1.0.

CO2 cals and spectra:  Spectra doesn't look very good.  It's doesn't
follow the line at all.

203: Weather, Site , Thu 06-Mar-2003 16:11:32 MST, snowing
It actually was nice about 13:30, but about 15:00 got some light snow pellets.
Now it has changed back to moderate snow (blowing, of course...)

202: Kryptons, Site lake, Thu 06-Mar-2003 15:08:56 MST, cleaned krypton
Just cleaned krypton too.

201: Radiometers, Site lake, Thu 06-Mar-2003 15:08:21 MST, clean rads
Have just cleaned all rads at lake.  Again, net looked okay.

200: CO2, Site tower, Thu 06-Mar-2003 13:49:25 MST, CO2 cal gas amounts
I just remembered that I checked the CO2 cylinders at about 09:00 this morning.
Hi cal: 1170
Lo cal: 1100
These look like plenty to me.

Note that there is a lot of water/ice in the pump box -- now to the base of
the subpanel in the floor of the box.  At the moment, the front panel is
completely clogged with blown snow, so more snow won't get in, but eventually
the front panel will be open again.

Because it was mostly ice, I couldn't see a good way to get rid of the ice
that was there.  Also, I haven't thought of a good way to block the holes.



199: Kryptons, Site sage, Thu 06-Mar-2003 13:48:53 MST, krypton cleaned
While I was there, I cleaned the krypton as well at sage.

198: Radiometers, Site sage, Thu 06-Mar-2003 13:46:46 MST, rads cleaned
Just cleaned all rads at sage about 12:15.  Again, I couldn't get the net 
apart, so I just cleaned it.  There are a few small (few mm) scratches on the 
upward dome, and it looks a bit milky-grey, but this looks normal to me.  It 
was pretty level, though I had to adjust it a bit after cleaning.


197: Kryptons, Site tower, Thu 06-Mar-2003 11:32:18 MST, cleaned krypton at 2m
Generally, the kryptons have looked okay, so I didn't think it was a high
priority to clean all of them -- especially since it is windy.  This done
~09:30??

196: Radiometers, Site tower, Thu 06-Mar-2003 11:31:01 MST, rad cleaned
As one of the 3 tasks at the tower this AM, I cleaned all rads (incl 25&4m).
I was going to replace the net rad dome at grass, but couldn't get it off.
I just cleaned and releveled it instead.  This done ~09:00-09:30.



195: Weather, Site , Thu 06-Mar-2003 11:05:06 MST, still windy with blowing snow
There have been lulls, but for the last 26 hours or so there has been a lot of
wind moving the snow pack around.  Some spots are clear, others are ~0.5m.

194: Weather, Site , Wed 05-Mar-2003 17:10:12 MST, blowing snow
It has been blowing like crazy most of the day (8+ m/s at 2m) and snow has
been drifting all over.  However, there was no snow, except for a possible
dusting early this morning (pre 05:30).  This is a good case if anyone wants
to look at drifting snow (to check the rain gauge??).

193: Log, Site , Wed 05-Mar-2003 07:33:37 MST, KingAir flew this AM
flight from only ~6:00-7:30 with Larry guiding.

192: Radiometers, Site all stns, Wed 05-Mar-2003 07:29:19 MST, some clouds this AM
When Kate was here on Monday, I was concerned that Rsw.in was significantly
different at times between the sites.  Some of this could be snow/ice on the
domes, but it seemed like there were more cases with different radiation than
could be explained just by snow/ice.  Now is an example where there are patches
of sun slowly propogating across the surface, which could make differences
(at least instantaneously) of Rsw.in between lake/sage/tower.  I'm still not
completely convinced that the 5-min ave values should be as different as the
400 W/m2 that we sometimes get.

P.S. The attached graph shows several problems (that are real!).  [The red line
is: 1300*sin(dat("azel.sun")[,2]*pi/180)
1. A semi-stationary cloud pattern (probably orographically influenced) shadows
grass and lake at about noon, but doesn't move over sage until 45min later.
2. When this cloud shifts off at about 14:00, it apparently ducts some of the
sun to have even more than a clear-sky irradiance.
3. At about 16:00, thick clouds built up along the Gore Range cause "sunset" to
occur early, though some light is still present until real sunset.

It is going to be difficult to use these data for QCing snow/ice cases!
It may also indicate that using Rlw.in.sage as a surrogate for Rlw.in.grass may
be a mistake.

191: Temp_RH, Site tower, Tue 04-Mar-2003 18:35:51 MST, swapped 0.5-1.0 TRH
In our ongoing effort to nail down biases in the T/RH profiles, I've swapped
0.5 and 1.0 (recall that this morning I put a "new" sensor in at 1.0, so now
the new one is at 0.5.  Immediately, the RH profile at least developed a +2%
large value at 1.0.  I'll have to get more data to see how the T profile 
changed.  In any case, I think the old 0.5m (new 1.0m) T/RH has some biases.
(Steve's post-cals of these sensors will be critical!)

190: Log, Site lake, Tue 04-Mar-2003 15:14:57 MST, Lake site survey
For reference, I've surveyed the stuff at lake:
All positions are referenced to the center of the PAM tripod:
				Bearing		Range (m)
Center of Power tripod		36		12.4
Center of rad stand		102		11.10
Soil Mux box			40		4.20
Soil site A stake		91		4.55
Soil site B stake		43		6.00
Soil site C stake		couldn't locate (didn't try!)

189: Soil, Site all stns, Tue 04-Mar-2003 13:37:58 MST, Gravimetric soil samples
17:00 Mar 4 Sage
Near B.
                   Fresh(4 Mar) Dry(5 Mar)	%(wet/total)
A sample taken 16:00 which broke, but I recovered some:
   0-~5	   	   80.3		56.9		29.1
		   56.9		49.7		12.7

The real sample:
   0-3 cm  EI 	   28.0	    	18.4 		34.3
           EII     30.0		19.8		34.0
           EIII    28.7		19.8		31.0
   Mean:					33.1
   Total sample:   88.6         [59.3]
   Density:			[0.86]

   3-8 cm  FI      52.8		44.1		16.5
           FII     54.0		44.5		17.6
           FIII    48.8		38.4		21.3
   Mean:					18.5
   Total sample:  179.3		[146.1]
   Density:			[1.28]

12:45 Mar 4 Lake
Close to B!
Dug down through 28 cm of snow - broke up ice coating surface

                   Fresh(4 Mar) Dry(5 Mar)
(has some ice/snow on the top, but mostly solid frozen)
  0-3 cm  AI       38.0		28.5		25.0
          AII      35.9		27.3		24.0
          AIII     23.3		17.8		23.6
  Mean:						24.2
  Total sample:	   99.8		[75.6]
  Density			[1.10]

(peaty: root and stem)
  3-8 cm  BI       19.3		14.7		23.8
          BII      15.7		11.3		28.0
          BIII     16.0		11.2		30.0
  Mean:						27.3	
  Total sample:    87.8         [63.8]
  Density:	   		[0.56]

15:00 Mar 5 Grass site:
Close to B:
Dug down through 28 cm of snow - broke up ice coating surface

                   Fresh(5 Mar) Dry(6 Mar)
(has some ice/snow on the top, but mostly solid frozen)
  0-3 cm  CI       18.3		4.4		76.0
          CII      10.9		3.1		71.6
          CIII     12.9		2.9		77.5
  Mean:						[75.0]
  Total sample:	   42.9		[10.7]
  Density	   0.62		0.16

(peaty: root and stem)
  3-8 cm  DI       28.3		22.1		21.9
          DII      19.8		13.1		33.8
          DIII     29.4		24.0		18.4
  Mean:						[24.7]
  Total sample:    111.2	[83.7]
  Density:	   0.97		0.73

17:00 Mar 5 Lake site:

Close to C:

                   Fresh(5 Mar) Dry(
(has some ice/snow on the top, but mostly solid frozen)
  0-3 cm  EI       32.8		23.5		28.4	[had a grub inside!]
          EII      26.9		19.7		26.8
          EIII     28.4		20.6		27.5
  Mean:						[27.5]
  Total sample:	   92.5		[67.1]
  Density	   1.35		0.98

(peaty: root and stem)
  3-8 cm  FI       39.4		24.7		37.3
          FII      44.8		32.3		27.9
          FIII     39.8		22.7		43.0
  Mean:						[36.1]
  Total sample:    157.9	[100.9]
  Density:	   1.38		0.88

09:30 Mar 6 Grass site:
Close to A:
                   Fresh(6 Mar) Dry(7 Mar)
(has some ice/snow on the top, but mostly solid frozen)
  0-3 cm  AI       14.5		1.6		89.0
          AII      10.6		1.0		90.6
          AIII     11.1		1.1		90.1
						89.8
  Mean:			
  Total sample:	   37.0 	3.8		(but not complete volume)
  Density	   0.54		0.06

(peaty: root and stem)
  3-8 cm  BI       22.2		9.8		55.9
          BII      16.5		7.5		54.5
          BIII     17.9		4.6		74.3
  Mean:						61.3
  Total sample:    87.8		34.0
  Density:	   0.77		0.30

13:00 Mar 6 Sage site:
Close to A:
I took two cores that broke (along with the corer), so I only recovered about 
6cm deep cores twice.
                   Fresh(6 Mar) Dry(7 Mar)
  0-~6 cm EI       44.2		34.3		22.4
          EII      34.5		27.3		20.9
          EIII     42.4		34.0		19.8
  Mean:						21.1
  Total sample:	   138.7 	109.4		(very crude estimate)
  Density	   1.01		0.80

  0-~6 cm FI       42.8		34.5		19.4
          FII      37.7		31.4		16.7
          FIII     39.4		31.8		19.3
  Mean:						18.5
  Total sample:    168.1 	137.0		(very crude estimate)
  Density:	   1.22		1.00


[] values are calculated from other variables
Diameter of corer is 5.4cm













188: Temp_RH, Site tower, Tue 04-Mar-2003 10:14:55 MST, T/RH change
In an attempt to understand temperature profile (and RH) biases, I've just
put spare T/RHs in at 2m and 1m:
2m: 004 -> 104
1m: 703 -> 001
0.5m still appears to be lower than 1m, which isn't expected.  Thus, I'll 
probably do some more swapping soon.


187: Weather, Site all stns, Tue 04-Mar-2003 09:00:17 MST, snowing
It started snowing ~0500 this morning and is still coming down -- I estimate
about 1" has fallen so far.  About 15 min ago, I checked sage and all sensors
were snow-free (T/RH, sonic, krypton, 4-comp rad, and net).

186: Status, Site , Sun 02-Mar-2003 11:55:33 MST, Daily status, feb 28
Weather Plots:  OK

Rain Gauge: OK

Vbatt: OK. Got down to ~12.3

Sonics 2m: OK

Sonics profile: OK

Kryptons: OK. Showing diurnal variation again.  kh2o.20m had the highest
reading over night at 20 g/m3.

H2O-kh2o: OK

Tv and P:  OK.

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2:  Tony and Steve had done some manual calibrations yesterday which
caused variations in the data (see earlier log entry for more detail).
Look OK today.

TRHdiff 2m:  OK

TRH_diff_profiler:  OK.

Thermocouples, Ice Detectors:  Ttc.h.stn0 varies from the other sensors
quite a bit.  Ttc.ref.stn0 is OK. It is a reference
temperature and will vary from the other Ttc's.

Long Wave Rad:. OK

|Tdome-Tcase| : .OK

Tsoil:  . Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil:  Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
Gsoilc.10cm.lake still bad.

Soil Thermal Properties:  Are looking better.  Steve has been keeping a
close eye on the data and the sensors appear to working properly, and
are giving us some good data.

u,v spectra, stn 1-3:   OK

w,tc spectra, stn 1-3:  tc.2m between 1200-1230 is slightly lower than
lake and sage.

u,v spectra tower:  OK.

w,tc spectra tower:  OK

kryton spectra:  OK, some white noise at frequencies greater than
0.1-1.0

CO2 cals and spectra:  some white noise at frequencies greater than 1.0.


185: Status, Site , Sun 02-Mar-2003 10:47:54 MST, Daily status, Feb 27
Weather Plots:  OK

Rain Gauge: OK

Vbatt: Got pretty low last night.  It dropped to ~12.1.

Sonics 2m: OK

Sonics profile: OK

Kryptons: OK. Showing diurnal variation again.

H2O-kh2o: OK

Tv and P:  OK.

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2:  Between 1000 and 1200 MST it jumped from 248 to 254, then dropped
to 237, jumped back up to 250 and then began behaving.

TRHdiff 2m:  at 1200 MST lake RH increased to almost 15 % while sage was
recording ~4% RH, other than that they are in good agreement.

TRH_diff_profiler:  OK. 2m tower RH sensor seems to be OK.  5m - .5m
trace dropped to -28% at ~1430 MST Feb 26, and then increased slowly
over next few hours.

Thermocouples, Ice Detectors:  Ttc.h.stn0 varies from the other sensors
quite a bit.  Ttc.ref.stn0 is OK. It is a reference
temperature and will vary from the other Ttc's.

Long Wave Rad:. OK

|Tdome-Tcase| : .OK

Tsoil:  . Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil:  Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
Gsoilc.10cm.lake still bad.

Soil Thermal Properties:  Are looking better.  Steve has been keeping a
close eye on the data and the sensors appear to working properly, and
are giving us some good data.

u,v spectra, stn 1-3:   No plots as of 4:00.

w,tc spectra, stn 1-3:  No plots as of 4:00.

u,v spectra tower:  No plots as of 4:00.

w,tc spectra tower:  No plots as of 4:00.

kryton spectra:  No plots as of 4:00.

CO2 cals and spectra:  No plots as of 4:00.


184: Status, Site , Sun 02-Mar-2003 10:46:47 MST, Daily status, Feb 26
Weather Plots:  OK

Rain Gauge: OK

Vbatt: Got pretty low last night.  It dropped to ~12.1.

Sonics 2m: OK

Sonics profile: OK

Kryptons: OK. Showing diurnal variation again.

H2O-kh2o: OK

Tv and P:  OK.

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  OK

TRH_diff_profiler:  Temp profile is good.  RH profile looks a little
different most likely because of the snow.  2m meter sensor is still
reading ~3-4% lower than other sensors, may be malfunctioning.

Thermocouples, Ice Detectors:  Ttc.h.stn0 varies from the other sensors
quite a bit, and lacks the features present in the other traces??
Ttc.ref.stn0 is OK. It is a reference
temperature and will vary from the other Ttc's.

Long Wave Rad:. OK

|Tdome-Tcase| : .OK

Tsoil:  . Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil:  Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
Gsoilc.10cm.lake still on the fritz.

Soil Thermal Properties:  Are looking better.  Steve has been keeping a
close eye on the data and the sensors appear to working properly, and
are giving us some good data.

u,v spectra, stn 1-3:   Good

w,tc spectra, stn 1-3:  OK

u,v spectra tower:  OK

w,tc spectra tower: OK

kryton spectra:  A lot of white noise between 0000-0030 higher
than 1.00 Hz.  Seems to keep increasing every day...might want to check?

CO2 cals and spectra:  Missing data that was there yesterday?

183: Log, Site , Sun 02-Mar-2003 10:10:52 MST, king air flew today
Supposedly, the King-Air flew this morning, taking off at 6AM, though there
was no sign of it by the time we arrived on site at 0830.  (We went to 
Moose Creek for our pancake/french toast breakfast.)

182: Weather, Site all stns, Sun 02-Mar-2003 09:58:56 MST, fog/frost
There was heavy fog last night and associated rime this morning on all sensors.
On the drive in (~0830), I brushed the rime (but couldn't get at all of the
underlying frost) off of all rads, sonic, and krypton at sage, then at lake.

About 0900, I brushed off bare.1.5m, then all grass rads.  I then brushed off
the 1&2m sonic and krypton at the tower.  Next, I "twanged" the NW guy wires
to knock rime off all sonics and kryptons.  Nevertheless, I decided to climb
the tower to check on the effectiveness of the "twanging".  I twanged the
side boom guy wires and was able to knock off a bit more rime, though I think
that the tower guy twanging would have been adequate by itself.  All this
twanging was sufficient to get the 20&30m sonics happy again (not cycling diag).

While on the tower, I noticed that rime had clogged the T/RH inlets.  I blew/
wiped the rime off of all of the T/RHs, in the order: 30, 25, 20, 15, 10, 5,
0.5, 1, 2m.  Note that I didn't open the tubes to get at any ice that may
be on the Gortex cap.  Also note that I didn't think of doing the T/RHs at
lake and sage.

181: Temp_RH, Site tower, Sat 01-Mar-2003 16:06:34 MST, analysis of T.5m fan
I'm starting to look at the T/RH aspiration problems.  During the last part
of Dec and early Jan, T.5m sometimes looks okay and sometimes bad as its
fan became more or less effective.  The attached plot shows Rsw.in on the top,
spd on the bottom, and a flag, ib = abs(T.5m-T.2m)>abs(T.10m-T.2m).  ib should
be 0.  On every day during this period, ib was 1 sometime during the daytime
and 1 during the whole daytime period when the wind was calm.

There may be good data during some of this time period, but we'll never be sure.

180: Weather, Site all stns, Sat 01-Mar-2003 14:37:43 MST, fog this AM
We took the morning off to go to Steamboat, but noticed that there was still
fog in low-lying areas at 8:40 this morning.  Thus, I'm not surprised that the
kryptons had their usual failures.

We also saw 2 coyotes and a large herd of elk along the road driving into the
site just now.

179: Log, Site , Fri 28-Feb-2003 16:27:46 MST, analysis thoughts
I've spent the afternoon looking at tc and kh2o data over the last 100 days.
Some prelim. conclusions:
1. There are significant differences between tc and T at certain times.
Many of these probably are due to bad fans on the T/RHs!
2. When the kryptons work, they generally are within 1 g/m3 of Q (except
30m which is up to 1.5 g/m3).  This is quite good, considering that we've
only cleaned every month or so.
3. Using the difference kh2o-Q as a QC flag for kh2o is the best start.
4. Using a threshold on kh2oV is a mess and misses "edges".
5. Using a threshold on skewness or kurtosis doesn't remove many points (<<1%)
6. There are events in tc that cause large variance which are not "internal
instrumental" problems since they are seen to a lesser extent in both T and 
kh2o.  Nevertheless, they many still be exposure artifacts rather than
atmospheric phenomena. [They look similar to rain on a thermocouple or in
my sonic sprinkler tests.)

178: GPS, Site sage, Fri 28-Feb-2003 09:31:18 MST, GPS testing
At John's advice, I'm logging the GPS signal using EVE-talk.  Thus, data from 
sage already has had one outage and I will make another outage now.  (Since
I note that the GPS always goes bad at 39 minutes past the hour.) 
177: Sonics, Site tower, Fri 28-Feb-2003 09:13:59 MST, weird 1m sonic spectra
While looking over plots last night, Dean and I noticed that the 1m sonic
has a velocity spectra that are significantly different than the other sonics
on the tower -- with more energy and shifted to higher frequencies, especially
during the day. This has been going on for as long as we've generated spectra.

This may be due simply to the lower height, or it may indicate turbulence 
generation from the tower or other local clutter.  I'll try to produce a 
normalized spectral plot to see if it is just a height issue.

P.S. I've generated and attached the plot.  Everything looks great (I only used
nominal heights -- hopefully it would look even better with actual sensor 
heights.)  I wonder if the WWW plots are being produced correctly?  I'll 
check that now.

P.P.S. Yes, the WWW plots are okay.  It is just an optical illusion that the
1m is different when the 2m data aren't on this plot.  I may add it...



176: Weather, Site , Fri 28-Feb-2003 08:43:22 MST, low clouds just breaking up
I notice that all kryptons died again last night.  The radio had reports of fog,
though the site just had low clouds this morning, i.e. slightly elevated fog.
This fog is just now breaking up.

175: GPS, Site sage, Thu 27-Feb-2003 16:20:12 MST, Intermittent GPS outages at sage
Today we've had several "Lost Data" messages from the GPS at sage.  A plot of
the raw data indicate that we lost at least one sample from this sensor every
hour, and John's config is only reporting a sample every minute.  (In contrast,
lake's GPS has only lost one sample in the last 8 hours.)

John speculates that this GPS is having problems locking onto the satellites,
is reporting that its data may be bad, and EVE is rejecting these messages.

I just went to sage for an inspection of the GPS.   All looked okay, but I
cycled power on the GPS and rotated the antenna 90 degrees, so that the walk
there wasn't wasted.  We'll have to see what happens.

(John has suggested using EVE_talk to this sensor to look for "A" or "Z?" as
the first character, but I think this would be hard for such an intermittent
problem.)

174: Sonics, Site tower, Thu 27-Feb-2003 16:16:30 MST, USFS station further South
Instead of being near the SE tower guy points, Bill Massman's station is more
to the south of our tower.  Using the datascope, I estimate that his station
is between 152-160 degrees from our sonics.  A wind direction histogram
indicates that this range of directions occurs only 2% of the time (and all
directions occur at least 1% of the time), so we will let him keep this
station where it is.

173: Temp_RH, Site tower, Thu 27-Feb-2003 11:49:18 MST, brushed snow caps from 15-30m
From ~9:45-10:30, removed inlet tubes and then brushed snow caps off the Gortex
shields from T/RHs at 30, 25, 20, and 15m.

172: CO2, Site tower, Thu 27-Feb-2003 11:26:23 MST, CO2 cal flows and supplimentary cals
10:35 CO2 intake flow 6.75 lpm
      Hi-cal xs flow 2.2 lpm
      Lo-cal xs flow 2.2 lpm
There was no need to adjust the bypass nor secondary regulator pressures.

Hi-cal 1220/10    Lo-cal 1310/12

Arranged plumbing for suplimentary calibrations

10:40 to 10:50 N spare 1650 psi xs flow 2 lpm
10:55 to 11:05 S spare 1760 psi xs flow adjusted to 2 lpm but wouldnt stay 
                                   const. had to cont incr pressure at regulator
171: Radiometers, Site tower, Thu 27-Feb-2003 11:24:25 MST, Boom trolley
Last evening after fixing the fans on the 25m pyrgs the trolley was moved out.
Unfortunately it wasn't moved to the end of the boom
At 10:15 the trolley was moved to the end. Photo's to prove it
170: Log, Site sage, Thu 27-Feb-2003 08:26:33 MST, station visit
Since there was patchy fog this morning, we decided to stop by sage on our way
in about 15min ago and photograph the krypton -- there was feather rime on it,
the sonic, and all rads (especially the net, which actually has snow on the
uplooking dome).  See img_2669-2672.jpg.

While there, we also did the same T/RH cleaning off of a 4mm thick crust of
snow which has packed onto the tip of the gortex filter and removed the snow
from the "bell".

We didn't clean the ice off the radiometer domes.

169: Temp_RH, Site tower, Wed 26-Feb-2003 15:37:08 MST, T/RH's on tower
14:40 5m T/RH replaced, 204 removed, ??? installed
15:00 10 m disassembled and an icicle removed from the teflon filter. 
            Reinstalled 10 m
16:25 - 16:30 Removed icicles from teflon filters of 0.5m, 1m, and 2m T/RH

168: Radiometers, Site all stns, Wed 26-Feb-2003 15:28:33 MST, Removing radiometer snow doughnuts
The ventilation prevented the fine powdery snow from settling upon the 
radiometer domes. Doughnuts of snow formed, up to 7-8cm high, with a central
hole and an annulus extending from the ventilator slot to the edge of the
radiation shield. All upward-looking radiometers as well as the ice detectors
had the same problem. Some downlooking rads had an icicle from their dome.
14:00 Sage, removed doughnuts and cleaned all domes.NB Netrad was already clear.
14:10 Lake, ditto
14:21 Grease, cleaned dome
14:28 Bare, cleaned dome.
14:30 4m scaffold, removed doughnut and cleaned domes
14:32 Ice.a and Ice.b ditto
14:45 - 14:55 Drew in boom. S pyrg already clear, removed doughnut from N pyrg.
              The fans were unplugged with the N fan's wires cut.
              Left trolley inboard until we fix fan wiring.
16:10 -16:20  Remounted fans and powered up. Fans operated. 
              Ran out the trolley BUT THE TROLLEY DIDN'T GO ALL THE WAY



167: Status, Site all stns, Tue 25-Feb-2003 16:35:00 MST, Daily status, Feb 25
Weather Plots:  OK

Rain Gauge: OK

Vbatt: Got a little low last night.  It dropped to ~12.1.

Sonics 2m: OK

Sonics profile: OK

Krytons: kh2o.20m  stn0  spiked to almost 14 g/m3 between 1000 and 1100

MST Feb  Other sensors ranged between 2-4 g/m3 during that same time.
No diurnal variation last night.

H2O-kh2o: this plot also shows data spike from 20m sensor.

Tv and P:  OK.

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  OK

TRH_diff_profiler:  Temp profile is good.  RH profile looks a little
different most likely because of the snow.  30m sensor is recording
about 2% more moisture then other sensors, which is probably fine.
Should keep an eye on 2m meter sensor to see how it behaves.  Currently
it's reading ~3-4% lower than other sensors, may be malfunctioning.

Thermocouples, Ice Detectors:  Ttc.ref.stn0 is OK. It is a reference
temperature and will vary from the other Ttc's.

Long Wave Rad:. OK

|Tdome-Tcase| : .OK

Tsoil:  . Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil:  Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
Gsoilc.10cm.lake still on the fritz.

Soil Thermal Properties:  Are looking better.  Steve has been keeping a
close eye on the data and the sensors appear to working properly, and
are giving us some good data.

u,v spectra, stn 1-3:   Good

w,tc spectra, stn 1-3:  OK

u,v spectra tower:  OK

w,tc spectra tower: OK

kryton spectra:  A lot of white noise between 0000-0030 higher
than 1.00 Hz

166: , Site all stns, Tue 25-Feb-2003 10:59:37 MST, Site inspection after snow
Persistant snow fall last night and on into the daytime. Only ~10 cm of low 
density powder.
Saw well over 100 elk near the road near Sage..
Arrived at 8:45: check_aster, Daisy and Cosmos, Dailies only completed to 22 Feb.
9:30 - 10:00 Yutaka Hagimoto and I inspected tower site.
Ice detecors have snow doughnuts, 6 or so cm high with a hole down to the domes.
Photograped and also applied handkerchief test. Light lit (and also later saw
cockpit response.
Cleaned snow from cal gas guages:
Hi-cal = 1250/10
Lo-cal = 1400/12
Walked in front of tower photographed lower levels of sensors. Neither sonics 
nor Kr's snow clogged..
Climbed 20m or so up tower. Snow on steps and walkways.

10:30 visited Sage: cleaned solar panels
From ladder photographed LW.in and SW.in. Same snow doughnuts as ice detectors.
Neither sonic nor Kr snow clogged.
165: Status, Site all stns, Mon 24-Feb-2003 17:08:00 MST, Daily status, Feb 24
Weather Plots:  OK

Rain Gauge: OK

Vbatt: Great.

Sonics 2m: OK

Sonics profile: OK

Krytons: kh2o.2m  sage increased to almost 6 g/m3 between 2100 and 2200
MST Feb 23 and then slowly dropped back down.  Other sensors ranged
between 1-2 g/m3 during that same time.

H2O-kh2o: this plot shows sensor at 2m sage measuring more moisture
also.

Tv and P:  OK.

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2:  OK

TRHdiff 2m:  RH plot shows sage-tower is about 2% more moist than
lake-tower between 1900 MST Feb 23 and 1300 MST Feb 24.

TRH_diff_profiler: looks like we had a little inversion between
0500-0800 MST

Thermocouples, Ice Detectors:  Ttc.ref.stn0 is doing it's own thing (and
has been since plotting began), and doesn't look good. Ttc.h.stn.0
deviates a bit from the other sensors also (~2-5 deg) between 2000 MST
Feb 23 and 0400 Feb 24.

Long Wave Rad:. OK

|Tdome-Tcase| : .OK

Tsoil:  . Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil:  Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
Gsoilc.10cm.lake still on the fritz.

Soil Thermal Properties:  Are looking better.  Steve has been keeping a
close eye on the data and the sensors appear to working properly, and
are giving us some good data.

u,v spectra, stn 1-3:   Good

w,tc spectra, stn 1-3:  OK

u,v spectra tower:  OK

w,tc spectra tower: w.1m isn't really following along with other
profiles?

kryton spectra:  Quite a bit of white noise between 0000-0030 higher
than 1.00 Hz

164: Soil, Site lake, Mon 24-Feb-2003 16:58:38 MST, Deep soil moisture at Lake
At Lake site at 15:00- 16:00. Level with the middle soil installations, 6 - 7m
north of the center of the radiometer stand.
Cleared the 15 cm snow for a 1m diameter circle. 
There was an ice crust a cm or so thick. This was broken up and discarded.
Then carefully excluding
loose snow dug a 20 cm wide, 35 cm deep pit using a digging bar. The loose dirt
was sccoped out by hand and thrown to the north, away from the rad stand.
Using trowel took samples of 0-5cm,5-10cm,10-15cm,15-20cm,20-25cm,and 25-30 cm 
Returned to base and weighed aliquots
Into oven 16:40 on 24 Feb''03
Out of oven for weighing 09:00 on 26 Feb''03

         Aliquots   Fresh       Dry      Different    Gravmoist

 0-5cm    A I       33.4        25.4      8.0        0.315   
          A II      36.1        27.5      8.6        0.313     30.4%
          AIII      30.6        23.8      6.8        0.286

 5-10cm   BI        28.4        23.7      4.7        0.198
          BII       26.5        22.2      4.3        0.194     20.1%
          BIII      27.5        22.7      4.8        0.211

 10-15cm  CI        26.4        23.0      3.4        0.148
          CII       29.1        25.4      3.7        0.146     14.9%
          CIII      25.8        22.4      3.4        0.152

 15-20cm  DI        24.0        20.7      3.3        0.159
          DII       23.0        19.7      3.3        0.168     16.4%
          DIII      22.0        18.9      3.1        0.164

 20-25cm  EI        28.3        23.8      4.5        0.189
          EII       26.2        22.0      4.2        0.191     18.9%
          EIII      24.1        20.3      3.8        0.187

 25-30cm  FI        30.6        25.1      5.5        0.219
          FII       24.7        20.3      4.4        0.217     21.8%
          FIII      24.1        19.8      4.3        0.217

There is a minimum soil moisture at ~ 15 cm. Note that below this depth the soil
changes from "sandy" to "clayey".


163: Status, Site all stns, Fri 21-Feb-2003 15:00:00 MST, Daily status, Feb 21
Weather Plots:  OK

Rain Gauge: OK

Vbatt: Great.

Sonics 2m: OK

Sonics profile: OK

Krytons: Good.

H2O-kh2o: OK

Tv and P:  OK.

Scalar variance: for 2m tc'tc' plot tower and sage show a bit more
scalar variance than lake during the daytime (1.0-1.5 deg).

Fluxes at 2m: tower and sage also show slightly more flux than at lake,
although difference is not as great (~.3)

Flux profiles: OK

CO2:  Large spike (from 241 up to 253 and back down) at 1400 MST on Feb
20.

TRH:OK

Long Wave Rad:. OK

|Tdome-Tcase| : .OK

Tsoil:  . Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil: Gsoila.10cm.lake(3) began to drift off on it's own at ~2100 MST
Jan 19.  It kept decreasing until 0700-0800 MST on the 20th when it
slowly began to rise again.  By noon it was back in line with the other
sensors.
Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
Gsoilc.10cm.lake still on the fritz.

Soil Thermal Properties:  Are looking better.  Steve has been keeping a
close eye on the data and the sensors appear to working properly, and
are giving us some good data.

u,v spectra, stn 1-3:   There appears to be some white noise at the
lower frequencies (.001 Hz) and also at frequencies greater than 10.0.

w,tc spectra, stn 1-3:  OK

u,v spectra tower:  Some white noise at frequencies greater than 10.0 Hz

w,tc spectra tower: OK

kryton spectra:  OK
162: Status, Site all stns, Wed 20-Feb-2003 15:00:00 MST, Daily status, Feb 20
Weather Plots:  OK

Rain Gauge: OK

Vbatt: Great.

Sonics 2m: OK

Sonics profile: OK

Krytons: Good.

H2O-kh2o: OK

Tv and P:  OK.

Scalar variance: for 2m tc'tc' plot tower and sage show a bit more
scalar variance than lake during the daytime (1.0-1.5 deg).

Fluxes at 2m: tower and sage also show slightly more flux than at lake,
although difference is not as great (~.2)

Flux profiles: OK

CO2:  OK

TRH:OK

Long Wave Rad:. rlw.out.grease.stn0 measurements are lower than other
sensors by between 10-40 W/m2.

|Tdome-Tcase| : .OK

Tsoil:  . Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil: Gsoila.10cm.lake(3) began to drift off on it's own at ~2100 MST
Jan 19.  It kept decreasing until 0700-0800 MST on the 20th when it
slowly began to rise again.  By noon it was back in line with the other
sensors.
Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
Gsoilc.10cm.lake still on the fritz.

Soil Thermal Properties:  Are looking better.  Steve has been keeping a
close eye on the data and the sensors appear to working properly, and
are giving us some good data.

u,v spectra, stn 1-3:   There appears to be some white noise at the
lower frequencies (.001 Hz).

w,tc spectra, stn 1-3:  OK

u,v spectra tower:  No plots as of 4:25 pm

w,tc spectra tower: No plots as of 4:25 pm

kryton spectra:  No plots as of 4:25 pm

161: Status, Site all stns, Wed 19-Feb-2003 15:00:00 MST, Daily status, Feb 19
Weather Plots:  OK

Rain Gauge: OK

Vbatt: Great.

Sonics 2m: OK

Sonics profile: OK

Krytons: Good. Some diurnal variation with four of the sensors.

H2O-kh2o: OK

Tv and P:  OK.

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2:  At 1200 MST CO2 jumped from ~234 ppmV to ~246 ppmV

TRH:OK

Long Wave Rad:. rlw.out.grease.stn0 measurements are lower than other
sensors by about 20 W/m2.

|Tdome-Tcase| : .OK

Tsoil:  . Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm.sage still bad, the cable was probably cut.  Needs
to be fixed.
Gsoilc.10cm.lake still on the fritz.

Soil Thermal Properties:  Are looking better.  Steve has been keeping a
close eye on the data and the sensors appear to working properly, and
are giving us some good data.

u,v spectra, stn 1-3:   There appears to be some white noise at the
lower frequencies (.001 Hz).

w,tc spectra, stn 1-3:  OK

u,v spectra tower:  OK

w,tc spectra tower: OK

kryton spectra:  kh2o.10m and kh2o.2m sort of trail off on their own
between 00:00-00:30

160: Power, Site tower, Sat 22-Feb-2003 17:14:54 MST, power changes
I put the ice detector breakout box on a separate 12V supply.  To plug this
supply in, I installed a power strip in the CO2 pump box and rerouted the
pump to plug into it.  The current status is:
transformer breaker 1 - unused
		2 - used for daisy&cosmos
		3 - going to Bill Massman's setup (using an ADAM power supply)
		4 - to the electric fence charger
This was done about 12:30

159: Soil, Site all stns, Sat 22-Feb-2003 17:12:02 MST, 10V regulators added
From 1300-1400 added regulators to the switched 12V signals from the CR10s at
all 3 sites (grass, then sage, then lake).  I'm not sure about some of the
electrical contacts -- the daily plots show that I did make a change -- 
hopefully, it was for the good!

158: Temp_RH, Site tower, Sat 22-Feb-2003 17:10:02 MST, TRH work - finding why 10m is a max
Dean and Larry have noted that the 10m T was a bit (~0.1C) high.  I changed the fan at about 12:45, but this didn't help.  Now, we've just swapped 5m and 10m
and now see that 5m is too high.  Thus, there is a bit of a bias with the sensor
which is now at 5m.  (And the one which was at 5 and now 10 looks a bit low!).

Since the fan I removed appears to have been working, I've put it back in the
cabinet in the trailer.



157: , Site tower, Fri 21-Feb-2003 14:06:37 MST, Tasks completed
Before leaving on Friday:
  Deep soil core weighings done
  CO2 flow checked
  Frost detector operation checked
  14:00 Hical 1360/10, local 1590/12
Tasks completed this last week:
  All radiometers cleaned including Boom rads
  All Kr's cleaned including tower Kr
  Deep soil samples done
  Supplimentary CO2 calibrations
156: Adams, Site tower, Fri 21-Feb-2003 14:04:08 MST, Cosmos ADAM fan replaced
At 13:45 the old Cosmos ADAM fan was definately not operating. It could be 
started by poking it with a screwdriver but would soon stop.
The spare fan was employed as a replacement. 
155: Weather, Site all stns, Fri 21-Feb-2003 08:13:55 MST, Weather conditions
A light snow occured in the early morning.
By daybreak there were only scattered clouds.
A good opportunity to check the functioning of the frost detector.

154: Snow_Precip, Site tower, Thu 20-Feb-2003 16:37:01 MST, Frost detectors installed
The Analog breakout box arrived by UPS.
Installed at 16:20.Thu, 20 Feb
Small dome 0/1
Big dome  2/3
ABOB to ADAM 2
Fans running, on handkerchief covering, the LED's light and the heaters come on.

Fri, 21 Feb''03, Gordon reconfigured the cosmos config so I could see the
Ice detector on the cockpit.At 10:50 I performed the handkerchief test, first
on the small dome detector , then on the big dome detector. LED's and heaters OK
On cockpit:
 Chan 108, little dome Ice.a went from 0 to 5 vdc, Chan109, stability,no change
 Chan 110, big   dome  Ice.c went from 0 to 5 vdc, Chan 111,stability,no change

The frost detectors are operational

153: CO2, Site tower, Thu 20-Feb-2003 15:02:24 MST, Supplimentary CO2 calibrations
A supplimentary calibration system was assembled. It consisted of:
 a regulator with a shut-off valve, 10 m of 1/4" tubing leading into a ~ 250 cm3
 Savillex three connector teflon jar. The other two tubes connected to a
 floating ball flow meter and to either a close-off or the Licor intake.

The Licor intake flow was tested = 6.5 lpm

The regulator was connected to a cylinder, the flow thro the jar with the 
shut-off installed was adjusted to 10 lpm and run for a few minutes to flush 
the system and then the shut-off was removed and the line connected to the Licor
intake. The excess flow of a few lpm was checked. The suplimentary cal gas
was fed to the Licor intake for ~ 10 minutes.
The system was disconnected and after set-up with the second cylinder the 
process was repeated.

First calgas K694978, ppm = 371.5 , Pressure 1700 psi, start 13:50, stop 14:00
Secnd calgas 57677  , ppm = 392.55, Pressure 1840 psi, start 14:08, stop 14:19 

152: Soil, Site sage, Thu 20-Feb-2003 14:46:55 MST, Deep soil moisture at Sage
To investigate the soil moisture down to the 35 cm depth, a pit was dug at the
Sage site. The pit was 5m or so south of the Darkhorse. near the fence.
It was in a small clear patch with 20-30cm sage bushes some 20-30 cm away.
Several cm of snow was brushed away and with digging bars the ground was broken
to a depth, then the loose dirt collected.
Samples were collected for:
                            0 -  6 cm
                            6 - 10 cm
                           10 - 15 cm
                           15 - 20 cm
                           20 - 25 cm
                           30 - 35 cm
Triplicate aliquots were separated and introduced into the oven at 12:00
The dried aliquots were removed and weighed the next day at 09:45.
Note that the ordering of the aliquots C, B, A, D, E, F is correct

   Sample    Aliquot   Fresh      Dry     Difference      Gravmoist
                         g         g          g
  0 -  6 cm   C I      22.7       19.0       3.7        0.195
              C II     27.5       23.4       4.1        0.175   18.7%
              C III    25.6       21.5       4.1        0.191
  6 - 10 cm   B I      21.8       19.0       2.8        0.147
              B II     18.2       15.8       2.4        0.152   15.0%
              B III    24.3       21.1       3.2
 10 - 15 cm   A I      24.5       21.5       3.0        0.140
              A II     23.0       20.3       2.7        0.152   13.6%
              A III    20.9       18.4       2.5        0.152
 15 - 20 cm   D I      28.0       24.7       3.3        0.134
              D II     25.5       22.5       3.0        0.133   13.2%
              D III    26.1       23.1       3.0        0.130
 20 - 25 cm   E I      21.3       19.0       2.3        0.121
              E II     24.0       21.3       2.7        0.127   12.7%
              E III    24.7       22.0       2.7        0.123
 30 - 35 cm   F I      19.9       17.9       2.0        0.112
              F II     26.2       23.6       2.6        0.110   11.2%
              F III    25.7       23.1       2.6        0.113


















151: Radiometers, Site tower, Wed 19-Feb-2003 16:22:40 MST, Cleaned boom pigs
At 14:30 climbed tower to boom level.
The trolley rolled in with only a small glitch when the trolley had to be
reversed and pulled in again. Note that the catch/release worked.
Used handkerchief to polish domes. Ran out again and "catched"
Inspected from the ground and telephotoed.
150: Sonics, Site tower, Wed 19-Feb-2003 16:01:55 MST, New sonic boom angles at tower
New sonic angles

sonic height        angle             diff from Nov
                    deg min sec       deg min sec
1m                  88 47 35          00 17 15
2m                  88 26 10         -00 04 10
5m                  88 26 10         -00 04 10
10m                 88 41 50          00 11 30
15m                 88 58 15          00 27 55
20m                 88 49 10         -00 09 55
30m                 89 08 05         -00 15 55
149: Base, Site base, Wed 19-Feb-2003 13:37:12 MST, change disk drives
Hard drive 2 removed, replaced with drive 1.
Time was 1:30 pm local.
148: Adams, Site tower, Wed 19-Feb-2003 12:26:10 MST, COSMOS rebooted
COSMOS was rebooted to download the new channel config defining
channel 211. This took place around 11:10 am local.

147: Thermocouples, Site sage, Wed 19-Feb-2003 12:24:10 MST, thermocouples running
The thermocouples are now working at Sage. The data path is
freewave to the tower then through channel 211 of cosmos.

146: CO2, Site tower, Wed 19-Feb-2003 12:13:01 MST, CO2 flow rates adjusted
At 11:30 the CO2 flows were checked and adjusted.
The intake flow was 8 lpm. It was adjusted to 6.5 lpm by adjusting the carbon 
vane pump bypass valve.
The cal gas overflows were then adjusted to 2lpm. The vernier valves were each 
at max and so the adjustment had to be done by adjusting the regulator output
pressure, Hical = 10 psi, lo-cal 12 psi.

    ttl_on  cosmos0       Hi-cal solenoid on
       _off`      0                       off
       _on  cosmos1       Lo-cal solenoid on
       _off       1                       off
    ttl_reset cosmos0     Return to sequence
    ttl_reset cosmos1     Return to sequence

Went along the cylinder to Licor cal gas lines and tightened all swagelocks.
Didnot mess with connections inside Licor box
145: Kryptons, Site all stns, Wed 19-Feb-2003 12:07:08 MST, Kr cleaned
On 19 Feb .Between 11:00 and 11:30 all the tower Kr's were cleaned, starting 
at the lowest and working to the uupermost.
Yesyerday, 18 Feb the Kr's at the outlying sites were all cleaned during the 
radiometer cleaning visits, see logbook # 140
144: Thermocouples, Site sage, Tue 18-Feb-2003 17:57:07 MST, Thermocouples at Sage
Juilen set up her thermocouple array out at Sage today. The Freewave
interface was tested in the base and worked ok. Tomorrow the Freewave
will be moved out to the tower and connected up to a serial channel
on cosmos.

143: Adams, Site tower, Tue 18-Feb-2003 17:55:50 MST, cosmos down at 1:30 pm local
Cosmos was powered down at about 1:30 pm local to add
analog boards. It was back up around 1:45 pm local.

142: Logistics, Site tower, Tue 18-Feb-2003 17:23:04 MST, Guy tensions checked
Between 13:30 and 14:15 checked all guy tensions.
Noted tensions in logbook # 2
141: Logistics, Site tower, Tue 18-Feb-2003 17:18:38 MST, Tower beacon
At noon we climbed to the top of the tower and installed the red blinking beacon
AC power from the 20m ADAM power box.
Beacon operating 
Still operating at 17:30
140: Radiometers, Site all stns, Tue 18-Feb-2003 16:51:56 MST, Radiometer cleaning
Made a tour of the radiometers and cleaned domes and inspected sites.In no case
was there any dirt apparent on cleaning tissue.

14:04 Grass: 10% snow cover only a few cm deep over flat laying long grass 
Downlooking LW and SW have housing but no fans
Up looking LW and SW have housings and fans. LW.in fan stopped. I spun it and 
it started up.

15:00 Sage: Helped Jielun et al locate TC system
           50% snow cover, patches 6-8 cm deep between and among the sage.
Both incoming SW and LW housings and fan operating
Outgoing SW housing, no fan, outgoing LW no housing
All domes cleaned,
Net rad still level
Cleaned Kr, psych fan operating

15:30 Lake 90% snow cover around site.
SW.in fan operating, SW.out housing no fan, LW.out no housing
cleaned all domes, cleaned Kr, psych fan operating

15:40 Scaffold: 50% snow in dappled patches.
4m LW.in fan operating, LW.out no housing
1m LW.out no housing
Cleaned all domes

15:45 Grass (again) SW.in and LW.in fans operating
Only housings, no fan for SW.out and LW.out
Cleaned all domes

15:50 Greasewood: Snow drift 30 -50 cm deep under radiometer
LW.out has no housing.
Cleaned dome without breaking up the snow.
 
139: CO2, Site tower, Tue 18-Feb-2003 16:38:41 MST, CO2 calibration gases
The Lo-cal gas had run out before we arrived on Tues, 18 Feb''03
See notes in Logbook # 136 as to the details of the cylinders dispositions.

The new Lo-cal was attached at 13:15 
The previous two Lo-cal cylinders were transported to the base and will be 
returned to NCAR and then to US Welding.
The cylinders at the tower are, from S to N

Spare Hi-cal    Online Hi-cal    Online Lo-cal    Spare Lo-cal
57677            8/3 F98206         K361224          K 694978
 392               389               283              371
Full (1810)     1380/10          1680/12             Full (1700)

Note that the pressures in the two spare cylinders were determined when at 14:00
on Feb 20 the supllimentary calibrations were carried out.





138: Status, Site all stns, Thu 06-Feb-2003 15:04:38 MST, Daily status, Feb 5
Weather Plots: OK

Rain Gauge: OK

Vbatt: OK

Sonics 2m: OK

Sonics profile: Ok, but w.10m.stn0 dropped out of sync with other sensors between 1600 on Feb 5 and 0200 on Feb 6 (down to ~0.5).  

Krytons: OK, diurnal variation for all sensors.

H2O-kh2o: OK

Tv and P:  OK.

Scalar variance 2m: OK

Fluxes at 2m: OK.

Flux profiles: OK

CO2: Sensor was reset OK

TRH: OK 

Long Wave Rad:. OK. 

|Tdome-Tcase| : .OK.

Tsoil:  . OK

Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake is still not working properly and is showing quite a bit of variation. 

Gsoil: Gsoil.c.10cm.sage and Gsoil.c.1m lake are both still down (see earlier entries).

Soil Thermal Properties:  Starting to look a little better.

U,V spectra stns 1-3: OK.

w,tc spectra stns 1-3: OK

U,V spectra, tower: small amount of white noise in higher frequencies (10+)

w,tc spectra, tower: OK


Kryton Spectra:  The power plots have some white noise at higher frequencies but all in all the signals follow the 2/3 line fairly well.
 

CO2 spectra: quite a bit of white noise in the power plot. 











137: Status, Site all stns, Wed 05-Feb-2003 13:33:55 MST, Daily status, Feb 5
Weather Plots: OK

Rain Gauge: OK

Vbatt: OK

Sonics 2m: OK

Sonics profile: OK

Krytons: OK, diurnal variation for all sensors.

H2O-kh2o: OK

Tv and P:  OK.

Scalar variance 2m: kh2o'kh2o'.2m sensor at tower went off the charts at ~0700 and came back around 0900 MST

Fluxes at 2m: OK.

Flux profiles: OK

CO2: Sensor was reset (see earlier log entry).  Appears to be working fine.

TRH: OK 

Long Wave Rad:. OK. 

|Tdome-Tcase| : .OK.

Tsoil:  . OK

Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake is still not working properly and is showing quite a bit of variation. Qsoil.c.5cm lake(3) appears to be fine.

Gsoil: Gsoil.c.10cm.sage and Gsoil.c.1m lake are both still down (see earlier entries).

Soil Thermal Properties:  Starting to look a little better.

UW spectra: 10.00 f(Hz) between 0000-0030 does not look right, way too low.

Kryton Spectra:  The power plots have some white noise at the higher end of the spectra but the lower frequencies follow the 2/3 line fairly well.

2m sonic spectra: OK

Tower sonic spectra:OK 

CO2 spectra: quite a bit of white noise in the power plot. 










136: CO2, Site tower, Wed 05-Feb-2003 09:46:04 MST, CO2 cal gas status
THIS ENTRY IS NOT YET COMPLETE. NEED TO LOOK THROUGH LOGBOOK SOME MORE
The CO2 Calibration gases used with the Licor 6252 at FLOSS02 were:

Log  Date                Lo-cal                              Hi-cal
             Cylinder Mix rat  Pressure Flow   Cylinder  Mix rat  Pressure Flow
                 
 68  Start   LK438863  323.16   ~1800    Yes   F98206     389.38   ~1800    No
     16 Dec      "        "      ~0      Yes     "          "               No
     16 Dec  LK387670  346.11   ~1700    Yes     "          "               No
 96  06 Jan     "        "        575/15 Yes     "          "       1800/15 No
122  28 Jan     "        "        950/14 Yes     "          "       1580/10 Yes
139  18 Feb     "        "        000/12 No      "          "       1380/10 Yes
139  18 Feb  K 361224   283      1680/12 Yes     "          "         "      "
 ~   21 Feb     "        "       1530/12 Yes     "          "       1320/10  "  




FLOSS02 Cal gas cylinders, 29 Jan''03

57677     392.55   To be used as next Hi-cal          Hi-cal spare
F98206    389.38   Init Hi-cal                        Hi-cal on,28 Jan 1580/10
K694978   371.5                                       Near gate
LK387670  346.11   In Dec Steve O connected as Lo-cal Lo-cal on, 28 Jan 950/14
LK438863  323.16   Init Lo-cal, now empty             EMPTY at base of tower
K361224   283.38   To be used as next Lo-cal          Near gate
K135850   269.83                                      At NCAR
K247699   249.77                                      At NCAR


FLOSS02 Cal gas cylinders, 18 Feb''03 (see LOGBOOK # 139)
---------------------------------------
At base of tower, from S to N 
57677      392.55  To be used as next Hi-cal       Hi-cal spare
8/3 F98206 389.38  Init Hi-cal, still online       Hi-cal still on 18 Feb''03
K361224    283     New Lo-cal , online                     Lo-cal started 18 Feb''03
K694978    371     To be used as next "Lo-cal"     Spare(will use as intake cal)

To be taken back to NCAR and returned to US Welding
LK 387670
LK 438863

At NCAR to be used for lab testing
K135850 269.83
K247699 249.77





135: Status, Site all stns, Tue 04-Feb-2003 14:39:47 MST, Daily status, Feb 4
Weather Plots: OK

Rain Gauge: OK

Vbatt: OK

Sonics 2m: OK

Sonics profile: OK

Krytons: OK

H2O-kh2o: OK

Tv and P:  OK.

Scalar variance: OK

Fluxes at 2m: OK. Lake w'tc'.2m has not been following the general pattern during the day as tower and sage do. Instead of increasing, it basically stays within the same range while the others increase. Gordon believes it is because the lake bed is absorbing more heat than the other sights.

Flux profiles: OK

CO2: Sensor was reset (see earlier log entry).  Appears to be working fine.

TRH: OK 

Long Wave Rad:. OK. 

|Tdome-Tcase| : .OK.

Tsoil:  . Ok. Tsoil.b.2.5.cm sage had a little bleep (~1 deg) between 1300-1500 MST Feb 2. 

Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake is still not working properly and is showing quite a bit of variation. Qsoil.c.5cm lake(3) appears to be fine.

Gsoil: Gsoil.c.10cm.sage and Gsoil.c.1m lake are both still down (see earlier entries).

Soil Thermal Properties:  Starting to look a little better.

Kryton Spectra:  The power plots have some white noise at the higher end of the spectra but the lower frequencies follow the 2/3 line fairly well.

2m sonic spectra: OK

Tower sonic spectra:OK 

CO2 spectra: quite a bit of white noise in the power plot. 








134: Status, Site all stns, Mon 03-Feb-2003 13:46:41 MST, Daily status, Feb 3
Weather Plots: OK

Rain Gauge: OK

Vbatt: OK

Sonics 2m: OK

Sonics profile: w.15m moved in line with other sensors at ~1400 MST Feb 2, and still looks good. 


Krytons: Show increased moisture between 0600-0900 MST at all levels ranging between 2-12 g/m^3.

H2O-kh2o: OK

Tv and P:  OK.

Scalar variance: OK

Fluxes at 2m: Lake w'tc'.2m has not been following the general pattern during the day as tower and sage do. Instead of increasing, it basically stays within the same range while the others increase?

Flux profiles: OK

CO2: Sensor was reset (see earlier log entry).  Appears to be working fine.

TRH: OK, difference between 1m sensor and others is not noticable today (wasn't really over the weekend either). 

Long Wave Rad:. Rlw.out.grease stn0 is lower than other sensors (during daylight) by ~5-10 W/m^2. 

|Tdome-Tcase| : .OK.

Tsoil:  . Ok. Tsoil.b.2.5.cm sage had a little bleep (~1 deg) between 1300-1500 MST Feb 2. 

Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake is doing some crazy stuff.  It came back up at ~1300 on Jan 31, and shows quite a bit of variation over the last two days. At about that same time, Qsoil.c.5cm lake(3) began behaving badly.  It climbed from ~8 %vol to ~32 %vol and over the last 3 days it has slowly come back down.  It looks like it came back into line around 0600 MST today.

Gsoil: Gsoil.c.10cm.sage and Gsoil.c.1m lake are both still down (see earlier entry).

Soil Thermal Properties:  All of these signals look bad, despite fixing
the wiring problem at the tower (grass).  It could take a while to get
good data from these, since there was a small air gap around the probes
after installation; however, the data look even worse.  In particular,
each "event" on Vheat should look like a square wave toggling between 0
and 1 V.  The "Vpile" signal at tower looks close to the signal expected
for Vheat, but at the wrong amplitude.









133: CO2, Site tower, Sat 01-Feb-2003 07:58:15 MST, Licor CO2 back to life
The licor CO2 sensor was reset (a favor from Bill Massman)
and the data looks reasonable starting at 14:00 MST Dec 31.


132: Status, Site all stns, Fri 31-Jan-2003 14:35:56 MST, Daily status, Jan 31
Weather Plots: OK

Rain Gauge: OK

Vbatt: OK

Sonics 2m: OK

Sonics profile: w.15m still reading lower than other sensors  It was checked with a level (see earlier log entry) but apparently that wasn't the problem.


Krytons: OK

H2O-kh2o: OK

Tv and P:  OK.

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2: Sensor crashed after 1130 MST Jan 28, don't know the cause.

TRH: Still concerned about the 1m sensor.  It is warmer and drier, as it should be, but lacks the same characterstics the other profiles have. 

Long Wave Rad:. Rlw.out.grease stn0 is lower than other sensors (during daylight) by ~5-10 W/m^2. 

|Tdome-Tcase| : .OK.

Tsoil:  . Ok. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil: 4 or the 6 Qsoilc.sensors from tower all (with the exception of the 5m sensor)increase to ~15-20 around 1500 MST. This may be a result of melting snow? Iwill keep and eye on it.
Qsoil.c.10cm lake is still both down.

Gsoil: Gsoil.c.10cm.sage and Gsoil.c.1m lake are both still down (see earlier entry).

Soil Thermal Properties:  All of these signals look bad, despite fixing
the wiring problem at the tower (grass).  It could take a while to get
good data from these, since there was a small air gap around the probes
after installation; however, the data look even worse.  In particular,
each "event" on Vheat should look like a square wave toggling between 0
and 1 V.  The "Vpile" signal at tower looks close to the signal expected
for Vheat, but at the wrong amplitude.









131: Status, Site all stns, Thu 30-Jan-2003 13:57:13 MST, Daily status, Jan 30
Weather Plots: OK

Rain Gauge: OK

Vbatt: OK

Sonics 2m: OK

Sonics profile: w.15m still reading lower than other sensors  It was checked with a level (see earlier log entry) but apparently that wasn't the problem.


Krytons: OK

H2O-kh2o: OK

Tv and P:  OK.

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2: Sensor crashed after 1130 MST Jan 28, don't know the cause.

TRH: OK

Long Wave Rad:. Rlw.out.grease stn0 is lower than other sensors by ~5-10 W/m^2. 

|Tdome-Tcase| : .OK.

Tsoil:  . Ok. Tsiol.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil: Qsoilc.1m sage rose to about 13% and slowly dropped, over the last day, down to ~5% (still a little higher than other sensors). Will keep and eye on.
Qsoil.c.10cm lake is still both down.

Gsoil: Gsoil.c.10cm.sage and Gsoil.c.1m lake are both still down (see earlier entry).

Soil Thermal Properties:  All of these signals look bad, despite fixing
the wiring problem at the tower (grass).  It could take a while to get
good data from these, since there was a small air gap around the probes
after installation; however, the data look even worse.  In particular,
each "event" on Vheat should look like a square wave toggling between 0
and 1 V.  The "Vpile" signal at tower looks close to the signal expected
for Vheat, but at the wrong amplitude.







130: Status, Site all stns, Thu 30-Jan-2003 09:38:12 MST, Daily status, Jan 29
Weather Plots: OK

Rain Gauge: OK

Vbatt: OK

Sonics 2m: OK

Sonics profile: w.15m still reading lower than other sensors  It was checked with a level (see earlier log entry) but apparently that wasn't the problem.


Krytons: Ok, no large diurnal variation. There were a couple of spikes (~8-14 g/m^3). 

H2O-kh2o: OK

Tv and P:  OK.

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2: Sensor crashed after 1130 MST Jan 28, don't know the cause.

TRH: OK

Long Wave Rad:. Rlw.out.grease stn0 is lower than other sensors by ~5-10 W/m^2. 

|Tdome-Tcase| : .OK.

Tsoil:  . Ok. Tsiol.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil: Qsoilc.1m sage is acting up. 
Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm.sage became bad as Tony and Steve were working at
sage on 12/18.  The cable was cut and needs to be fixed.

Soil Thermal Properties:  All of these signals look bad, despite fixing
the wiring problem at the tower (grass).  It could take a while to get
good data from these, since there was a small air gap around the probes
after installation; however, the data look even worse.  In particular,
each "event" on Vheat should look like a square wave toggling between 0
and 1 V.  The "Vpile" signal at tower looks close to the signal expected
for Vheat, but at the wrong amplitude.





129: Status, Site all stns, Thu 30-Jan-2003 08:37:51 MST, Daily status, Jan 28
Weather Plots:  There is a small gap in the P, MR, 1.2m, spd, and dir
data from the tower at 1000 MST this morning, it was back up shortly
after.

Rain Gauge: OK

Vbatt: Great. Only down to 13.0 over night.

Sonics 2m: OK

Sonics profile: w.15m still reading lower that other sensors prior to
1000 MST when data gap took place.  Seems to have come in line with
others, may be fixed now.

Krytons: No large diurnal variation but there are some very interesting
spikes in the data.  The first was between 1600-1800 MST Jan 27 where
the 2m, 10m, 20m, and 30m sensors briefly rose to ~5-~8 g/m3.  The
larger of the two happened between 1000-1200 MST Jan 28 when all sensors
shot up (as much as 20 g/m3) and dropped back down (I think) after 1200
MST.

H2O-kh2o: This plot looks bad at times mentioned above in the Kryptons
section.  H2O-kh2o.20m.stn0 is still lower than other sensors by ~-1.0
g/m3.

Tv and P:  OK.

Scalar variance: show large increases corresponding to changes with
krytons, water vapor etc, at ~1600 MST Jan 27, and ~ 1100 MST Jan 28.

Fluxes at 2m: ""

Flux profiles: ""

TRH: RH at sage is a little higher that at lake (~2%) , but there
doesn't appear to be much difference in the temperatures.

Long Wave Rad:. Rlw.in.b.25m and Rlw.out.a.25m are up and running
again.  There were some glitches yesterday, I assume while it was being
worked on, but is looking
good now.

|Tdome-Tcase| : .in.a.26m stn and .in.b.25m stn0 were both off for a
while but began working again around 1600 MST Jan 27.  out.a.25m stn0
was acting up, reading higher than others, but appears to be corrected.

Tsoil:  . Ok. Tsiol.2.5cm.a.sage still looks like it may be exposed to
air.  Im just ignoring this, since we should have adequate Tsoil
sampling.  All Tsoil tower(grass) have a bit nit of noise from the
radiometer fans, but this is small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also be a cable
cut (like the Gsoil.10.cm.a problem found at lake), but will require
more digging up of snow to check.  Currently, I'm ignoring this as well
since there is probably enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm.sage became bad as Tony and Steve were working at
sage on 12/18.  The cable was probably cut.  Needs to be fixed.

Soil Thermal Properties:  All of these signals look bad, despite fixing
the wiring problem at the tower (grass).  It could take a while to get
good data from these, since there was a small air gap around the probes
after installation; however, the data look even worse.  In particular,
each "event" on Vheat should look like a square wave toggling between 0
and 1 V.  The "Vpile" signal at tower looks close to the signal expected
for Vheat, but at the wrong amplitude.




128: Soil, Site lake, Tue 28-Jan-2003 15:07:15 MST, Sage / Lake soil cable problem checks
7/28/03 we examined the soil problems noted in the 'todo' list.

SAGE: 
Very little snow on ground, mostly bare with patches of snow.
gsoil.c.10cm yes indeed the cable was spliced and it appears
dicey to try to fix it where it enters the frozen ground.
Tsoil.a.2.5cm appeared ok and not disturbed even though someone
said it looks 'wierd'  took pictures.

LAKE:
crusted snow cover all over.  Perhaps 4-8" deep.
Gsoil.c.10cm has a wiring break and we were supposed to resolder
it.  We played a bit trying to pull out the cable to find the
'tape job' but decided we were impacting the surface and jeopardizing
the other sensors so stopped.
Qsiol.c.10cm ditto.
Logger connections are secure and look good.
127: Sonics, Site lake, Tue 28-Jan-2003 15:05:15 MST, Sonic Level offsets: lake, 15m
Using the electronic level we observed:

Lake Sonic: .3 degrees sag.
15m Sonic:  .4 degrees sag.
126: Adams, Site tower, Tue 28-Jan-2003 15:00:59 MST, Cosmos Cooler / overheating fixed
1/28/03 The outboard fan on the TECA coolers was bad
and was replaced.  This hopefully should fix the problem
so we won't get anymore lockups.

SteveS thinks that the cooler was working ok otherwise
based on the data.

NOTE: several adam power cycles this morning because of this
work.  Blew breaker trying to heat test the thermostat.  then
a few other cycles while playing around and swapping the fan.
125: Radiometers, Site all stns, Tue 28-Jan-2003 14:57:32 MST, Radiometers Cleaned
7/27/03 25m radiometers cleaned.
7/28/03 All other radiometers cleaned with dry wipe.
Checked inside plenum at Sage and there was no snow packed
inside.  All domes looked very good although we did get
some snow showers this morning which left some moisture on
them.
I noticed something like a smudge inside the psp at Lake
on the small dome.  I don't know if that's a factor.
124: Kryptons, Site all stns, Tue 28-Jan-2003 14:54:04 MST, Kryptons cleaned
7/28/03 cleaned Sage, Lake
7/27/03 cleaned tower sensors.
Used alcohol with q-tip

NOTE: today SteveS mentioned kh2oV actually dropped on
20m unit after we cleaned it.   We examined it and it
was OK.  Re-cleaned with just a dry q-tip and after
looking at cockpit, it appears probably the best now.
I think the alcohol is a poor substitute for distilled
because it may leave a residue esp. if something comes
off the seal and/or qtip.  Plane dry q-tip is better imho.



123: Soil, Site tower, Tue 28-Jan-2003 14:46:32 MST, Grass TP01 soil properties fixed
7/28/03 ~1pm
TP01 should now be up.  We haven't seen a pulse on the Vheat yet
but it's still set to 3-hour cycling and we didn't change that.

Checked wiring after examining sage site which is working.
CR10 wiring OK.  I did note the C7 - Switched 12 goes to
the access point 2 positions away instead of the access
across the terminal blocks adjacent to the +12 input.

Amp-9 connector from probe was bad.  It could never have
worked and apparently wasn't checked in Boulder before
coming up here.
Pin 2 (gnd) was crimped with the insulation preventing
contact.
Pin 1/4 were in wrong positions.
Pin 8 with 150 ohm connector had and has a tenuous
solder to the wire.  We probably broke off when unscrewing
the backshell.

Wiring correct now.
NOTE: the 150ohm connection needs to be redone so it's more
secure.

122: Status, Site all stns, Mon 27-Jan-2003 21:17:58 MST, Jan 27 work: 25m rad. installed/fixed, t.10m, calgas note, guys
1/27/03

TRH.10m fan replaced.  Yes the other one was dead.
TRH.1m bell inlet partially broken (~40%).  This happened
       at the 10m but we decided to move the broken one down
       below where the wind is less.   On 1/28/03 we installed
       one of the unpainted metal bell inlets on this sensor
       to replace the broken bell.

25m Rad/Logger
    Removed logger. Replaced AM416: out=13529, in=13532.
    Wiring was apparently a problem in that:
    1. pyg.in.1 Amp pin-5 to 3-L1 was not seated in the shell
       so we pulled it out to make contact w/cable.
    2. pyg.out.2 Amp wires 7,4 to 6-H2,L2 not connected.
       so we connected them.
    3. CR10 4-H1/L1,H2/L2 were connected to Pyg.in.2
       CR10 5-H1/L1,H2/L2 were connected to Pyg.out.1
       These were apparently reversed per logger pgm
       so we swapped t  122           Status  Updated:28-Jan-2003 15:15  Jan 27 work  hem to
          CR10-4H1/L1 to pins 1,5 pyg.out.1
          CR10-4H2/L2 to pins 7,4 pyg.out.1
          CR10-5H1/L1 to pins 1,5 pyg.in.2
          CR10-5H2/L2 to pins 7,4 pyg.in.2

    Checked s/n of sensors:
          pyg.in.1  = 31975, south-side
          pyg.in.2  = 31979, north-side (just installed today)
          pyg.out.1 = 31974, south-side (just installed today)
          pyg.out.2 = 31977, north-side
    We verified that the cables from the sensors to the logger
    were consistent with the naming conventions.
    We then checked the 'pile' values on cockpit versus putting
    a hand over the sensors: all-4 names was consistent/correct.
    Cleaned domes.
    Per bullet Level check, the electronic sensor bias shows:
          y=-.039
          x=-.100

CalGas: Tony wanted to know what cylinders are here.  They need
        to be identifie  122           Status  Updated:28-Jan-2003 15:15  Jan 27 work  d with a visual marking such as colored
        tape bands to make this obvious !!!

        Hical Spare = 57677
        Hical ON    = IHOP site 8/3 (number was underneath
                      1580psi at 10psi setting
        Local Spare = LK438863
        Local ON    = LK387670
                      950psi, at 14psi setting

Kryptons:
        All tower kryptons cleaned.


Tower: all welds inspected looked OK
Tensioned upper Guy Wires 1-turn each.

         NE     SE       NW      SW
Top     550     500     520      600
4       510     600     610      580
3       430     520     570      500
2       425     580     500      470
Bottom  450     520     440      540


Base: swapped disk
      Added mouse bait
      Retrieved krypton 1101


Cosmos:
      cooler was not running, temp when we checked it was ~24
      We did not get back to trying the heat gun yet.
      The lid insul was left intact for the moment.





121: Status, Site all stns, Mon 27-Jan-2003 15:17:42 MST, Daily status Jan 27
Weather plots:OK

Rain: OK.

Pam power: Is doing excellent, it dropped to just below 13.0 over night.

Sonics 2m:  OK

Sonic profiles: Good. w.15m stn0 still seems to be measuring a tad bit lower than other instruments (~.2 m/s). Position of sensor may be off a little. It will be checked with a level on next visit.

Krytons: No diurnal variation for any of the sensors. kh2o.20m stn0 is reading higher than other sensors by ~.5 g/m^3

H2O-kh2o: HsO-kh2o.2m stn0 is measuring between -.6 and -.9 while other sensors read between -.1 and .4.

Tv and P: pressure sensor data at tower reading higher than sage...will be fixed in post processing.

2m Sonic QC and profile: OK

Sonic variance (2m / profile): OK

Scalar variance 2m: OK

Scalar variance: OK


Fluxes at 2m:  OK

Flux profiles: OK

CO2: OK

CO2 Moments: OK

TRH 2m difference: OK

TRH difference profile: RH profile of sensor at 1m is reading drier, by about 5%, than the other sensors.  John and Tony seem to believe this may be real.

Short wave radiation: OK .

Long Wave Radiation: OK.

|Tdome-Tcase|: OK. 

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.
I'm just ignoring this, since we should have adequate Tsoil sampling.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more
digging
up snow to check.  Currently, I'm ignoring this as well since there
probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and Steve were working at sage
on
12/18.  Problem was not fixed because ground was frozen. Ignore for now.Gsoil.10cm lake has been removed...I think?
Gsoila.10cm tower took a large dip just after 0100 MST from ~0 to -100 W/m^2.


Soil Thermal properties: All of these signals look bad, despite "fixing"

the
wiring problem at tower (grass).  I've expected that it will take a
while to
get good data from these, since there was a small air gap around the
probes
after installation; however, the data look even worse.  In particular,
each
"event" on Vheat should look like a square wave toggling between 0 and 1.

Wind direction profile: OK, though there probably are biases of 1-2
degrees
(despite aligning the booms by eye up and down the tower and shooting
boom
angles that were mostly the same).  We'll need to check these angles
again.

T profile (5 min ave): 10m sensor has been acting up, and reading warmer than others lately.  It will be checked.



















  Kate Beierle 
  Associate Scientist I
  NCAR/ATD
  Research Technology Facilit
























120: Status, Site all stns, Fri 24-Jan-2003 14:20:52 MST, Daily status Jan 24
Weather plots:OK

Rain: OK.

Pam power: OK

Sonics 2m:  OK

Sonic profiles: Good. w.15m stn0 still seems to be measuring a tad bit lower than other instruments (~.2 m/s). Position of sensor may be off a little. It will be checked with a level on next visit.

Krytons: OK. Diurnal variation is back, for all sensors.  Went wacky between 1900 on Jan 23 and 0100 Jan 24.

H2O-kh2o:  OK 

Tv and P: pressure sensor data at tower reading higher than sage...will be fixed in post processing.

2m Sonic QC and profile: OK

Sonic variance (2m / profile): OK

Scalar variance 2m: all kh20'kh2o'.2m sensors spike to ~3-4 between 1800-0000 MST.

Scalar variance: kh20'kh2o'.2m sensors spike to 4 (or greater) between 1800-0000 MST.


Fluxes at 2m:  OK

Flux profiles: OK

CO2: OK

CO2 Moments: OK

TRH 2m difference: OK

TRH difference profile: 10m Temp and RH were off between 0900 and 1400 MST. Temp measurements were coming in ~2-5 deg above other sensors and the RH was reading a maximum of 15% more moist.

Short wave radiation: OK .

Long Wave Radiation: OK.

|Tdome-Tcase|: OK. 

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.
I'm just ignoring this, since we should have adequate Tsoil sampling.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more
digging
up snow to check.  Currently, I'm ignoring this as well since there
probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage
on
12/18.  Problem was not fixed because ground was frozen. Ignore for now.


Soil Thermal properties: All of these signals look bad, despite "fixing"

the
wiring problem at tower (grass).  I've expected that it will take a
while to
get good data from these, since there was a small air gap around the
probes
after installation; however, the data look even worse.  In particular,
each
"event" on Vheat should look like a square wave toggling between 0 and 1.

Wind direction profile: OK, though there probably are biases of 1-2
degrees
(despite aligning the booms by eye up and down the tower and shooting
boom
angles that were mostly the same).  We'll need to check these angles
again.

T profile (5 min ave): 10m sensor has been acting up, and reading warmer than others lately.  It will be checked on next visit (Mon. Jan 27).



















  Kate Beierle 
  Associate Scientist I
  NCAR/ATD
  Research Technology Facilit






















119: Status, Site all stns, Thu 23-Jan-2003 13:49:35 MST, Daily status Jan 23
Weather plots:OK

Rain: OK.

Pam power: OK

Sonics 2m:  OK

Sonic profiles: Good. w.15m stn0 still seems to be measuring a tad bit lower than other instruments (~.2 m/s). Position of sensor may be off a little. It will be checked with a level on next visit.

Krytons: OK

H2O-kh2o:  OK 

Tv and P: pressure sensor data at tower reading higher than sage...will be fixed in post processing.

2m Sonic QC and profile: OK

Sonic variance (2m / profile): OK

Scalar variance 2m:  OK

Scalar variance: OK

Fluxes at 2m:  was more activity than normal at sage on Jan 22 between 0700 and 1500 MST. w'tc'2m increased to .1, values are generally around ~-.1-.05

Flux profiles: OK

CO2: OK

CO2 Moments: OK

TRH 2m difference: sage is higher by a degree between 0900 and 1500 MST on Jan 22. (this difference is also shown in the fluxes at 2m)

TRH profile difference:1m RH is reading a bit more moist than 2m by about 5%. John and Tony agree that this looks reasonable.

Short wave radiation: OK .

Long Wave Radiation: OK.

|Tdome-Tcase|: OK. 

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.
I'm just ignoring this, since we should have adequate Tsoil sampling.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more
digging
up snow to check.  Currently, I'm ignoring this as well since there
probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage
on
12/18.  Problem was not fixed because ground was frozen. Ignore for now.


Soil Thermal properties: All of these signals look bad, despite "fixing"

the
wiring problem at tower (grass).  I've expected that it will take a
while to
get good data from these, since there was a small air gap around the
probes
after installation; however, the data look even worse.  In particular,
each
"event" on Vheat should look like a square wave toggling between 0 and 1.

Wind direction profile: OK, though there probably are biases of 1-2
degrees
(despite aligning the booms by eye up and down the tower and shooting
boom
angles that were mostly the same).  We'll need to check these angles
again.

T profile (5 min ave): 10m sensor has been acting up, and reading warmer than others lately.  It will be checked on next visit (Mon. Jan 27.



















  Kate Beierle 
  Associate Scientist I
  NCAR/ATD
  Research Technology Facilit




















118: Soil, Site all stns, Wed 22-Jan-2003 15:39:36 MST, Soil sensor configuration
To view the configuration of the soil sensors at Grass, Sage and Lake:

Hit "Graphics Viewer", and then the little diamond. 

117: Status, Site all stns, Wed 22-Jan-2003 15:31:46 MST, Daily status Jan 22
Weather plots:OK

Rain: No data shown in first plot.

Pam power: OK

Sonics 2m:  OK

Sonic profiles: Good. w.15m stn0 still seems to be measuring a tad bit lower than other instruments (~.2 m/s). Position of sensor may be off a little. It will be checked with a level on next visit.

Krytons: OK

H2O-kh2o:  OK 

Tv and P: pressure sensor data at tower reading higher than sage...will be fixed in post processing.

2m Sonic QC and profile: OK

Sonic variance (2m / profile): OK

Scalar variance 2m:  OK

Scalar variance: OK

Fluxes at 2m:  OK

Flux profiles: OK

CO2: OK

CO2 Moments: OK

TRH 2m difference: OK

TRH profile difference:1m RH is reading a bit more moist than 2m by about 5%. John and Tony agree that this looks reasonable.

Short wave radiation: OK .

Long Wave Radiation: OK.

|Tdome-Tcase|: OK. 

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.
I'm just ignoring this, since we should have adequate Tsoil sampling.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more
digging
up snow to check.  Currently, I'm ignoring this as well since there
probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage
on
12/18.  Problem was not fixed because ground was frozen. Ignore for now.


Soil Thermal properties: All of these signals look bad, despite "fixing"

the
wiring problem at tower (grass).  I've expected that it will take a
while to
get good data from these, since there was a small air gap around the
probes
after installation; however, the data look even worse.  In particular,
each
"event" on Vheat should look like a square wave toggling between 0 and 1.

Wind direction profile: OK, though there probably are biases of 1-2
degrees
(despite aligning the booms by eye up and down the tower and shooting
boom
angles that were mostly the same).  We'll need to check these angles
again.





















  Kate Beierle 
  Associate Scientist I
  NCAR/ATD
  Research Technology Facilit



















116: Soil, Site all stns, Wed 22-Jan-2003 12:57:12 MST, TP01 wiring
The graphics file shows the wiring ofr the TP01 probe. The color code
for the wires corrresponds to the probe wires and NOT the wires in the
logger box.
115: Status, Site all stns, Tue 21-Jan-2003 14:50:59 MST, Daily status Jan 21
Weather plots:OK

Rain: OK

Pam power: OK

Sonics 2m:  OK

Sonic profiles: Good. w.15m stn0 seems to be measuring a tad bit lower than other instruments (~.2 m/s).

Krytons: OK

H2O-kh2o:  OK 

Tv and P: pressure sensor data at tower reading higher than sage and lake by about a millibar.

2m Sonic QC and profile: OK

Sonic variance (2m / profile): OK

Scalar variance 2m:  OK

Scalar variance: OK

Fluxes at 2m:  OK

Flux profiles: OK

CO2: OK

CO2 Moments: OK

TRH 2m difference: OK

TRH profile difference:OK

Short wave radiation: OK.

Long Wave Radiation: OK.

|Tdome-Tcase|: OK. 

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.
I'm just ignoring this, since we should have adequate Tsoil sampling.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more
digging
up snow to check.  Currently, I'm ignoring this as well since there
probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage
on
12/18.  Problem was not fixed because ground was frozen. Ignore for now.


Soil Thermal properties: All of these signals look bad, despite "fixing"

the
wiring problem at tower (grass).  I've expected that it will take a
while to
get good data from these, since there was a small air gap around the
probes
after installation; however, the data look even worse.  In particular,
each
"event" on Vheat should look like a square wave toggling between 0 and 1.

Wind direction profile: OK, though there probably are biases of 1-2
degrees
(despite aligning the booms by eye up and down the tower and shooting
boom
angles that were mostly the same).  We'll need to check these angles
again.





















  Kate Beierle 
  Associate Scientist I
  NCAR/ATD
  Research Technology Facilit

















114: Status, Site all stns, Fri 17-Jan-2003 14:42:35 MST, Daily status Jan 17
Weather plots:OK

Rain: OK

Pam power: OK

Sonics 2m:  OK

Sonic profiles: OK. 

Krytons: OK

H2O-kh2o:  OK 

Tv and P: pressure sensor data at tower dropped below sage and lake between 1300 and 1600 Jan 16, then rose again to between 2-4 mb greater than sage and lake.

2m Sonic QC and profile: OK

Sonic variance (2m / profile): OK

Scalar variance 2m:  OK

Scalar variance: OK

Fluxes at 2m:  OK

Flux profiles: OK

CO2: OK

CO2 Moments: OK

TRH 2m difference: OK

TRH profile difference: 10m sensor measurements varied from the others between 
0900 MST Jan 16 and 1600 MST Jan 16 by as much as 6 deg C. Rebounded after 1600 and appears fine since.

Short wave radiation: OK.

Long Wave Radiation: OK.

|Tdome-Tcase|: all sensors fall out of sync after 0800 MST, as of 1400 they had not corrected themselves. 

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.
I'm just ignoring this, since we should have adequate Tsoil sampling.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more
digging
up snow to check.  Currently, I'm ignoring this as well since there
probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage
on
12/18.  Problem was not fixed because ground was frozen. Ignore for now.


Soil Thermal properties: All of these signals look bad, despite "fixing"

the
wiring problem at tower (grass).  I've expected that it will take a
while to
get good data from these, since there was a small air gap around the
probes
after installation; however, the data look even worse.  In particular,
each
"event" on Vheat should look like a square wave toggling between 0 and 1.

Wind direction profile: OK, though there probably are biases of 1-2
degrees
(despite aligning the booms by eye up and down the tower and shooting
boom
angles that were mostly the same).  We'll need to check these angles
again.

w'kh2o' profiles:  01:00 profiles looks bad, 2 and 10 m sensors read w'kh2o' below -.0005 while the 20m and 30m sensors are ready w'h2o' at about .0005??



















  Kate Beierle 
  Associate Scientist I
  NCAR/ATD
  Research Technology Facilit
















113: Status, Site all stns, Thu 16-Jan-2003 15:21:17 MST, Daily status Jan 16
Weather plots:OK

Rain: OK

Pam power: OK

Sonics 2m:  OK

Sonic profiles: OK. 

Krytons: kh2o.2m lake(3) spiked to ~7 (g/m^3) and kh2o.20m.stn0 spiked to ~4 at 2400 Jan 15 MST. 

H2O-kh2o:  OK 

Tv and P: pressure sensor data at tower is still slightly higher then at other sites. (~2-3 mb)  Tbox.cosmos.stn0 is consistently warmer than daisy.


2m Sonic QC and profile: OK

Sonic variance (2m / profile): OK

Scalar variance 2m:  OK

Scalar variance: OK

Fluxes at 2m:  OK

Flux profiles: OK

CO2: OK

CO2 Moments: OK

TRH 2m difference: OK

TRH profile difference: 10m, for both RH and T were measuring about the same as the other sensors until about 1000 MST when the temperature began to increase(RH decreased). It continued to warm to 7 deg (~peaked at 1100) and didn't come back into sync until 1600 MST. 

Short wave radiation: OK.

Long Wave Radiation: OK.

|Tdome-Tcase|: Tdome-Tcase.out...all sensors fall out of sync a bit between 0900 and 1700 MST 

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.
I'm just ignoring this, since we should have adequate Tsoil sampling.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more
digging
up snow to check.  Currently, I'm ignoring this as well since there
probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage
on
12/18.  Problem was not fixed because ground was frozen. Ignore for now.
Problems with Gsoil.c.10cm.lake, it should be checked.

Soil Thermal properties: All of these signals look bad, despite "fixing"

the
wiring problem at tower (grass).  I've expected that it will take a
while to
get good data from these, since there was a small air gap around the
probes
after installation; however, the data look even worse.  In particular,
each
"event" on Vheat should look like a square wave toggling between 0 and 1.

Wind direction profile: OK, though there probably are biases of 1-2
degrees
(despite aligning the booms by eye up and down the tower and shooting
boom
angles that were mostly the same).  We'll need to check these angles
again.





















  Kate Beierle 
  Associate Scientist I
  NCAR/ATD
  Research Technology Facilit
















112: Status, Site all stns, Wed 15-Jan-2003 15:07:47 MST, Daily status Jan 15
Weather plots:OK

Rain: OK

Pam power: OK

Sonics 2m:  OK

Sonic profiles: OK. w.15m stn0 is still reading a little lower (~-.2) than other sensors.

Krytons: plots are showing more moisture during early evening and early morning then throughout the night?  

H2O-kh2o:  OK 

Tv and P: pressure sensor data at tower is looking better, it matches sage and lake around 1200 MST Jan 14, then begins to break away a bit after 1800 MST jan 14.  After that, difference is about 2 mb.  Tbox.cosmos.stn0 is consistently warmer than daisy.


2m Sonic QC and profile: OK

Sonic variance (2m / profile): OK

Scalar variance 2m:  OK

Scalar variance: OK

Fluxes at 2m:  OK

Flux profiles: OK

CO2: OK

CO2 Moments: OK

TRH 2m difference: OK

TRH profile difference: 1m is more moist over night by ~5%. After 0700 MST all sensors measure about the same.

Short wave radiation: OK.

Long Wave Radiation: OK.

|Tdome-Tcase|: OK.

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.
I'm just ignoring this, since we should have adequate Tsoil sampling.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more
digging
up snow to check.  Currently, I'm ignoring this as well since there
probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage
on
12/18.  Problem was not fixed because ground was frozen. Ignore for now.
Problems with Gsoil.c.10cm.lake, it should be checked.

Soil Thermal properties: All of these signals look bad, despite "fixing"

the
wiring problem at tower (grass).  I've expected that it will take a
while to
get good data from these, since there was a small air gap around the
probes
after installation; however, the data look even worse.  In particular,
each
"event" on Vheat should look like a square wave toggling between 0 and 1.

Wind direction profile: OK, though there probably are biases of 1-2
degrees
(despite aligning the booms by eye up and down the tower and shooting
boom
angles that were mostly the same).  We'll need to check these angles
again.





















  Kate Beierle 
  Associate Scientist I
  NCAR/ATD
  Research Technology Facilit













111: Status, Site all stns, Tue 14-Jan-2003 14:24:22 MST, Daily status Jan 14
Weather plots:OK

Rain: OK

Pam power: OK

Sonics 2m:  OK

Sonic profiles: OK. w.15m stn0 is coming in a little bit lower (~-.2- -.4) than other sensors (.4 - -.1)

Krytons: Diurnal variation is not present for any sensors.  

H2O-kh2o:  H20-kh20.2m sage had glitch from .6 to 1.2 on Jan 13 at 1300 MST 

Tv and P: pressure sensor data at tower is a bit higher (by about 2 mb) than at other sites.  Tbox.cosmos.stn0 is consistently about 5 deg warmer than daisy.


2m Sonic QC and profile: OK

Sonic variance (2m / profile): OK

Scalar variance 2m:  OK

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2: OK

CO2 Moments: OK

TRH 2m difference: OK

TRH profile difference: 1m -.5m is about 4% drier than other differences measured.

Short wave radiation: OK.

Long Wave Radiation: OK, Rlw.in.b.25m and Rlw.out.a.25m stn0 were removed.

|Tdome-Tcase|: OK, out.grease stn0 no longer being plotted either.

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.
I'm just ignoring this, since we should have adequate Tsoil sampling.
All Tsoils.tower(grass) have a bit noise from the radiometer fans, but
this is
small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more
digging
up snow to check.  Currently, I'm ignoring this as well since there
probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage
on
12/18.  Problem was not fixed because ground was frozen. Ignore for now.
Problems with Gsoil.c.10cm.lake, it should be checked.

Soil Thermal properties: All of these signals look bad, despite "fixing"

the
wiring problem at tower (grass).  I've expected that it will take a
while to
get good data from these, since there was a small air gap around the
probes
after installation; however, the data look even worse.  In particular,
each
"event" on Vheat should look like a square wave toggling between 0 and 1.

Wind direction profile: OK, though there probably are biases of 1-2
degrees
(despite aligning the booms by eye up and down the tower and shooting
boom
angles that were mostly the same).  We'll need to check these angles
again.





















  Kate Beierle 
  Associate Scientist I
  NCAR/ATD
  Research Technology Facilit











110: Kryptons, Site tower, Mon 13-Jan-2003 14:18:15 MST, krypton at 20 meters

On Jan 6 at noon, SteveO/Cody removed krypton 1392 at 20 meters
and replaced it with 1390. 

The cal_file for kh2o.20m is 
$ASTER/projects/FLOSSII/cal_files/krypton.20m

At first an incorrect V0 for 1390 was used, which resulted in negative
densities, which the calib function set to 0.

On jan 13 the correct V0 was entered in the cal file and
the data since Jan 6 rerun and plots remade.

kh2o.20m from 1390 looks good now, and does not have the diurnal
"light-leak" appearance of 1392. 

109: Status, Site all stns, Fri 10-Jan-2003 14:56:44 MST, Daily status Jan 10
Weather plots:Pressure at tower dropped yesterday afternooon just after 1300, to 751. It rebounded at 1700 MST

Rain: Top plot has no data.

Pam power: OK

Sonics 2m:  OK

Sonic profiles: OK

Krytons: Diurnal variation is present for all sensors.  Sensors show small moisture increase from  0100-0130 MST .
 
H2O-kh2o:  20m tower still showing signs that it may have a light leak? 2m tower
shows a small variation from what the other sensors are outputting.  It increased ~1.5 g/m^3 between 0100-0500 MST jan 9.

Tv and P: pressure sensor data at tower is quite a bit higher (by about 2 mb) than at other sites.  Tbox.cosmos.stn0 jumped from ~22 deg to ~58 deg between 0100and 0500 MST Jan 9.  Doubt that its real because temp shot straight up to 58 deg, apparently stayed there for 4 hours and dropped straight back to ~20 deg.


2m Sonic QC and profile: Small but quick increase in 10m for all components (u,v,w,diag), between 0300 and 0500 MST.

Sonic variance (2m / profile): v'v'.2m shot up to almost 7 (m/s)^2 at 0500 MST./ OK

Scalar variance 2m:  kh2o'kh2o'.2m steadily climbed to 4+ over night, suddenly dropped to near 0 at 0600 MST. Remained at/near zero until 1200 MST when all the sensors increeased to between 2.5-3.5. tc'tc'.2m showed a lot of variation over night, up until about 0500 when it dropped to almost zero where it has stayed

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2: OK

CO2 Moments: OK

TRH 2m difference: Temp dropped from ~-1 to ~-13 just before 0500, RH also decreased at this time from ~2% to ~-18%. Both recovered fairly quickly.

TRH profile difference: fairly significant temp change (drops 5-10 deg), and decrease in RH (10-20%) at 0500 MST. 

Short wave radiation: OK.

Long Wave Radiation: OK, but Rlw.in.b.25m and Rlw.out.a.25m stn0, which were having having problems are not longer there?

|Tdome-Tcase|: OK, in.b.25m stn0 is no longer being plotted. out.grease stn0 no longer being plotted either.

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.
I'm just ignoring this, since we should have adequate Tsoil sampling.
All Tsoils.tower(grass) have a bit noise from the radiometer fans, but
this is
small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more
digging
up snow to check.  Currently, I'm ignoring this as well since there
probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage
on
12/18.  Problem was not fixed because ground was frozen. Ignore for now.
Gsoil.c.10cm is back on the fritz. Also, keep an eye in Gsoil.a.10cm.sage, it seems to be behaving starngely.

Soil Thermal properties: All of these signals look bad, despite "fixing"

the
wiring problem at tower (grass).  I've expected that it will take a
while to
get good data from these, since there was a small air gap around the
probes
after installation; however, the data look even worse.  In particular,
each
"event" on Vheat should look like a square wave toggling between 0 and 1.

Wind direction profile: OK, though there probably are biases of 1-2
degrees
(despite aligning the booms by eye up and down the tower and shooting
boom
angles that were mostly the same).  We'll need to check these angles
again.





















  Kate Beierle 
  Associate Scientist I
  NCAR/ATD
  Research Technology Facilit









108: Status, Site all stns, Thu 09-Jan-2003 13:06:39 MST, Daily status Jan 9
Weather plots:Pressure at tower dropped yesterday afternooon to, or below, 752. It rebounded, and then dropped until about 1700 MST, Jan 8, and came back up around 2300.

Rain: From looking at top plot, there's no moisture, no calibraiton bleeps, and no sign of life.  

Pam power: OK

Sonics 2m:  OK

Sonic profiles: w data from 10m drops from around 0 to -.5 m/s shortly before 0200 MST, and rebounds shortly after 0500 MST. None of other sensors behaved this way.

Krytons: Diurnal variation is present for all sensors except at 30m. 2m sensor misbehaves between 1700-2300  MST Jan 8.
 
H2O-kh2o: 20m and 30m show no diurnal variation. 2m jumps from -.8 to 1.3 at 1700, drops slightly over the next few hours, then returns to negative values. 

Tv and P: pressure sensor behavior at tower noted above in weather section.


2m Sonic QC and profile: Flares up a bit over night, reading dropped again around 0900 MST.

Sonic variance (2m / profile):  OK 

Scalar variance 2m: OK  

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

CO2: between 1700-2400 Jan 8 co2mr.2m, tlicor2m, and plicor.2m flatlined. Pressure also jumped from ~640 to ~840 mb during this time. 

CO2 Moments: no moments data shown during times listed above.

TRH 2m difference:  OK

TRH profile difference: .  OK

Short wave radiation: OK.

Long Wave Radiation: OK but Rlw.in.b.25m and Rlw.out.a.25m stn0, which were having having problems are not longer there?

|Tdome-Tcase|: OK, in.b.25m stn0 is no longer being plotted. out.grease stn0 no longer being plotted either.

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.
I'm just ignoring this, since we should have adequate Tsoil sampling.
All Tsoils.tower(grass) have a bit noise from the radiometer fans, but
this is
small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more
digging
up snow to check.  Currently, I'm ignoring this as well since there
probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage
on
12/18.  Problem was not fixed because ground was frozen. Ignore for now.
Gsoil.c.10cm lake was on the fritz prior to 1500 Jan 8 but seems to have
recovered.

Soil Thermal properties: All of these signals look bad, despite "fixing"

the
wiring problem at tower (grass).  I've expected that it will take a
while to
get good data from these, since there was a small air gap around the
probes
after installation; however, the data look even worse.  In particular,
each
"event" on Vheat should look like a square wave toggling between 0 and 1

V.
The "Vpile" signal at tower looks close to the signal expected for
Vheat, but
at the wrong amplitude.

Wind direction profile: OK, though there probably are biases of 1-2
degrees
(despite aligning the booms by eye up and down the tower and shooting
boom
angles that were mostly the same).  We'll need to check these angles
again.

u* profiles: First plot there is not data for sensors at 30 m between 1:00-6:00 MST?



















  Kate Beierle 
  Associate Scientist I
  NCAR/ATD
  Research Technology Facilit






107: Weather, Site tower, Thu 09-Jan-2003 09:09:08 MST, riming events
Here are some (probably not all) icing events I see in the Axis camera images:

Dec 15 08:00?
Dec 24 08:00
Dec 28 08:00?
Jan 1 08:00
Jan 7 08:00
Jan 8 08:00

106: Status, Site all stns, Wed 08-Jan-2003 14:03:31 MST, Daily status Jan 8
Weather plots:rainr tower(1) had a spike from -25 to -3 g/kg at 1530 on Jan 7.

Rain: OK

Pam power: OK

Sonics 2m:  w.2m tower was behaving strangely from 0700-0900 MST. 

Sonic profiles: data from 10m is quite different than that
coming from other sensors.  for u component Between 0600-0900 the
10m sensor was reading between  -2 and -4 m/s, while others were close
to (or above) 0 m/s. At one point it jumped to 2 m/s from -2 m/s.The v componentduring this time ranged between 2 and 9 m/s.  The w component for the 5m, 10m and 15m sensors at that time looks bad also.  All sensors seem to recover
shortly after 0900 MST.

Krytons: OK. Diurnal variation is present for all sensors.  

H2O-kh2o: 20m still shows a diurnal difference (light leak?).  This
didn't
change with cleaning on about 12/17.

Tv and P: tc.1m stn0 may have had a glitch just after 0800.  According to the sensor the temperature drops from ~-21 to ~-29 between 0800 and 0900 MST which is almost 10 deg colder than at other heights.


2m Sonic QC and profile: Flares up a bit over night, reading drop again around 0900 MST.

Sonic variance (2m / profile):  OK / Lots of spikes between 0400 and 0900 MST.

Scalar variance 2m: kh20'kh2o'.2m at tower steadily increased between 2100 MST Jan 7 and 0200 Jan 8 MST until it reached 4 where it has stayed.  There is very little evidence of sage...must be at or very close to zero.  

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: Very irregular between 0200 and 0900 MST.  

CO2: OK

CO2 Moments: OK

TRH 2m difference:  it seems that when RH.2m sage-tower drops significantly lake-tower increases by about the same magnitude.

TRH profile difference: .  OK

Short wave radiation: OK.

Long Wave Radiation: Rlw.inb.25m looks to have crashed. Rlw.out.a.25m stn0 is continueing to experiance problems. 

|Tdome-Tcase|: out.a.25m is reading higher than others. Also, b.25m.stn0 appears to have crashed.

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.
I'm just ignoring this, since we should have adequate Tsoil sampling.
All Tsoils.tower(grass) have a bit noise from the radiometer fans, but
this is
small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more
digging
up snow to check.  Currently, I'm ignoring this as well since there
probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage
on
12/18.  Problem was not fixed because ground was frozen. Ignore for now.


Soil Thermal properties: All of these signals look bad, despite "fixing"

the
wiring problem at tower (grass).  I've expected that it will take a
while to
get good data from these, since there was a small air gap around the
probes
after installation; however, the data look even worse.  In particular,
each
"event" on Vheat should look like a square wave toggling between 0 and 1

V.
The "Vpile" signal at tower looks close to the signal expected for
Vheat, but
at the wrong amplitude.

Wind direction profile: OK, though there probably are biases of 1-2
degrees
(despite aligning the booms by eye up and down the tower and shooting
boom
angles that were mostly the same).  We'll need to check these angles
again.




















  Kate Beierle 
  Associate Scientist I
  NCAR/ATD
  Research Technology Facilit


105: CO2, Site tower, Wed 08-Jan-2003 11:57:06 MST, Shortened CO2 cal periods
Jan 8, 11:40am

2 minute CO2 calibration periods are longer than necessary - the
signal settles to the new level within 10 seconds, so Tony
recommended shortening the cal periods to 30 seconds for each cal gas.

The sequence is now:

        XX:00:00 hi cal on
        XX:00:30 low cal on
        XX:01:00 normal sampling
        ...
        7+XX:00:00 repeat same thing 7 hours later.

Rebooted cosmos

104: Log, Site all stns, Wed 08-Jan-2003 11:09:25 MST, yesterday's tasks
Today was another day filled with tasks:

Condtions:: started foggy, cleared shortly after sunrise, but things stayed rimed for
several hours.  Very light winds even up the tower, though not as calm as yesterday.
Pretty clear skies, with scattered cirrus now at the end of the day.

Sage (~8:00-9:30)
- found everything rimed due to the fog.
- found where Tony had sliced through the Gsoil.10cm.c cable and spliced the cut, but
data were still bad.  Left it broken (even though we had brought a spare, I couldn't
see how to install it in the frozen ground).
- corrected swapped wires for TP01 (4H&5H) [which fixed this problem]
- cleaned rads
- cleaned krypton & blew rime away from sonic

Tower
- Gordon rebooted cosmos since the Datel channels were all constant overnight (~1800-)
- cleaned 4m.bare
- found by swapping 1.5m.bare and grease that problem was with adaptor connector.
brought cable back to base to fix.
- (found that we had to reboot the Toughbook to get it to talk to com1, in the process
rebooted cosmos again.)

Lake (While Cody was effecting the repair, I went to Lake)
- cleaned all rads (net has ice inside dome -- probably should replace dome next time)
- swapped 4H&5H on TP01 here as well
- Got Gordon and Steve to start 10 min TP01 cycling at lake and sage

Base
- Got printer to work by removing, shaking, and reinstalling color cartridge
- charged ScopeMeter

Sage (~1130-1200)
- after Steve and Gordon reported noisy data from TP01, I monitored the signals with a
DVM.  Found TP01 working okay, but main PAM power cycling between 13.3-14.9 V.  The
charge controller (began with "F"?) was clicking every 5 seconds corresponding to the
min and max of the voltage cycling.  Temporarily fixed problem by removing solar panels
(and wind generator, but WG by itself didn't stop cycling).  [Later reconnected ~1600]

Tower (Cody started these tasks while I was at Sage)
- reconnected bare.1.5m
- cleaned bare.1.5m
- reconnected grease
- cleaned grease
- cleaned all rads at grass
- checked that AMP connector on TP01 was wired properly (it was, though later found that
the colors are wrong)
- asked Gordon to turn off TP01 10-min cycling (back to 3 hr) at sage and lake and
turn on 10-min cycling at grass.
- checked that all logger screws were tight at 25m (they were)
- pulled in.b.25m and out.a.25m after doing "hand warming" test of (in order): 
out.b.25m, out.a.25m, in.a.25m, in.b.25m [have confirmed that the channels were connected
as we expected: "a"="1"=south; "b"="2"=north]
~1500
- removed hauling rope and secured davit
- more tests of TP01 found that cable colors from sensor itself are probably wrong.
I couldn't make sense of them by measuring the resistance between wires.  Left probe
connected with a guess at wiring, but probably wrong.  Asked Gordon to reset cycling back
to 3 hrs.
- Dumped the calibrated amount of water into the rain gauge.  The fluid level is about
at the midpoint of the cylinder (about 50cm below the top and about 8 cm above the
top of the "box").
- noted that only 490psi was left in Lo-Cal, so changed to new Lo-Cal bottle.  We forgot
to note the new pressure, we think 1700-1800.  My plan is to switch back to the first
Lo-Cal during the IOP, when we can get to it quickly once it runs out.
- finally removed stepladder from bare

Base (~1540)
- Stopped charging ScopeMeter
- Swapped USB disk
- Turned off oven

Sage (~1600)
- Reconnected solar panels

Left site.


103: Status, Site all stns, Wed 08-Jan-2003 09:44:45 MST, Daily status Jan 7
Weather plots: Ptower is having problems and should be checked
out.

Rain: No data shown in first plot (raina stn0 and rainr.grass stn0).
Evaporation plot don't look good either.

Pam power: OK, but power at sage dropped from ~14 to ~12.7 at about
0100...battery was temporarily disconnected,hould be coming back up
soon.

Sonics 2m:  u.2m at lake began to increase at after 0400 MST and peaked
at 19 m/s at 0800 MST while other stations remained near ~2m/s.  w.2m
lake is looking bad.  It drops to ~-1.0 m/s between 0400 and 0800,
during this time sage and tower spiked to ~-.8 and ~-.9.

Sonic profiles: data from 10m  and 15m  is quite different than that
coming from other sensors.  for u component Between 0600-0900 the
sensors were reading between  -10 and -15 m/s, while others were close
to 0 m/s. However the v component at these heights climbed to between
+10 and +15 m/s during this time.  w component for most of the sensors
around this time frame looks bad too.  All sensors seem to recover
shortly after 1000 MST.

Krytons: OK. Diurnal variation is present except at 2m.  There appears
to be very little moisture over night at that level and voltage never
drops until 0900 MST.  It seems to be working properly again after 0900.

H2O-kh2o: 20m still shows a diurnal difference (light leak?).  This
didn't
change with cleaning on about 12/17.

Tv and P:  Looks like pressure sensor was down over night.  Brought back
up at 0900 MST

2m Sonic QC and profile: Flares up a bit between 0400 and 1000 MST

Sonic variance (2m and profile):  All componnets flare up a bit between
0400 and 1000 MST with values reaching 10 m/s^2.

Scalar variance 2m:  tc'tc'.2m lake got up to ~3.0 between 0400-1000
MST.  kh20'kh2o'.2m got up to ~3.5 (g/m^3)^2 between
0400-0600 MST

Scalar variance: OK

Fluxes at 2m: Same deal, flares and glitches between 0400-1000 MST

Flux profiles: ""

CO2: Im not really sure what to make of this data.  Sensor seems to
have gone down between 1800 on Jan 6 and 0930 on Jan 7, and then
recovered.

CO2 Moments: ""

TRH 2m difference:  RH.2m sage-tower drops about 25% between 1000 and
1400 MST.

TRH profile difference: .  OK

Short wave radiation: OK.

Long Wave Radiation: Rlw.in.a25m and Rlw.inb.25m were acting up
yesterday afternoon, calmed over night and then 25 m data jumped quite a
bit higher than others and remains there.  Rlw.out.grease was not
looking good, looked like it was only reading 100
w/m^2 rather than 300 w/m^2 like the others, but recovered around 1100
MST.  Rlw.out.a.25m stn0 seems to have started experiencing problems at
0200 MST.

|Tdome-Tcase|: out.grease.stn0 should be checked, also should check both
in and out of  b.25m.stn0.

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.
I'm just ignoring this, since we should have adequate Tsoil sampling.
All Tsoils.tower(grass) have a bit noise from the radiometer fans, but
this is
small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more
digging
up snow to check.  Currently, I'm ignoring this as well since there
probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage
on
12/18.  One of us probably cut a cable.  This should be fixed next
visit.

Soil Thermal properties: All of these signals look bad, despite "fixing"

the
wiring problem at tower (grass).  I've expected that it will take a
while to
get good data from these, since there was a small air gap around the
probes
after installation; however, the data look even worse.  In particular,
each
"event" on Vheat should look like a square wave toggling between 0 and 1

V.
The "Vpile" signal at tower looks close to the signal expected for
Vheat, but
at the wrong amplitude.

Wind direction profile: OK, though there probably are biases of 1-2
degrees
(despite aligning the booms by eye up and down the tower and shooting
boom
angles that were mostly the same).  We'll need to check these angles
again.




















  Kate Beierle 
  Associate Scientist I
  NCAR/ATD
  Research Technology Facilit

102: Status, Site all stns, Tue 07-Jan-2003 16:31:04 MST, Daily status Jan 6
One of the computers died on Jan 4, 2002 and affected all datasets
collected from sensors mounted near the bottom of the tower, through
0900 MST on Jan 6.

Pam power: OK.

Sonics:  No uvw data from tower at or below 5m.  Others look fine.

Krytons: OK. Diurnal variation is back.  kh20.20m.stn0 spikes to 24
g/m^3at noon MST, appears to be a data glitch.

H2O-kh2o: 20m still shows a diurnal difference (light leak?).  This
didn't
change with cleaning on about 12/17.

Tv and P: OK

Sonic variance 2m: OK

Scalar variance 2m:  kh20'kh2o'.2m got up to +4 (g/m3)2 between
0400-0600 MST

Scalar variance: tc'tc'.10m has a small gap of  missing data at 0000
MST.  kh2o'kh2o'.10m stn0 has missing data just after 0500 MST and then
no data after 0830 MST??

Fluxes at 2m: OK

Flux profiles: OK

CO2: Very little data because of downed computer.

CO2 Moments: Very little data.

TRH profile difference: .  5m and 30m may not have been working
correctly.  Between  1000 and 1300 MST, when computer was brought back
up, 5m sensor was reading between 2 and 6 degrees higher than others and
was drier by 10-25%.  30m tower looks like it had a glitch just before
noon.  Both seem to have recovered.

Short wave radiation: OK.

Long Wave Radiation: glitch in Rlw.out.4m.bare.stn0 and
Rlw.out.bare.stn0 of ~100(W/m2), both recovered almost immediately.
Rlw.out.grease is not looking good, looks like it's only reading 100
w/m2 rather than 300 w/m2 like the others.

|Tdome-Tcase|: out.grease.stn0 should be checked.

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to
air.
I'm just ignoring this, since we should have adequate Tsoil sampling.
All Tsoils.tower(grass) have a bit noise from the radiometer fans, but
this is
small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more
digging
up snow to check.  Currently, I'm ignoring this as well since there
probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage
on
12/18.  One of us probably cut a cable.  This should be fixed next
visit.

Soil Thermal properties: All of these signals look bad, despite "fixing"
the
wiring problem at tower (grass).  I've expected that it will take a
while to
get good data from these, since there was a small air gap around the
probes
after installation; however, the data look even worse.  In particular,
each
"event" on Vheat should look like a square wave toggling between 0 and 1
V.
The "Vpile" signal at tower looks close to the signal expected for
Vheat, but
at the wrong amplitude.

Wind direction profile: OK, though there probably are biases of 1-2
degrees
(despite aligning the booms by eye up and down the tower and shooting
boom
angles that were mostly the same).  We'll need to check these angles
again.

101: Soil, Site all stns, Tue 07-Jan-2003 13:40:58 MST, testing of thermal properties probes
At sage, 11:59 am, Jan 7, set the thermal properties cycle to 10 minutes
(from 180 minutes) so that Steve/Cody could check the voltages.

At 12:05, did the same at lake.

Steve determined at sage that the variation seen on Vheat is due to the
charging system.


At 1:30 pm, changed sage back to 180 minute cycle.  At 1:32, changed
lake back.

At 1:37pm, changed grass to 10 minutes.
At 3:51pm, changed grass back to 180 minutes.

The TP instructions are numbers 68 and 72

The values for normal 3 hour cycling are:
MODE 01:0000  68\r\n
68:P92  \r\n
01:0000  \r\n
02:0180  \r\n
03:30  \r\n
\r\n
MODE 01:0000  72\r\n
72:P92  \r\n
01:0003  \r\n
02:0180  10\r\n
03:30  \r\n
\r\n


For 10 minute cycling, change the second parameter from 180 to 10.
100: Needs, Site base, Tue 07-Jan-2003 12:01:48 MST, other needs
needle-nose pliers
drill index!
non-leaky alcohol bottle
clear-all button in cockpit (to clear all the "lost data")
new fan power connectors for 25m
truck credit card




99: Needs, Site base, Tue 07-Jan-2003 07:49:42 MST, print needs color cartridge
The printer is complaining about its color cartridge, and I couldn't get it
to print last night.

98: Weather, Site all stns, Tue 07-Jan-2003 07:42:15 MST, sunrise
sunrise 2 minutes ago ~7:40


97: Weather, Site all stns, Tue 07-Jan-2003 07:40:19 MST, fog
Fog now mostly clear, but was thicker earlier -- especially at low levels.
Still thin wisps at bottom ~3m.  Does this explain our krypton problems?
(but what is happening at 2m.tower?)


96: Status, Site tower, Mon 06-Jan-2003 18:29:31 MST, visit tasks
Here is an attempt to document what we did today [still a frustrating day]:

Conditions: snow deep (~40cm) some places, bare others, grass clumps poking
through.  Completely clear today until evening and still only a few wisps
to the EAST.  (We noted sunset at ~16:43, with no orographic clouds to the 
west. NO WIND!

We left the base for the tower about 9:30.

- Rebooted cosmos, which had crashed at 14:50 on 1/4 (2 days ago) for some
unknown reason.  Matrix LED was dull RED, Datel LED was off, Ironics LED was
1. and green.
- Installed grease [but data from Tdome still look strange]
- Installed Hi-cal regulator on CO2 system
	- Hi-cal is ttl 0, Lo-cal is ttl 1
	- Reset system flow rate (inside pump box) from 6 to 7 lpm
	- Set Hi-cal to -2 lpm
	- Reset Lo-cal to -2 lpm (was almost 0 after the above adjustment)
	- Cylinder values: Lo-cal 575psi, Hi-cal 1800psi, both output pressures
	~15 psi
	- Visually inspected (most of) inlet line -- no problems noted
- Swapped cables at bare (both 1.5 and 4m).  Noted that fan was working to
keep dome clear of icing that was present on the white shield, so didn't swap
fan.  Data don't look any different from before.  [We noted later that the
shadow of the main tower hits the bare site at noon, so dips due to 
short-wave radiation are expected.]
- Remounted snow depth gauge to center of grass darkhorse (to avoid having
it view the darkhorse legs).  Data are still very noisy.  This sensor is not
suited for this purpose!
- Started work on the top of the tower:
- Swapped T/RH fan at 30m.  Data still quite different.  We'll probably have
to swap the entire sensor:(
- Cleaned krypton at 30m
- Swapped krytpon at 20m -- now SN1390 with a 2.0cm path.
- While attempting to clean krypton at 10m (and investigate the funny tilt
observed in the data), we broke the weld of the boom-tower clamp.  It had
been rotated up and clearly was the source of the tilt.  Brought boom into
tower and laid it on its side.
- Cleaned krypton at 2m.

We returned to the base from about 13:00-14:00.

- Back out to work on 25m.  Pulled trolley in; removed top pygs; installed new
plate; reinstalled pygs; removed level; cleaned all pygs;
remounted level on new plate; tested
leveling [level reads as -.030, -.100]; ran trolley out - DOESN'T CLEAR;
brought trolley back; filed away edges of white shields [this took a while to
get right]; ran trolley out [FINALLY WORKS!]  However, the data still are
strange.  We'll have to do more tests (and may bring back a pyg or two) 
tomorrow.  I think the serial numbers are:
in.a 31975; in.b 31979; out.a 31974; out.b 31977 [Gordon says that these
correspond to Steve's cal-lab files]
- Figured out a fix to 10m -- remounted T/RH at 0.5 underneath 1m sonic boom
using hose clamps [now it is ~3cm higher]; stole the boom-tower clamp it was
using and remounted the 10m sonic; cleaned 10m krypton at this time.  Note that
knowlege of the boom angle is now lost at this site, since the side guy wires
were loose once it was remounted.
- Tried to shoot boom angles with the DataScope.  The angles definitely are
different, but the DataScope doesn't read accurate at large pitch angles and
you need to be pretty close to see the alignment.  We will have to do
theodolite shoots on everything.
- Checked guy wire tension (see updated old entry).
- Checked TP01 wiring at grass by "beeping".  The following table is from 
memory:
	Pin	Color	CR10
	1	Green	4H
	2	Black	5L
	3	White	5H
	4	Black	4L
	7	Black	Gnd
	8	Red	Switched 12V
	9 	Clear	Shield
	 	Red	12 V control to P8?
I think that these are correct, but the probe still doesn't work.
- Checked 2m krypton heater prototype.  Control voltage works, but no heat or
air flow.  Removed assembly for trailer testing [showed that pump is now dead,
but we melted the heaters demonstrating this!]

We got back to the base about 18:00.

- we still have a resident mouse

95: Status, Site all stns, Fri 03-Jan-2003 15:33:48 MST, Daily Status Jan 3
Pam power: OK.  

Sonics: OK.

Krytons: OK. No diurnal variation.

H2O-kh2o: 20m still shows a diurnal difference (light leak?).  This didn't
change with cleaning on about 12/17.

Tv and P: tc.2m.sage(2) drops below -8 deg C twice around 1900 MST (1/2/02) while others measure between -2 and -6 deg.  Pressure at tower still has temperature contamination.

Sonic variance 2m: OK

Scalar variance 2m: tc'tc'2m at sage(2) spikes up to 2.8 ~1900 MST (1/2/03) while other sensors measure <.5

Scalar variance: OK
 
Fluxes at 2m: OK

Flux profiles: OK

CO2: OK.  Has been working since our visit on 12/18.  However, only the 
"low-cal" (tty 1) is connected due to a bad regulator. co2mr.2m quickly dropped from ~257 to ~222, jumped back up, then dove to ~231 all within the time frame of about an hour and a half ((0203 - 0400 MST 1/2/03).

CO2 Moments: shows problems with data during same time frame. 

TRH: OK, though the RH profiles are not monotonic.  

Short wave radiation: OK.  

Long Wave Radiation: OK

|Tdome-Tcase|: in.4m.bare has about the same amount of variation  but values are slightly higher.

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to air.
I'm just ignoring this, since we should have adequate Tsoil sampling.
All Tsoils.tower(grass) have a bit noise from the radiometer fans, but this is
small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more digging
up snow to check.  Currently, I'm ignoring this as well since there probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage on 
12/18.  One of us probably cut a cable.  This should be fixed next visit.

Soil Thermal properties: All of these signals look bad, despite "fixing" the
wiring problem at tower (grass).  I've expected that it will take a while to
get good data from these, since there was a small air gap around the probes
after installation; however, the data look even worse.  In particular, each
"event" on Vheat should look like a square wave toggling between 0 and 1 V.
The "Vpile" signal at tower looks close to the signal expected for Vheat, but
at the wrong amplitude.

Wind direction profile: OK, though there probably are biases of 1-2 degrees
(despite aligning the booms by eye up and down the tower and shooting boom
angles that were mostly the same).  We'll need to check these angles again.



94: CO2, Site tower, Fri 03-Jan-2003 09:19:13 MST, Calibration gas regulator installation
How to install the regulator on the caltank and adjust the cal-gas flow

Insert the regulator male fitting into the cylinder female screw thread.
Hand tighten the notched nut. REMEMBER THAT IT IS A LEFT-HANDED THREAD
Use the wrench from the pump box to tighten the nut, really tight.
Unscrew the regulator adjust until loose.
Close the regulator outlet valve.
Open the cylinder valve.
Note the cylinder pressure. Should be 1500 psi or so
Screw in the regulator adjust until the second stage reads 5 psi or so.

Connect, finger tight, the floating ball guage to intake
Monitor the intake flow of the CO2 intake. It should be 6-7 lpm
If it is too low close down the aspirator pump bypass valve
If it is too high open up the aspirator pump bypass valve

ASSUME THAT LO-CAL IS THE EXISTING CONNECTED CYLINDER
Command the Lo-cal gas solenoid to open; ttl_on cosmos 0  CHECK THIS
The Lo-cal gas flow of ~ 7-8 lpm should overflow out the intake at ~ 1-2 lpm
Reverse the flow connection of the floating ball guage
If overflow is too low increase the regulator adjust
If overflow is too high decrease the regulator adjust
Command the Lo-cal gas solenoid to close; ttl_off cosmos 0  CHECK THIS 

ASSUME THAT HI-CAL IS THE NEWLY CONNECTED CYLINDER
Command the Hi-cal gas solenoid to open; ttl_on cosmos 1  CHECK THIS
Open the regulator outlet valve
The second stage pressure will fall, screw in the regulator adjust  to ~ 5-10 psi
The Hi-cal gas flow of ~ 7-8 lpm should overflow out the intake at ~ 1-2 lpm
If overflow is too low increase the regulator adjust
If overflow is too high decrease the regulator adjust
Command the Hi-cal gas solenoid to close; ttl_off cosmos 0  CHECK THIS

Return Lo-cal solenoid to microprocessor control; ttl_reset cosmos0
Return Hi-cal solenoid to microprocessor control; ttl_reset cosmos1



93: Status, Site all stns, Thu 02-Jan-2003 18:58:44 MST, Daily Status Jan 2
Pam power: OK.  

Sonics: OK.

Krytons: Heating of 2m.tower is off.  The heating of other krytons still doesn't seem to be working.

H2O-kh2o: 20m still shows a diurnal difference (light leak?).  This didn't
change with cleaning on about 12/17.

Tv and P: Pressure at tower still has temperature contamination.

Sonic variance 2m: OK

Scalar variance 2m: Doesn't look good. kh2o'kh20'.2m for all sites shoots up to 4+ at various times.

Scalar variance: 10m and 30m seem noisy between 1800 MST on 1/1 and 0000 MST 1/2  
Fluxes at 2m: OK

Flux profiles: OK

CO2: OK.  Has been working since our visit on 12/18.  However, only the 
"low-cal" (tty 1) is connected due to a bad regulator.  We note that the
pressure drop down the line is almost exactly 10% (P.licor ~660, P ~730),
which is a reasonable value.

TRH: OK, though the RH profiles are not monotonic.  

Short wave radiation: OK.  

Long Wave Radiation: Rlw.in.4m.bare still looks bad to a greater or lesser
degree. 

|Tdome-Tcase|: in.4m.bare all seems likes it's off, it has much greater variation then other sensors. 

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to air.
I'm just ignoring this, since we should have adequate Tsoil sampling.
All Tsoils.tower(grass) have a bit noise from the radiometer fans, but this is
small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more digging
up snow to check.  Currently, I'm ignoring this as well since there probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage on 
12/18.  One of us probably cut a cable.  This should be fixed next visit.

Soil Thermal properties: All of these signals look bad, despite "fixing" the
wiring problem at tower (grass).  I've expected that it will take a while to
get good data from these, since there was a small air gap around the probes
after installation; however, the data look even worse.  In particular, each
"event" on Vheat should look like a square wave toggling between 0 and 1 V.
The "Vpile" signal at tower looks close to the signal expected for Vheat, but
at the wrong amplitude.

Wind direction profile: OK, though there probably are biases of 1-2 degrees
(despite aligning the booms by eye up and down the tower and shooting boom
angles that were mostly the same).  We'll need to check these angles again.

92: Radiometers, Site all stns, Tue 31-Dec-2002 13:38:02 MST, logger files
Here are all of the coeffient files loaded into the loggers.
(I got these from the cal-lab PC, so Steve S should verify that they
are the correct files.)

25m:
31975	pi1	06	251.193168
31974	po1	07	257.334020
31979	pi2	08	263.157895
31977	po2	09	261.711594
PI1TC-A0	x	10	2500
A1	x	11	3.355702
A2	x	12	.279378
A3	x	13	3.970700
A4	x	14	200.967000
PI1TD-A0	x	15	2500
A1	x	16	3.054151
A2	x	17	.270750
A3	x	18	3.531936
A4	x	19	172.539000
PO1TC-A0	x	20	2500
A1	x	21	3.354452
A2	x	22	.2795544
A3	x	23	3.870320
A4	x	24	213.662200
PO1TD-A0	x	25	2500
A1	x	26	3.053245
A2	x	27	.2706839
A3	x	28	3.462000
A4	x	29	179.035300
PI2TC-A0	x	30	2500
A1	x	31	3.355302
A2	x	32	.2796132
A3	x	33	3.907393
A4	x	34	205.226200
P12TD-A0	x	35	2500
A1	x	36	3.055463
A2	x	37	.271608
A3	x	38	3.538717
A4	x	39	170.026300
PO2TC-A0	x	40	2500
A1	x	41	3.355048
A2	x	42	.2793915
A3	x	43	3.969477
A4	x	44	201.022200
PO2TD-A0	x	45	2500
A1	x	46	3.052511
A2	x	47	.2706538
A3	x	48	3.516783
A4	x	49	176.020500
2018-A0	level-x	50	1.219419589
A1	x	51	10.90459926
A2	x	52	0.098392445
A3	x	53	-0.007021535
A4	x	54	0.000000
A0	level-y	55	-0.3726148
A1	x	56	10.4045871
A2	x	57	-0.1230141
A3	x	58	0.511059
A4	x	59	0.000000


4m:
31981	pi1	06	267.1653754
31980	po1	07	277.0466824
31976	p01m	08	255.102041
PI1TC-A0	x	10	2500
A1	x	11	3.354002
A2	x	12	.27995
A3	x	13	3.89791
A4	x	14	222.6
PI1TD-A0	x	15	2500
A1	x	16	3.050935
A2	x	17	.271316
A3	x	18	3.382922
A4	x	19	190.3757
PO1TC-A0	x	20	2500
A1	x	21	3.354756
A2	x	22	.279427
A3	x	23	3.948847
A4	x	24	215.9851
PO1TD-A0	x	25	2500
A1	x	26	3.052869
A2	x	27	.271025
A3	x	28	3.449757
A4	x	29	189.982
PO1mTC-A0	x	30	2500
A1	x	31	3.354748
A2	x	32	.279406
A3	x	33	3.926898
A4	x	34	202.8886
PO1mTD-A0	x	35	2500
A1	x	36	3.053435
A2	x	37	.270616
A3	x	38	3.557806
A4	x	39	167.4648

Grass (tower):
Q7-97069	NET(A+)	00	9.27
x	A-	01	11.48
P26416	x	02	290.866783
P29260	x	03	308.261406
K940181	x	04	82.3723229
K970379	x	05	45.51661356
H993563	x	06	38.9
H993562	x	07	37.3
H87034	x	08	36.8
Tc-A0	x	10	2500
A1	x	11	3.354472
A2	x	12	.2797812
A3	x	13	3.912974
A4	x	14	215.5208
Tdin-A0	x	15	2500
A1	x	16	3.052399
A2	x	17	.270981
A3	x	18	3.408667
A4	x	19	188.8687
Tcot-A0	x	20	2500
A1	x	21	3.353832
A2	x	22	.278595
A3	x	23	4.02801
A4	x	24	180.4266
Tdot-A0	x	25	2500
A1	x	26	3.053253
A2	x	27	.2706527
A3	x	28	3.530291
A4	x	29	166.282
HUX	x	50	180
baud	x	99	0

Greasewood:
31978	pygout	00	283.527077
PINTC-A0	x	10	2500
A1	x	11	3.354223
A2	x	12	.27981
A3	x	13	3.963779
A4	x	14	206.898
PINTD-A0	x	15	2500
A1	x	16	3.052301
A2	x	17	.270621
A3	x	18	3.499677
A4	x	19	177.414

Lake:
Q7-94197	NET(A+)	00	9.0
x	A-	01	11.22
P29137	x	02	240.211386
P29136	x	03	234.741784
K970377	x	04	43.55400697
K940186	x	05	82.71298594
H983213	x	06	38.2
H983211 x	07	36.4
H943149	x	08	39.7
Tc-A0	x	10	2500
A1	x	11	3.350057
A2	x	12	.279229
A3	x	13	3.147866
A4	x	14	-193.854
Tdin-A0	x	15	2500
A1	x	16	3.354167
A2	x	17	.278817
A3	x	18	3.653292
A4	x	19	222.2646
Tcot-A0	x	20	2500
A1	x	21	3.354344
A2	x	22	.278532
A3	x	23	3.845153
A4	x	24	185.129
Tdot-A0	x	25	2500
A1	x	26	3.354233
A2	x	27	.278537
A3	x	28	3.802989
A4	x	29	184.2122
HUX	x	50	180
baud	x	99	1

Sage:
Q7-94916	NET(A+)	00	9.21
x	A-	01	11.54
P26417	x	02	212.68
P27907	x	03	286.53
K970378	x	04	50.02501251
K940187	x	05	101.4198783
H983212	x	06	38.4
H993560	x	07	37.7
H943147	x	08	42.1
Tc-A0	x	10	2500
A1	x	11	3.355204
A2	x	12	.2795166
A3	x	13	3.950246
A4	x	14	209.3092
Tdin-A0	x	15	2500
A1	x	16	3.354969
A2	x	17	.2793838
A3	x	18	3.85506
A4	x	19	212.913
Tcot-A0	x	20	2500
A1	x	21	3.353713
A2	x	22	.2787871
A3	x	23	3.778048
A4	x	24	177.421
Tdot-A0	x	25	2500
A1	x	26	3.354732
A2	x	27	.2787514
A3	x	28	3.651471
A4	x	29	190.8095
HUX	x	50	180
baud	x	99	1








91: Sonics, Site all stns, Sun 29-Dec-2002 22:05:31 MST, sonic tilts
Attached are tilt plots over the last month.  All show small biases (3cm max?),
and reasonable angles (1.3deg max?).  There are some bad periods (icing?) in
the upper sonics.  More serious is that the 10m level had a bias shift sometime.
I'll have to investigate this more.

90: Snow_Precip, Site tower, Fri 27-Dec-2002 16:51:36 MST, no rain tips from Nov 20 to Dec 16
There were no rain tips between Nov 20 and Dec 16, even during
the snowstorm of Nov 23/24.  

There is data from the grass logger over this time.
Rnet is the next field after rainr, and we have complete Rnet coverage.

date/time MST   tips
Nov 20 13:32:30  61
Nov 20 13:37:30 189
...zilch...
Dec 16 01:07:30 1

After Dec 16, sporatic tips are seen - which looks reasonable -
showing an accumulation of 5.3 mm between Dec 16 and Dec 25.


I re-ran covar, with rainr by itself in a covar group, so it does not
appear to be a software issue.




89: Status, Site tower, Fri 27-Dec-2002 16:27:46 MST, tower stuff
I never did log the events of the afternoon of 12/18 at the end of our last
visit:

CO2: It took us a while to diagnose a cabling problem between the adam ttl
control and Tony's CO2 box.  (They weren't wired 1-to-1, though the 
interconnecting cable was after its last repair!)  We exchanged the power and
ground pins inside the CO2 box which got things working.  Tony adjusted flows.
However, one of the regulators was bad, so we only left lo-cal working.  I
note that, with the current wiring, lo-cal is controlled as ttl channel 1
(not 0).

Thermal properties: We found a deviation in the CR10 wiring from what we
expected, so Cody switched wires to 4H and 5H at grass.  The data show that 
something is still wrong.

Grease: We installed the darkhorse, leveled everything, and had it running,
but the data only showed good Tdome values.  Thus, I pulled the radiometer
and cable and brought it back to Boulder.  (This was easier than disconnecting
the cable!)  It turned out to be swapped pins in the new cable we had made.
We'll reinstall this when we go back up.


88: Status, Site all stns, Fri 27-Dec-2002 15:53:27 MST, Daily Status Dec 27
None of us have been systematically looking at data over X-mas, but here are
the things Gordon and I have observed:

--------------------------------

Pam power: OK.  The (new) lake wind generator is producing enough to keep
the batteries happy, especially with the high winds we have been getting.
Iload still is different between lake and sage -- cause still unknown.

Sonics: OK!

Krytons: Heating of 2m.tower is now off.  The heating didn't seem to be as
effective as it should have (perhaps the tubes became misaligned?) and Gordon
may have melted it by now by overheating to compensate.

H2O-kh2o: 20m still shows a diurnal difference (light leak?).  This didn't
change with cleaning on about 12/17.

Tv and P: Pressure at tower still has temperature contamination.

Sonic variance 2m: OK

Scalar variance: 10m and 30m seem noisy.  2m tower has a low level of noise 
(non-zero variance).  

Fluxes at 2m: OK

Flux profiles: OK

CO2: OK.  Has been working since our visit on 12/18.  However, only the 
"low-cal" (tty 1) is connected due to a bad regulator.  We note that the
pressure drop down the line is almost exactly 10% (P.licor ~660, P ~730),
which is a reasonable value.

TRH: OK, though the RH profiles are not monotonic.  This could be fetch 
related, but it is more likely that there are up to 2% biases in the sensors.
I'm also not completely convinced that there aren't some small biases on
the T values, though the data have looked pretty good lately.

Short wave radiation: OK.  We've noticed a sharp drop in incoming in the
evenings.  We suspect that this is caused by the elevated horizon of the
mountains, though it could also be effected by a lee-wave cloud caused by
the mountains.  It would be fun to plot the horizon (using the sun_angle from
the various days).  More effective probably would be to log the conditions
and time when the sun disappears during the next visit.

Long Wave Radiation: Rlw.in.4m.bare still looks bad to a greater or lesser
degree.  I've thought this was snow build-up due to a weak fan, but may be
wrong.  Grease is still not installed (but now ready to go on the next trip).

|Tdome-Tcase|:  25m.in.a, 25m.out.b, and bare in.4m.bare all seem worse than
they should be (up to 0.5C).

Tsoil:  OK. Tsoil.2.5cm.a.sage still looks like it may be exposed to air.
I'm just ignoring this, since we should have adequate Tsoil sampling.
All Tsoils.tower(grass) have a bit noise from the radiometer fans, but this is
small enough that I'm ignoring it too.

Qsoil:  Qsoil.c.10cm lake sensor still down.  It might also a cable cut
(like the Gsoil.10cm.a problem found at lake), but will require more digging
up snow to check.  Currently, I'm ignoring this as well since there probably
is enough redundancy in the Qsoil measurements.

Gsoil: Gsoil.c.10cm sage became bad as Tony and I were working at sage on 
12/18.  One of us probably cut a cable.  This should be fixed next visit.

Soil Thermal properties: All of these signals look bad, despite "fixing" the
wiring problem at tower (grass).  I've expected that it will take a while to
get good data from these, since there was a small air gap around the probes
after installation; however, the data look even worse.  In particular, each
"event" on Vheat should look like a square wave toggling between 0 and 1 V.
The "Vpile" signal at tower looks close to the signal expected for Vheat, but
at the wrong amplitude.

Wind direction profile: OK, though there probably are biases of 1-2 degrees
(despite aligning the booms by eye up and down the tower and shooting boom
angles that were mostly the same).  We'll need to check these angles again.







87: Kryptons, Site tower, Tue 24-Dec-2002 22:25:47 MST, stopped krypton_heat script
The 2 meter krypton heater was turned on for much of last night
but is didn't have any effect - apparently it has burned out.

SteveO said it wasn't designed to be run for lengthy periods - 
perhaps the resistor strip melted the hose.

I have stopped the krypton_heat script.  

86: Kryptons, Site tower, Sun 22-Dec-2002 22:40:06 MST, krypton heater at 2m on tower
kh2oV at the 2 meter kryptons started falling at about 18:00 MST tonight,
Dec 22, showing the signature of a frost event.

The krypton_heat script worked correctly and turned on the heater
control line for 2 minutes starting at 19:35:15 MST.  The voltage
kept dropping and so every 20 minutes it turned  the heater on again.

However I can see no effect in the data - the voltage kept falling.
At 22:24:30 I turned it on manually, until 22:37:15 when the script
turned it off.  No effect was seen in the data.  10 second averages 
of kh2oV.2m stayed around 0.0035 V and didn't show any upward trend.

At 22:51:00 MST I turned it the heater line on again. The krypton
voltage is still bottomed out at 0.003 V.

At 23:00 MST I changed the script to leave the heater on until
the voltage goes above 0.05 V.  Each level is checked every
5 minutes.

FYI: To create 10 second averages of kh2oV.2m
	prep -r 0.1 -D kh2oV.2m

The history of heater control is saved in files on 
$ASTER/projects/FLOSSII/cal_files/kh2oV.*

It is also written to syslog, and so it is in the isff logs.

Added, morning of Dec 23:  It took 40 minutes to bring kh2oV from
.004 at 23:10 MST to .05 V at 23:50 last night.


85: Status, Site all stns, Thu 19-Dec-2002 14:10:16 MST, Daily Status Dec 19
 Pam power:
  Vbatt: lake down to ~12.2 V overnight.

Sonics: OK.

Krytons: spike of kh20.30m 20m and 10m at 0500 MST.

H2O-kh2o: spike in the data (basically all sensors, but largest at 30m (-2.7)) at  0500 MST

Tv and P: Pressure at tower is bad.

Sonic variance 2m: OK

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

TRH: OK

Long Wave Radiation: Rlw.in.4m.bare stn0 still looks bad. Should be checked.

|Tdome-Tcase| : Sensors at bare in.4m.bare stn0 and out.grease stn0 are misbehaving. 

Tsoil:  Sensor error at soil.2.5cm.  Others look good.

Qsoil:  Qsoil.c.10cm lake sensor still down.  Need to check.

Gsoil: Gsoil.c.10cm sage(2) isn't acting right. Need to check.





84: Soil, Site lake, Thu 19-Dec-2002 11:56:01 MST, Corrected therm prop. timing at lake
The logger at lake was looping through the thermal properties code every
3 minutes, instead of every 3 hours.  The code was changed to 3 hour
looping at 11:30am, Dec 19


83: Radiometers, Site tower, Thu 19-Dec-2002 11:54:00 MST, Checked logger coefs at 25m
We're seeing wrong-looking Tdome-Tcase on in.b.25m and out.a.25m

Displayed the logger coefs at 25m.  They all match SteveS's sheet.
He will check his files.
82: Status, Site all stns, Thu 19-Dec-2002 11:51:40 MST, Daily Status Dec 18
 Pam power:
  Vbatt: lake down to ~12.2 V overnight.

Sonics: The 2m sonic QC (last plot) and the Profile sonic QC (last plot)
look different than I've seen before.  Values are below 100% and there
appear to be a few data gaps.  Please take a look at.

Krytons: OK.

H2O-kh2o: sensor at 2m.lake is OK, others may not be.  Between 0900 and
1100 MST there were downward spikes and then they suddenly jumped up ~2
g/m3 (or more)

Tv and P: Pressure at tower is bad.

Sonic variance 2m: Large spike (~5.8 (m/s)2) in u'u.2m lake on Dec 17
at 2050 MST

Scalar variance: kh20'kh20'.2m lake had spikes of 2.7 and 3.8 at about
Dec 17 0100 MST.  kh20'kh20'.2m tower had spikes of  2 and 3.9 close to
1300 MST.

Fluxes at 2m: OK

Flux profiles: OK

TRH: OK

Long Wave Radiation: Rlw.in.4m.bare stn0 looks bad on three spots. It
drops significantly on Dec 17, 0300MST and again between Dec 17 2300-Dec
18 0200.

|Tdome-Tcase| : OK. Sensors at bare and grease, I assume, are still be
worked on.

Tsoil:  Sensor error at soil.2.5cm.  Others look good.

Qsoil:  Qsoil.c.10cm lake sensor still down.  Need to check.

Gsoil: Gsoil.c.10cm sage(2) and Gsoil.a.10cm lake(3) aren't acting
right. Need to check.




81: Status, Site all stns, Thu 19-Dec-2002 11:51:08 MST, Daily status Dec 17
Pam power:
  Vbatt: lake down to ~12.5 V overnight.

Sonics: data missing from tower site between 1300-1500 - ASTER reboot

Krytons: Heaters appear to be working well.   kh2o.2m lake had unusually
large spike (~11.5 g/m3) at  1300 MST.

H2O-kh2o: sensor at 2m.lake is acting up. It's reading more moist than
the other sensors, and has a few spikes.  One in particular reaches
about -7.6 g/m3.  Might want to check out.

Tv and P: Pressure at tower is bad.

Scalar variance: OK

Fluxes at 2m: OK

Flux profiles: OK

TRH: OK

Radiation: Rlw.in at sage and lake the same.

|Tdome-Tcase| : OK

Tsoil:  Sensor error at soil.2.5cm.  Others look good.

Qsoil:  Qsoil.c.10cm lake sensor still down.  Need to check.

Gsoil: OK

80: Soil, Site all stns, Wed 18-Dec-2002 14:13:04 MST, Soil gravimetric/bulk density
The Sage, Lake and Grass C soil samples were not treated the same as were 
the A and B. Rather, the cores were extracted in toto and will be used for 
bulk density determination. Because of the limitation of the balance the cores
were broken in two. The core portions were put into larger Al container, were
weighed and then left to dry until the next visit.
In all cases the ~20 cm of snow was removed and the ice crust atop the soil 
surface was broken and removed. At Grass dead grass stalks were also removed.

Sage C            Tare               Fresh(18 Dec)           Dry(1/7/03)
0 - 8cm  ZI        8.0               143.2		     122.0
         ZII       8.0               149.0		     115.2

Lake C            Tare               Fresh(18 Dec)           Dry(1/7/03)
0 - 8 cm ZIII      8.0               120.5		     87.0
         ZIV       8.0               168.8		     127.2

Grass C           Tare               Fresh(18 Dec)           Dry(1/7/03)
0 - 8 cm ZV        8.0                76.2  top 4 cm peaty   34.3
         ZVI       8.0               136.1  bottom 4 cm      106.4


79: Status, Site sage, Wed 18-Dec-2002 11:42:12 MST, sage visit
Visited sage from ~0945-1100:
- turned on rad fans with cap
- cleaned rads
- did pyg shadow tests
- cleaned krypton
- installed TP01
- tested e-fence (showed 1000V, but not a good ground connection)
- removed SDR disk 1, inserted disk 5.
- Q1&Q2 are installed on EVE front panel

78: Status, Site lake, Wed 18-Dec-2002 11:39:27 MST, Lake visit
Visited lake from ~0800-0945.
- staked down new solar panel
- spliced break in hft.a cable (but just by twisting wires and with e-tape,
  since I couldn't get the butane soldering iron to stay on
- all data now look okay on EVE console
Cody went back and taped up the hft cables below the soil mux box.

77: Radiometers, Site tower, Wed 18-Dec-2002 10:08:57 MST, Boom PYG mount Modification
When initially we tried to extend the trolly out on the
boom we had the problem that we could not extend it all the way out because
the radiometer housings were too close together for the locking tab to fit
through them at the end.  To fix this problem we tried unscrewing all but
one of the four screws holding in the radiometer housings on each of
the top two radiometers.  We then rotated them outward and attached the
housings to the trolly using panduit ties going from the unsecured screws
to the braces of the trolly.  This held well, but when we tried to extend
the trolly once more, the fan on the left or southern side of the trolly hit
the corresponding guide wire.  We left the trolly for the night and tried
again to fix it this morning (12/18/2002).  We changed the position of the
problem radiometer by rotating it differently so that it could get by the
guide wire and was also far enough from the center so that the locking
tab could fit through.  We also replaced the panduit ties holding the
radiometer down with some wire so that we didn't have to worry about
the panduit ties breaking.  When we now attempted to send out the trolly
we were successull in reaching the end of the boom.  We are still having
some problems with the locking mechanism grabbing onto the trolly, but that
problem was not one that could be fixed without taking the entire boom
down and readjusting the lock.
76: Snow_Precip, Site all stns, Tue 17-Dec-2002 17:36:16 MST, snowing today
About 2 today, it started snowing.  I would guess that we got at least 5cm by 
5pm.

75: Eve, Site , Tue 17-Dec-2002 17:30:54 MST, EVE Q1&Q2
Both spare EVE boxes in the trailer have Q1 & Q2 installed.  I still haven't
checked sage.

74: Logistics, Site tower, Tue 17-Dec-2002 17:04:58 MST, electric fence ok
We see 5000V at the main tower electric fence.

73: Power, Site sage, Tue 17-Dec-2002 17:03:18 MST, 4th solar panel installed
I just installed a 4th solar panel at sage and put the 3rd panel in parallel
with the 2 main panels, as I did at lake yesterday.

72: Radiometers, Site tower, Tue 17-Dec-2002 16:58:14 MST, 4m rads up
~1330? the 4m rads were up and connected.  I leveled and cleaned them.  Fan
is on to the top housing.

We still haven't plugged in the 1.5m bare due to a lack of a cable.

71: Radiometers, Site tower, Tue 17-Dec-2002 16:37:04 MST, Installation of boom pyrgeometers
From 8:00 - 10:30 hauled up the boom trolley, the big white box and the cabling.
The pyrg leveling platform fitted on the trolley cross-wise. Even so the trolley
could not pass the latching device as the pygs in their housing were too close
together. 
From 14:30 - 16:60 modified the mounting of the pyrgs in their housings to yield
a greater separation. The level mounting was also modified. The south pyrg fan
then would not clear the lower stabilizing guy.

An ad hock arrangement will be contrived.

For a later time: Bring a drill and bits, centerpunch and tap to make new 
mounting holes.This can be done on the tower.

70: Soil, Site all stns, Tue 17-Dec-2002 16:21:07 MST, Gravimetric soil samples
Weighed the soil samples collected on Mon 16 Dec (see log entry 67)

11:40 Sage site. .At B, dug away 10 cm of snow. There was a 2 -3 cm ice crust 
on the soil surface. Very hard going to sink the core as the ground was frozen.
Although the snow was completely removed from the surface before coring the ice 
incorporated in the soil melted in the base to yield mud.
The corer stem broke and had to be replaced with a pice of pipe.

                   Fresh        Dry (18 Dec)
   0-3 cm  AI      25.3         17.6
           AII     20.9         14.7
           AIII    21.9         16.1

   3-8 cm  BI      26.1         21.6
           BII     22.6         18.8
           BIII    22.3         18.8

13:30 Lake site At B dug down through 30 cm of snow and brushed soil surface
clear of snow
                   Fresh        Dry (18 Dec)
  0-3 cm  CI       13.3         11.5
          CII      16.6         14.4
          CIII     14.6         12.5

  3-8 cm  DI       13.9         11.3
          DII      15.7         12.7
          DIII     19.1         15.3

14:00 Grass site. At B 25 - 30 cm of snow. Dug down and then broke up crust
incorporating dead grass stems and ice. Cleaned this away and started core
at peaty surface.
                   Fresh       Dry (18 Dec)
  0-3 cm  EI        8.2         2.5
          EII       7.4         2.4
          EIII      6.7         2.4

  3-8 cm  FI       15.9        12.3
          FII      17.3        12.7
          FIII     19.7        15.3




69: Temp_RH, Site tower, Mon 16-Dec-2002 18:31:26 MST, T/RH swapped back
We swapped the T/RH at 2m on the tower back with the sensor that was pulled
by Kurt 2 weeks ago, since it was reading about 0.2C high.  (The one we just
installed has a new fan (which is quieter than the one we just removed.)

This done about 11:00.

68: CO2, Site tower, Mon 16-Dec-2002 18:13:33 MST, CO2 setup
Rekeyed the two ADAMs at tower at 14;45

Set up the CO2 system

See doc for diagram of system
Checked all swageloks, tightened regulators but did not open cylinder valves.
Powered CO2 box, switched on Licor Green light on later
Powered up pump, monitored the intake flow = 6 lpm
Control cable from serial connector assosiated with the new fiber board circuit
CO2 to 101
(Did not yet connect T to 102 nor P to 103 as lacking the analog cables)
67: Soil, Site all stns, Mon 16-Dec-2002 17:55:08 MST, Gravimetric soil samples
Gravimetric soil samples are to be:
                                   top  3 cm
                                   next 5 cm
The plan is to do Sage A, Lake A and Grass A on Mon, 16 Dec, dry overnight
 ......................B.......B           B    Tue  17..................
.......................C ......C.......... C....Wed..18    leave in oven

Also to take total cores from three sites for bulk density and leave in oven

11:20 Sage site. Approx 10 cm of snow between bushes with almost no snow under some bushes.At A, under a bush with no snow cover, core sample.

                   Fresh        Dry (17 Dec)
   0-3 cm  AI      16.6         11.7
           AII     12.4          8.5
           AIII    13.1          8.8

   3-8 cm  BI      22.7         20.6
           BII     23.7         21.6
           BIII    24.2         21.9

12:00 Lake site Fairly uniform 20 -25 cm snow. Couldnt locate A so went approx 
the correct distance from box. Dug down through snow to soil and brushed the surface clear.
                   Fresh        Dry (17 Dec)
  0-3 cm  CI       15.6         11.0
          CII      14.5         10.4
          CIII     17.1         12.5

  3-8 cm  DI       23.4         21.5
          DII      25.8         23.8
          DIII     25.4         23.2

12:40 Grass site. Grass sticking out of 0-8 cm snow.
Note that the top 1 cm was grass stalks. I discarded this and separated the dirtr component from the next 3 cm. The lower 5 cm was mostly dirt.
                   Fresh       Dry (17 Dec)
  0-3 cm  EI       5.5          2.9
          EII      4.6          2.4
          EIII     5.0          2.7

  3-8 cm  FI       9.5          6.8
          FII     11.3          8.3
          FIII    12.7          9.2

66: Status, Site lake, Mon 16-Dec-2002 17:51:13 MST, stuff
Several other tasks at lake:
- capacitor added to fan (only upward PSP) and fan turned on ~16:30
- radiometers cleaned
- Rnet releveled (off a bit)
- Krypton cleaned
- Q1 and Q2 are installed in EVE
- Electric fence shows 5000V
- Freewave Tx power set to 5 (I'm not sure if this was what is was or not)
- SDR was on, seemed to be working, seemed quiet, #2



65: Power, Site lake, Mon 16-Dec-2002 17:48:00 MST, Wind generator swapped
We tried playing with the pot on the old wind generator (clockwise, counter-
clockwise, mid point) but were unable to get the prop to turn much.  Thus,
we replaced the entire unit with a new one (405, rather than 403?).  This
done from about 14:30-16:30.

Also, we put an extra solar panel for the electric fence and connected the
old 3rd panel with the others for more station power.  We still need to
stake this down.

64: Kryptons, Site tower, Mon 16-Dec-2002 17:46:35 MST, krypton heater added
We mounted the heater on the krypton at about 13:00.  It was connected through
cosmos about 16:00 (Tony may elaborate).  Thus, we should have some data on
this system tonight.

63: Base, Site base, Mon 16-Dec-2002 17:43:45 MST, aster rebooted
When testing the krypton heater script, a member of our software team 
accidentally killed several aster processes about 20:10Z.  ASTER was rebooted
here about 21:30Z and the adams rebooted about 30min later.

62: Eve, Site lake, Mon 16-Dec-2002 17:42:10 MST, EVE rebooted
EVE could not talk to /card or /ramdisk.  After talking to John, I pulled the
PCMCIA module from the backplane and rebooted.  The station appears to be fine
now.

61: Soil, Site lake, Mon 16-Dec-2002 17:40:33 MST, TP01 installed
TP01 SN 200235 installed at lake site ~1645.  It is installed about at 5cm in
the frozen sand.  It still has an air gap, though somewhat neater (and thus
smaller) than at the tower site -- I'm learning.





60: Soil, Site tower, Mon 16-Dec-2002 17:34:16 MST, TP01 installed
TP01 SN 200234 installed at tower site ~1345.  (It was connected an hour
earlier, but it was on top of the snow.)  It is installed about at 5cm in
the (peat-like) dirt.  I'm sure that there is an air-gap beside it, but 
hopefully this will heal with time.


59: Status, Site all stns, Mon 16-Dec-2002 14:40:53 MST, Daily status, Dec 16

Pam Power:
  Vbatt: lake down to ~12.3 V overnight.

Sonics: OK.

krytons: air at 30m sensor more moist then at other levels by ~.5 g/m^3. Power glitch apparent where data is missing.

H2O-kh2o: H2O-kh2o.20m has drastic changes compared with other measurements. 

Tv and P: Pressure is incorrect at tower.  

Scalar variance: OK. 

Fluxes a 2m : OK.

Flux profiles: OK.

TRH: OK

Radiation: Rlw.in at sage and lake the same.
      

|Tdome-Tcase|: OK.(Tdome-Tcase) at lake temporarily removed. 

Tsoil:  other than soil.2.5cm sensor problems, looks good.

Qsoil: Qsoil.c.10cm lake sensor still down. Need to check on next trip to site.
       
Gsoil:  spikes and glitches in Gsoil.c.10cm.lake





















58: Base, Site tower, Mon 16-Dec-2002 13:21:58 MST, krypton heater script
Commands for krypton heaters on the tower:
  start_krypton_heat    starts script which controls krypton heaters on tower
  stop_krypton_heat     stops script


The krypton_heat script runs itself every 5 minutes.  It cycles
through the 4 kryptons on the tower, and therefore will visit
each kryton every 20 minutes.

krypton_heat gets the latest 5 minute value of kh2oV.2m, kh2oV.10m, 
kh2oV.20m or kh2oV.30m, using ncprint on the netcdf file. If
the value is less than 0.05 V, then the appropriate ttl_on
command is sent to the adam, and after 2 minutes, ttl_off
is sent.

The adam name, the ttl port of the krypton heater, the trigger voltage,
and the  amount of time to leave the heater on is set in the script,
for each kh2oV variable.  If these are changed, do a a stop/start
to restart the script:

	stop_krypton_heat
	start_krypton_heat

57: Adams, Site tower, Mon 16-Dec-2002 12:18:20 MST, Power glitch caused adam crash
Daisy and cosmos crashed last night due to a power glich. Here's
what the UPS reported in the system log.  These times are MST:

Dec 16 00:59:37 aster upsd[1546]: UPS on battery
Dec 16 00:59:37 aster upsd[1546]: Normal power restored: UPS on line
Dec 16 00:59:40 aster upsd[1546]: UPS on battery: Deep momentary sag 108.5 V
Dec 16 00:59:40 aster upsd[1546]: Normal power restored: UPS on line

56: Status, Site all stns, Fri 13-Dec-2002 14:10:10 MST, Daily status, Dec 13

Pam Power:
  Vbatt: lake down to ~12.1 V overnight.

Sonics: OK. v'v'2m spiked to almost 6 m/s^2 at 2230 MST on Dec 12.

krytons: air at 30m sensor more moist then at other levels by ~1 g/m^3.

Tv and P: Pressure is incorrect at tower.  

Scalar variance: OK. 

Fluxes a 2m : w'tc'.2m went bad just after 0300 MST, but came back around 0430 MST. w'h2o'.2m spiked to .09 a little after 0400 MST.

Flux profiles: OK.

TRH: OK

Radiation: No Rlw.in at sage?
      

|Tdome-Tcase|: OK.(Tdome-Tcase).in lake still bad. 

Tsoil:  other than soil.2.5cm sensor problems, looks good.

Qsoil: Qsoil.c.10cm lake sensor still down. Need to check on next trip to site.
       
Gsoil:  spikes and glitches in Gsoil.c.10cm.lake


















55: Radiometers, Site lake, Thu 12-Dec-2002 18:12:19 MST, Added dat function to correct Rlw.in at lake

Since Rlw.in at lake has not been correct since the beginning
of the project (due to a bad Tcase.in) a special version of
dat.Rlw.in has been written which just sets Rlw.in at lake equal
to Rlw.in at sage.  

Using the corrected Rlw.in, then Rsum compares well with Rnet at lake,
so I don't believe Rnet needs to be replaced.


54: Soil, Site lake, Thu 12-Dec-2002 16:46:33 MST, Changed Tsoil sampling at sage,lake

Dec 12, 16:45 MST

As in logbook entry #51 at the grass site, changed the CR10x logger code 
at lake and sage so that there is a 10 msec delay before sampling
the Tsoil probes.




53: Status, Site all stns, Thu 12-Dec-2002 14:13:44 MST, Daily status, Dec 12

Pam Power:
  Vbatt: lake down to ~12.1 V overnight.

Sonics: w.2m spiked to -.9 at about 0330 MST.  Judging from 2m sonic, Profile Sonic QC, and Sonic Anemometer uvw plots anemometer may be having a bit of trouble in cold temperatures. 

krytons: for all sensors, except a 30m, voltage dropped to almost 0 over night.

Tv and P: Pressure is incorrect at tower.  

Scalar variance: kh2o'kh2o' tower went a bit crazy between 0300 and 0700 to. 

Fluxes a 2m : w'tc'.2m went bad just after 0300 MST, but came back around 0430 MST. w'h2o'.2m spiked to .09 a little after 0400 MST.

Flux profiles: OK.

TRH: OK

Radiation: Looks like Rlw.in.tower(1) and Rlw.out.tower(1) are missing a couple chunks of data.
      

|Tdome-Tcase|: missing chunk of data from tower between 1500 and 1700 MST in Dec 11.  Gordon was fixing things which caused the data gap.
 

Tsoil:  other than soil.2.5cm sensor problems, looks good.

Qsoil: Qsoil.c.10cm lake sensor still down. Need to check on next trip to site.
       
Gsoil:  spikes and glitches in Gsoil.c.10cm.lake

















52: Status, Site all stns, Wed 11-Dec-2002 14:44:12 MST, Daily status, Dec 11

Pam Power:
  Vbatt: lake down to ~12.1 V overnight.

Sonics: OK

krytons: OK

Tv and P: Pressure is off at tower.  

Scalar variance: kh2o'kh2o' lake spiked to .24 (g/m^3)^2. Krytons show lake site had the most moisture over night.

Fluxes a 2m : OK

Flux profiles: Spike in w'h2o'.10 m stn0 of .18 a little after 0700 MST.

TRH: OK

Radiation:
     Absence of pergeometer at lake site causes spikes, and missing values. It affects Rsum lake and the long wave incoming lake site trace. 

|Tdome-Tcase|: lake plot also affected by removal of pergeometer. 
 

Tsoil:  Sensors at the tower are noisy because fans are running but problem should be fixed (see "Modified grass logger Tsoil sampling" log entry).  soil.2.5cm sage still having problems.

Qsoil: Qsoil.c.10cm lake sensor still down. Need to check on next trip to site.
       
Gsoil:  spikes and glitches in Gsoil.c.10cm.lake














51: Soil, Site tower, Wed 11-Dec-2002 10:27:54 MST, Modifed grass logger Tsoil sampling
Dec 11, 10:27 MST:

Added 10 millisecond delay to P11 instruction (#50) which is used to 
sample T107 Tsoil probes. This is the 03 parameter, which we 
changed from 1 to 21.

Lab testing showed that this reduces
noise, and so we hope this will reduce the noise on Tsoil that
we see when the radiometer fans are on.

This changed reduced the noise significantly, except on the 10cm
probes, which are sampled at another point in the CR10 code.


Dec 11, 15:13 MST

Changed the P11 instruction for the 10cm probes too (instruction 60). 
Here is the  rserial output:

MODE 01:0000  60\r\n
60:P11  \r\n
01:01  \r\n
02:05  \r\n
03:01  \r\n
04:0027--\r\n
05:+1.0000 \r\n
06:+0.0000 \r\n
MODE 01:0000  60\r\n
60:P11  \r\n
01:01  \r\n
02:05  \r\n
03:01  21\r\n
04:0027--\r\n
05:+1.0000 \r\n
MODE 01:0000  60\r\n
60:P11  \r\n
01:01  \r\n
02:05  \r\n
03:21  \r\n
04:0027--\r\n
05:+1.0000 \r\n
06:+0.0000 \0x7f\r\n


Oops: looks like I didn't restart the logger with a "*0" and it didn't 
resume until I rserial'd to it again at about 16:20.






50: Status, Site all stns, Tue 10-Dec-2002 14:58:17 MST, Daily status, Dec 10

Pam Power:
  Vbatt: lake down to ~12.1 V overnight.
  Wind speeds at sage (lwind) still have not been high enough to trigger the current. Steve recommends checking/adjusting the pot on next visit.


Sonics: OK
MST.

krytons: Noraml diurnal variation in all sensors except those at 20 and 30 m.

Tv and P: Pressure is off at tower.  

Scalar variance: kh2o'kh20'.2m tower shot up to 4 (g/m^3)^2 at 0700 MST.  It recovered at 0900 MST. Most likely a result of frost in the kryton. 

Fluxes a 2m : spikes in the data were seen for both w'kh2o (.058) and w'tc (.28)
 at 0330 MST.

Flux profiles: OK

TRH: OK

Radiation:
     Absence of pergeometer at lake site causes spikes, and missing values. It affects Rsum lake and the long wave incoming lake site trace. 

|Tdome-Tcase|: lake plot also affected by removal of pergeometer. 
 

Tsoil:  Sensors at the tower are noisy because fans are running (they need to be turned off). Tsoil.2.5cm sage still having problems.

Qsoil: Qsoil.c.10cm lake sensor still down. Need to check on next trip to site.
       
Gsoil:  spikes and glitches in Gsoil.c.10cm.lake












49: Status, Site all stns, Mon 09-Dec-2002 10:15:41 MST, Daily status, Dec 9

Pam Power:
  Vbatt: lake down to ~12.2 V overnight.
  lwind sage dropped off after 2200 MST Dec 6, and has not been seen since. Wind  speeds have not been high enough to trigger the current. Steve recommends 
  checking/adjusting the pot on next visit.


Sonics:  small glitch in w.2m at sage, dropped to -.20 at little after 5:00 am 
MST.

krytons: voltage of sensor at kh2o.2m tower recorded negative values between about 0500-0900 MST 12/8, and after 2200 MST 12/8 until 0900 MST the morning of the 9th. Heaters?

Tv and P: Pressure is off at tower.  

Scalar variance: kh2o'kh20'.2m tower shot up to 4 (g/m^3)^2 after 2000 MST 12/8.  It recovered at about 0900 MST 12/9. Probabley frost in the kryton. 

Flux profiles: w'kh2o.30m stn0 spiked to .24 at 0800 MST.

TRH: OK

Radiation:
     Absence of pergeometer at lake site causes spikes, and missing values. It affects Rsum lake and the long wave incoming lake site trace. 

|Tdome-Tcase|: lake plot also affected by removal of pergeometer. 
 

Tsoil:  Sensors at the tower are noisy because fans are running (they need to be turned off). Tsoil.2.5cm sage still having problems.

Qsoil: Qsoil.c.10cm lake sensor still down. Need to check on next trip to site.
       
Gsoil:  spikes and glitches in Gsoil.c.10cm.lake

RH: For first two plots, second sensor above ground, between 1:00-12:00, reads more moist (~5%) than the other in all but two of the profiles.










48: Base, Site base, Mon 09-Dec-2002 09:17:56 MST, replaced disk drive
replaced disk drive, and returned it to Gordon on
Wed, Dec 4th
 
47: Radiometers, Site lake, Mon 09-Dec-2002 09:15:02 MST, removed uplooking pir sensor
removed uplooking PIR sensor from the lake site
on Wed, Dec 4th
46: Soil, Site tower, Mon 09-Dec-2002 09:09:34 MST, soil profile array cable installed
the soil profile array interface cable was installed on
Wed, Dec 4th
we are now receiving data
45: Temp_RH, Site tower, Mon 09-Dec-2002 08:38:50 MST, 2m TRH
replaced 2m TRH (bad fan on original) Wed, Dec 4th


44: Barometer, Site tower, Fri 06-Dec-2002 17:35:47 MST, Pressure error
We've seen this problem before (CASES99/OASIS98): a temperature error on the 
ASTER barometers (but not PAM).  Attached is a plot which shows the difference
between the ASTER and PAM barometers on the top panel.  The bottom panel is the
same after correcting using:

x <- dat("P")                            
x[,1] <- x[,1] + (1/7)*dat("T")[,1] - 0.8

(i.e., a 1mb change per 7degC change and a small bias).

In the future (even now??), we really should use the PAM barometers for ASTER 
-- 5mb is getting to be a large error.


43: Status, Site all stns, Fri 06-Dec-2002 13:43:50 MST, Daily status, Dec 6

Pam Power:
  Vbatt: lake down to 12.2 V overnight


Sonics:  OK

krytons: No diurnal variation. kh20.30m is again reading more moist than the other sensors, by about 1 g/m^3.

Scalar variance: OK.  

TRH: OK

Radiation:
     Absence of pergeometer at lake site causes spikes, and missing values. It affects Rsum lake and the long wave incoming trace. 

|Tdome-Tcase|: lake plot also affected by removal of pergeometer. 
 

Tsoil:  Sensors at the tower are noisy because fans are running (they need to be turned off). Tsoil.2.5cm sage still having problems.

Qsoil: Qsoil.c.10cm lake sensor still down. Need to check on next trip to site.
       
Gsoil:  spikes and glitches in Gsoil.c.10cm.lake

 







42: Status, Site all stns, Thu 05-Dec-2002 13:15:31 MST, Daily status, Dec 5

Pam Power:
  Vbatt: lake down to 12.2 V overnight


Sonics:  OK

krytons: All sensors show diurnal variation, but to a smaller degree than shown
in previous nights with variation. kh20.30m is reading slightly more moist than the other sensors.

Scalar variance: tc'tc'.2mlake shot up to 1.9 (degC)^2 at about 9 am.  

TRH: OK

Radiation:
     Instrument (pergeometer?) was removed caused spikes in places, and missing values in others. It affected net radiation plot and the long wave incoming trace. 

|Tdome-Tcase|: also affected by removal of instrument from lake site. 
 

Tsoil:  Sensors at the tower site are up and running. They are noisy because fans are running (they need to be turned off).

Qsoil: Qsoil.c.10cm lake sensor died. Need to check on next trip to site.
       Sensors at tower are up and running.

Gsoil:  spikes and glitches in Gsoil.c.10cm.lake

 






41: Status, Site all stns, Thu 05-Dec-2002 09:32:59 MST, Daily status, Dec 4

Pam Power:
  Vbatt: lake down to 12.3 V overnight


Sonics:  OK

krytons: sensors at 10, 20, and 30 m showed no diurnal variation over night.  The moisture at 2 m dropped significantly at around midnight.

Scalar variance: OK.  

TRH: 2m sensor was replaced.

Radiation:
     |Tdome-Tcase| problem at lake should be fixed as of today.


Tsoil:  Sensor at 2.5 cm may be exposed, tracking T-air rather than T-soil. 
Gsoil:  spikes and glitches in Gsoil.c.10cm.lake

1/2 hr ave. w'tc': small variations at 15 m (.007), and at 20m (~.14) at 1:00 MST noted. 






40: Status, Site all stns, Tue 03-Dec-2002 14:24:34 MST, Daily status, Dec 3

Pam Power:
  Vbatt: lake down to 12.4 V overnight


Sonics:  OK

krytons: kh20V 2m tower voltage dropped below 0.  Other sensors read at 0 as a resulst of frost/condensation in the sensor, this is the first time I've seen negative voltage values.  It recovered at around 9 am.

Scalar variance: kh20'kh2o' 2m tower shot up to four over night as a result of the kryton getting wet, or the presence of frost.  

TRH: 2m on tower is an outlier.  Should be replaced. New 5 minute T profile 
shows sensor error very well.

Radiation:
  Rlw.in.lake is bad. Tdome-Tcase.in.lake averaged 3 degC all night.  
  |Tdome-Tcase| for other radiometers is less than 0.5 degC at night.


Tsoil:  Sensor at 2.5 cm may be exposed, tracking T-air rather than T-soil. 
Gsoil:  spikes and glitches in Gsoil.c.10cm.lake







39: Status, Site all stns, Mon 02-Dec-2002 15:37:21 MST, Daily status, Dec 2

Pam Power:
  Vbatt: lake down to 12.2 V overnight


Sonics:  OK
krytons: OK. Plot of kh20 at 30m was added. Is reading a little high (~4.5-5.0) compared with others (~3.5-4.0). I will keep an eye on it.

Scalar variance: tc'tc'.2m sage was acting up a bit after 9am, Gordon says it appears to be ok.

TRH: 2m on tower is an outlier.  Should be replaced.

Radiation:
  Rlw.in.lake is bad. Tdome-Tcase.in.lake averaged 3 degC all night.  
  |Tdome-Tcase| for other radiometers is less than 0.4 degC at night.


Tsoil:  Sensor at 2.5 cm may be exposed, tracking T-air rather than T-soil. 
Gsoil:  spikes and glitches in Gsoil.c.10cm.lake

Gordon added profiles of T, Q, Wind spd, w'tc', w'kh20', and u*, to the web site.




38: Radiometers, Site lake, Sun 01-Dec-2002 07:48:56 MST, Tcase&Tdome.in coefs at lake
Fixing the coefficients for Tcase&Tdome.in at lake (logbook #32)
did not improve the data (perhaps it made it worse - haven't checked).

Tdome-Tcase.in at lake is still about 3 degC at night, whereas 
it is usually less than 0.5 degC for all other pyrgeometers, including 
the down-looker at lake.

Here are the current coefficients from the logger at lake from doing
an EVE talk with the logger.  They agree with logbook entry #14



>
MODE 04:00  
00:+9.0000 
01:+11.220 
02:+240.21 
03:+234.74 
04:+43.554 
05:+82.713 
06:+38.200 
07:+36.400 
08:+39.700 
10:+2500.0 
11:+3.3542 
12:+.27882 
13:+3.6533 
14:+222.26 
15:+2500.0 
16:+3.3500 
17:+.27923 
18:+3.1479 
19:-193.85 
20:+2500.0 
21:+3.3543 
22:+.27853 
23:+3.8451 
24:+185.13 
25:+2500.0 
26:+3.3542 
27:+.27854 
28:+3.8030 
29:+184.21 
50:+180.00 
99:+1.0000 
MODE 00
LOG 12
*


37: Camera, Site tower, Wed 27-Nov-2002 11:25:54 MST, New settings for camera
Changed these settings today (Nov 27 on the camera to try to improve the
pictures.


Image settings:
  Brightness: 11
  Color balance: Automatic fast
  Color Level: 15


Nov 30, 10:00 pm MST
  Images were looking a bit saturated
  Color level: 14
  Color balance: Automatic wide 

Dec 2, 14:30 pm MST
  Color balance: automatic normal
  Color level: 13
  
Dec 3, 12:00 noon
  Color balance: automatic normal  
  Color level: 10

Dec 17, 9:30 am
  Shutter speed 1/120
  Brightness: 8
  Color balance: automatic wide
  Color level 6



36: Camera, Site tower, Wed 27-Nov-2002 11:24:14 MST, Camera white balance documentation
From the help page for the Axis camera:

White Balance

White balance can be described as the reference color against which all other
colors in an image are compared. The resident white balancing system in the 
AXIS 2420 automatically detects white colors in any image and uses these
intelligently as a reference for other colors. From the drop-down list,
select a suitable balance setting for your camera's local environment.
The available white balance options are described below:

    * Automatic Normal:  Appropriate across the entire color bandwidth, this
automatic setting causes the AXIS 2420 to search for a white reference area
within the image, and all other colors are then adjusted proportionally.
Generally suitable for most environments filled with both artificial white
light, and sunlight.
    * Automatic Narrow: This setting causes the AXIS 2420 to search for a
white reference area in the image using narrow limits of the color range.
All other colors are adjusted proportionally. The automatic color adjustment
is turned off if an appropriate white reference area can not be identified.
    * Automatic Wide: This setting causes the AXIS 2420 to search for a
white reference area in the image, using no limits at all.  All other colors
are adjusted proportionally. This setting gives natural colors in most
environments.
      Note: In environments that do not contain white areas, another color
will be defined as white; i.e. a red object that covers the entire image
will appear as gray.
    * Automatic Fast: The same as Automatic Wide (described above),
but with faster adjustment time.
    * Indoor Light Bulbs: Fixed color adjustment; ideal for a room with
some artificial light other than fluorescent lighting; for example, areas
illuminated using standard light bulbs. Good for normal light temperatures
up to 3200K.
    * Indoor Fluorescent 1: Fixed color adjustment; good for fluorescent
tube light with color temperatures of 4200K
    * Indoor Fluorescent 2: Fixed color adjustment; good for fluorescent
tube light with color temperatures between of 4700K
    * Outdoor: General fixed color adjustment that is appropriate for
outdoor applications under good skylight (color temperatures of 6300K).

Note: In certain situations, the white balancing system will not operate
effectively. Problems generally occur if the image does not contain any
white color at all, or if the dominant color is something other than white.
In these circumstances, the AXIS 2420 can erroneously base the white balance
on another visible color in the image, and consequently, the colors may be
distorted: a pale background picture with foreground objects of an acute
reddish or blue hue are very symptomatic of this condition, and in such
cases we recommend that you operate the AXIS 2420 in a fixed white balance
mode.


35: Status, Site all stns, Tue 26-Nov-2002 16:35:36 MST, Daily status, Nov 26

Pam Power:
  Vbatt: lake down to 12.2 V overnight


Sonics:  OK
krytons: OK, presence of frost at night obvious.

TRH: 2m on tower is an outlier.  Should be replaced.

Radiation:
  Rlw.in.lake is still bad, caused by bad Tdome-Tcase.in.lake
  Tdome-Tcase.in.lake averaged 2 degC all night.  
  |Tdome-Tcase| for other radiometers is less than 0.4 degC at night.


Tsoil:  Sensor at 2.5 cm may be exposed, tracking T-air rather than T-soil. 
Gsoil:  spikes and glitches in Gsoil.c.10cm.lake


34: Camera, Site tower, Tue 26-Nov-2002 11:16:16 MST, color balance on axis camera

Using a color balance of "outdoor" on the axis camera gives us
green snow.  Yesterday, Nov 25, at about 4 pm I set the color balance to
"automatic normal".  Things look a bit better, I'll wait a bit
before trying others, like automatic narrow and automatic wide.
33: Status, Site all stns, Mon 25-Nov-2002 11:28:32 MST, Daily status, Nov 25

Pam Power:
  Vbatt: lake down to 12.2 V overnight


Sonics:  OK
krytons: OK

TRH: 2m on tower is an outlier.  Should be replaced.

Radiation:
  Rlw.in.lake is bad, probably caused by bad Tdome-Tcase.in.lake
  Tdome-Tcase.in.lake was around 2 degC all night.  
  |Tdome-Tcase| for other radiometers is less than 0.4 degC at night.


Gsoil:  spikes in Gsoil.c.10cm.lake
32: Radiometers, Site lake, Fri 22-Nov-2002 09:41:11 MST, Fix lake coefficients for pygin
A difference plot showed questionable data from the dome and case temperatures
for Lake pygin. It was discovered that the coefficients had been reversed.
Corrections have been made both at the PAM site and the logbook entry 14.

31: Base, Site base, Thu 21-Nov-2002 13:57:57 MST, Swapping USB data disks
We have two 120Gb Maxtor USB disks, labled USB#1 and USB#2 that we
alternate on the base system. They provide a backup copy of the high 
rate data and a way to to shuttle the data back to Boulder.

Here is how to swap the USB disks on aster.  You do not need
to reboot to do this.


0. Check that the red light on the front of the disk is not lit. 
   For a minute or two at about 10:15 am and 6:15 pm MST, files 
   are copied to it, so you may have to wait a bit.

1. From the aster account (does not have to be root), enter umount
   command in a shell window:

   umount /mnt/usbhd

   If the disk is in use, you will get an error to that effect. Wait
   a minute, then try again.

2. Wait until red light on front of drive stops flashing.

3. Unplug silver USB cable from back of drive

4. Unplug power from drive.
 
5. Plug power into second drive. Wait 5 seconds or so for it to spin up.

6. Plug USB cable into second drive.

7. Enter mount command:   
   mount /mnt/usbhd

You're done.


On Nov 21, the drive labeled USB#2 is in use and USB#1 was taken back
to Boulder.





30: Snow_Precip, Site tower, Thu 21-Nov-2002 13:10:02 MST, Snow depth gauge

Snow depth gauge is  1.75 m above the grass, which agrees with
the sensor output.
29: Radiometers, Site lake, Thu 21-Nov-2002 12:35:58 MST, Rnet at lake

SteveS checked serial numbers at lake, all OK.

He noticed water droplets inside net radiometer dome, perhaps
that explains the Rnet data?
28: Radiometers, Site sage, Thu 21-Nov-2002 12:31:47 MST, PSPs swapped at sage
The PSPs at sage were actually placed as follows:

  940187 is PSPIN,
  970378 is PSPOUT

which is the reverse of what the logger program expected.

The logger program was changed.
27: Kryptons, Site all stns, Thu 21-Nov-2002 10:04:49 MST, Krypton Placement
11/21/02
Krypton placement with respect to CSAT:
Horizontal orientation on all installations:

		From center of CSAT array:
Stn2 (sage)	26cm back	20cm below
Stn3 (lake)	26cm back	20cm below


Tower:
2m   27cm
10m  24cm
20m  24cm
30m  24cm


26: Status, Site all stns, Thu 21-Nov-2002 09:56:35 MST, Stn2 / Stn3 Rad Fans / Icing
11/20/02
Turned off fans at Lake:  Noise on Tsoil data eliminated through
			morning.

11/21/02-am
Lake 9:20 local. (no fans)
	Fairly heavy ice/rime on 2 up looking rads and rnet
	A few crystals also on csat but the Krypton appeared clean.
Sage 9:30 local (with fans)
	Heavy ice/rime on rsw.in and rnet.  The rsw.in appeared more crusty
	than at Lake, which may be perhaps worse than simple crystals.
	The rlw.in appeared clean.
	The rsw/rlw.out had a small amount of crystals forming on them.
	Csat had quite a bit of crystals on lower transducers.  Gordon reports
	minimal flagging of the data and it overall looks good.
25: Sonics, Site tower, Thu 21-Nov-2002 09:42:42 MST, Sonic boom angles
Boom angles at tower, as shot with the theodilite:
    d   m  s
1m  88 30 20
2m  88 30 20
5m  88 30 20
10m 88 30 20
15m 88 30 20
20m 88 58 05
30m 89 24 00

		DataScope	Compass
Stn2 (Sage) 	  78.8		  79
Stn3 (Lake)	  79.0		  80-
(0-declination set in compass/datascope)

Added, Dec 26, by GDM:
Using a declination of 11.5 deg east, the values for sage and lake become:
Stn2 (Sage)       90.3
Stn3 (Lake)       90.5

To get our normal boom azimuth angle, add 180.



24: Sonics, Site sage, Thu 21-Nov-2002 08:49:43 MST, PAM Stn2 Reconfiguration
11/18/02
CSAT3 s/n 536 changed from 20hz straight to 20hz oversampling mode.

Swapped TRH: new=101  removed=1
Swapped Krypton: new=1395  removed=1133
Swapped Electronics box new=2  removed=6
New config: Freewave s/n2105 fastout fmt0141.dat
Swapped Logger, installed 4-component Rad, config file FLOSSlog.dat


23: Sonics, Site lake, Thu 21-Nov-2002 08:43:20 MST, PAM Stn3 Reconfiguration
11/19/02
CSAT3 s/n 537 changed from 20hz straight to 20hz oversampling mode.

11/18/02
Swapped TRH: new=502  removed=5
Swapped Krypton: new=1397  removed=1393
Swapped Electronics box new=9  removed=7
New config: Freewave s/n2088 fastout fmt0141.dat
Swapped Logger, installed 4-component Rad, config file FLOSSlog.dat

22: Radiometers, Site tower, Wed 20-Nov-2002 13:24:51 MST, rad stand up
Lynette and I just installed the 4-comp rad on the grass rad stand with its
logger.  Data should be coming in now.   I'm going back out to install the
rain gauge cable.  The soil sensors are not coming in because we don't think
that we have the MUX cable for it here:(

21: Temp_RH, Site tower, Wed 20-Nov-2002 09:38:03 MST, TRH heights on tower
TRH heights in meters on tower, as measured by SteveS.


 0.5
 1.0
 1.95
 5.0
10.0
15.5
20.25
25.2
30.2






20: Status, Site all stns, Mon 18-Nov-2002 21:10:33 MST, November setup
Nov 18:
  SteveS, Kurt, John, Kyle and Gordon on site.

  PAM stations (sage, lake) reconfigured to send fast data
  over freewaves, and no slow data over GOES.

  Base up, with GPS, UPS, and PPP to Earthlink configured.
  Copying IP address, logbook and netcdf files to Boulder.

  Power laid to main tower, bottom two sonic booms on.

Nov 19:
  Pam stations in good shape - need notes from John.

  Tower:
    Adams up, breezecom's working.
    All TRHs on main tower, but 50Ys covered and without aspiration horns
    Sonics at 1 and 2m, krypton at 2m
    Other sonic booms on tower.
    no radiation
    no snow depth
    webcam operational

19: Radiometers, Site tower, Fri 15-Nov-2002 14:26:23 MST, Greasewood 1.5m radiation CR10X code
NOTE: comments in this code may not be up to date


;{CR10X}
; CASES99 - 1.5 meter radiation system at greasewood site
; This is a special program for the FLOSS-II project
; The logger handles 1 pygreometer at the 1.5m greasewood site.
;
; CR10 CHANNEL ASSIGNMENTS
; ANALOG CHANNELS
;        1H - comm 1H of AM416
;        1L - comm 1L of AM416
;        2H - comm 2H of AM416
;        2L - comm 2L of AM416
;
; CONTROL PORTS
;        C1 - reset signal to AM416
;        C2 - clock to AM416
;
; AM416 MULTIPLEXER CHANNELS
;        2-1H     +  PYGOUT
;        2-1L     -
;
;        2-2H     +  PYGOUT Tcase
;        2-2L     -
;
;        3-2H     + PYGOUT Tdome
;        3-2L     -
;
; OUTPUT MESSAGE FORMAT
;       LOGGER ID
;       PYGOUT
;       PYGOUT Tcase
;       PYGOUT Tdome
;       
;
; EEROM  SPACE USED FOR CALIBRATION COEFFICIENTS
; the coefficients below are examples!

;location      value        Serial Number   description
;  0           255.10       xxxxxF3         Epply pyrgeometer, in_1 W/m^2/mV
;  10          2500                         PYGOUT Tcase,  Vref mV
;  11          .00335595                    A0
;  12          .277486                      A1
;  13          4.8993                       A2
;  14          10.4799                      A3
;  15          2500                         PYGOUT Tdome, Vref mV
;  16          .00303674                    A0
;  17          .289526                      A1
;  18          -3.0453                      A2
;  19          841.413                      A3
;


*Table 1 Program
  01: 1.0       Execution Interval (seconds)

1:  Z=F (P30)
 1: 1        F
 2: 00       Exponent of 10
 3: 10       Z Loc [ LoggerID  ]

;  Start collecting data from the multiplexer

2:  Do (P86)
 1: 41       Set Port 1 High

;  loop thru 416mux to Epply Pile

3:  Do (P86)
 1: 9        Call Subroutine 9

4:  Beginning of Loop (P87)
 1: 0        Delay
 2: 2        Loop Count

     5:  Do (P86)
      1: 9        Call Subroutine 9

;  get pyg and pyr radiation data
     6:  Volt (Diff) (P2)
      1: 1        Reps
      2: 3         25 mV Slow Range
      3: 1        DIFF Channel
      4: 1     -- Loc [ pygin_raw ]
      5: 1.0      Mult
      6: 0.0      Offset

;  get Epply temperature data
     7:  Volt (Diff) (P2)
      1: 1        Reps
      2: 5         2500 mV Slow Range
      3: 2        DIFF Channel
      4: 12    -- Loc [ PYGOUT_Tc ]
      5: 1.0      Mult
      6: 0.0      Offset

8:  End (P95)

9:  Do (P86)
 1: 51       Set Port 1 Low

;  Compute some real values, pile first

10:  Z=X*F (P37)
 1: 1        X Loc [ pygin_raw ]
 2: 0.0      F @@0
 3: 11       Z Loc [ PYGOUT    ]

; now lets do the temperatures

;  convert raw PYGOUT Tdome

11:  Z=F (P30)
 1: 0        F @@10
 2: 00       Exponent of 10
 3: 20       Z Loc [ V_Vx      ]

12:  Z=X/Y (P38)
 1: 20       X Loc [ V_Vx      ]
 2: 12       Y Loc [ PYGOUT_Tc ]
 3: 20       Z Loc [ V_Vx      ]

13:  Do (P86)
 1: 2        Call Subroutine 2

14:  Polynomial (P55)
 1: 1        Reps
 2: 20       X Loc [ V_Vx      ]
 3: 20       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@12
 6: 0        C2 @@13
 7: 0        C3 @@14
 8: 0.0      C4
 9: 0.0      C5

15:  Z=F (P30)
 1: 0.0      F @@11
 2: -3       Exponent of 10
 3: 21       Z Loc [ B0        ]

16:  Z=X+Y (P33)
 1: 20       X Loc [ V_Vx      ]
 2: 21       Y Loc [ B0        ]
 3: 20       Z Loc [ V_Vx      ]

17:  Do (P86)
 1: 3        Call Subroutine 3

18:  Z=X (P31)
 1: 20       X Loc [ V_Vx      ]
 2: 12       Z Loc [ PYGOUT_Tc ]

;  convert raw PYGOUT Tdome

19:  Z=F (P30)
 1: 0        F @@15
 2: 00       Exponent of 10
 3: 20       Z Loc [ V_Vx      ]

20:  Z=X/Y (P38)
 1: 20       X Loc [ V_Vx      ]
 2: 13       Y Loc [ PYGOUT_Td ]
 3: 20       Z Loc [ V_Vx      ]

21:  Do (P86)
 1: 2        Call Subroutine 2

22:  Polynomial (P55)
 1: 1        Reps
 2: 20       X Loc [ V_Vx      ]
 3: 20       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@17
 6: 0        C2 @@18
 7: 0        C3 @@19
 8: 0.0      C4
 9: 0.0      C5

23:  Z=F (P30)
 1: 0.0      F @@16
 2: -3       Exponent of 10
 3: 21       Z Loc [ B0        ]

24:  Z=X+Y (P33)
 1: 20       X Loc [ V_Vx      ]
 2: 21       Y Loc [ B0        ]
 3: 20       Z Loc [ V_Vx      ]

25:  Do (P86)
 1: 3        Call Subroutine 3

26:  Z=X (P31)
 1: 20       X Loc [ V_Vx      ]
 2: 13       Z Loc [ PYGOUT_Td ]

;  Is it time to send data out?

27:  Time (P18)
 1: 0        Seconds into current minute (maximum 60)
 2: 5        Mod/By
 3: 22       Loc [ TimeHack  ]

28:  IF (X<=>F) (P89)
 1: 22       X Loc [ TimeHack  ]
 2: 4        <
 3: 1        F
 4: 10       Set Output Flag High

29:  Sample (P70)
 1: 1        Reps
 2: 10       Loc [ LoggerID  ]

30:  Resolution (P78)
 1: 1        High Resolution

31:  Average (P71)
 1: 3        Reps
 2: 11       Loc [ PYGOUT    ]

32:  Serial Out (P96)
 1: 52       Printer Comma/9600 Baud



*Table 2 Program
  02: 0.0000    Execution Interval (seconds)

*Table 3 Subroutines
1:  Beginning of Subroutine (P85)
 1: 2        Subroutine 2

;  compute x = ln[(V/Vx) - 1] / 1000

2:  Z=X+F (P34)
 1: 20       X Loc [ V_Vx      ]
 2: -1       F
 3: 20       Z Loc [ V_Vx      ]

3:  Z=LN(X) (P40)
 1: 20       X Loc [ V_Vx      ]
 2: 20       Z Loc [ V_Vx      ]

4:  Z=X*F (P37)
 1: 20       X Loc [ V_Vx      ]
 2: .001     F
 3: 20       Z Loc [ V_Vx      ]

5:  End (P95)

6:  Beginning of Subroutine (P85)
 1: 3        Subroutine 3

;  finish T (case/dome) computation

7:  Z=1/X (P42)
 1: 20       X Loc [ V_Vx      ]
 2: 20       Z Loc [ V_Vx      ]

8:  Z=X+F (P34)
 1: 20       X Loc [ V_Vx      ]
 2: -273.15  F
 3: 20       Z Loc [ V_Vx      ]

9:  End (P95)

10:  Beginning of Subroutine (P85)
 1: 9        Subroutine 9

; pulse the AM416 multiplexer to the next bank of channels
     11:  Do (P86)
      1: 72       Pulse Port 2

     12:  Excitation with Delay (P22)
      1: 1        Ex Channel
      2: 0        Delay W/Ex (units = 0.01 sec)
      3: 1        Delay After Ex (units = 0.01 sec)
      4: 0        mV Excitation

13:  End (P95)


End Program

-Input Locations-
1 pygin_raw 1 1 1
2 _________ 0 0 0
3 _________ 0 0 0
4 _________ 0 0 0
5 _________ 0 0 0
6 _________ 0 0 0
7 _________ 0 0 0
8 _________ 0 0 0
9 _________ 0 0 0
10 LoggerID  1 1 1
11 PYGOUT    5 1 1
12 PYGOUT_Tc 9 1 2
13 PYGOUT_Td 17 2 1
14 _________ 1 0 0
15 _________ 0 0 0
16 _________ 0 0 0
17 _________ 0 0 0
18 _________ 0 0 0
19 _________ 1 0 0
20 V_Vx      1 14 13
21 B0        1 3 2
22 TimeHack  1 2 1
23 _________ 1 1 0
24 _________ 1 1 0
25 _________ 0 0 0
26 _________ 0 0 0
27 _________ 0 0 0
28 _________ 0 0 0
-Program Security-
0000
0000
0000
-Mode 4-
0
10
11
12
13
14
15
16
17
18
19
-Final Storage Area 2-
0
-CR10X ID-
0
-CR10X Power Up-
3


18: Radiometers, Site tower, Fri 15-Nov-2002 14:24:46 MST, 4m radiation CR10X code
NOTE: comments in this code may not be up to date

;{CR10X}
; CASES99 - 4 and 1.5 meter radiation system
; This is a special program for the FLOS-II project
; The logger handles 3 pyrgeometers. 2 are located
; at 4m and the third at 1.5m
;
; CR10 CHANNEL ASSIGNMENTS
; ANALOG CHANNELS
;        1H - comm 1H of AM416
;        1L - comm 1L of AM416
;        2H - comm 2H of AM416
;        2L - comm 2L of AM416
;
;        4H       +  PYGOUT_1
;        4L       -
;
;        5H       +  PYGOUT_1 Tcase
;        5L       -
;
;        6H       +  PYGOUT_1 Tdome
;        6L       -
;
; CONTROL PORTS
;        C1 - reset signal to AM416
;        C2 - clock to AM416
;
; AM416 MULTIPLEXER CHANNELS
;        2-1H     +  PYGIN_4
;        2-1L     -
;
;        3-1H     + PYGOUT_4
;        3-1L     -
;
;        2-2H     +  PYGIN_4 Tcase
;        2-2L     -
;
;        3-2H     + PYGIN_4 Tdome
;        3-2L     -
;
;        4-2H     + PYGOUT_4 Tcase
;        4-2L     -
;
;        5-2H     + PYGOUT_4 Tdome
;        5-2L     -
;
;        6-1H     + level x
;        6-1L     -
;
;        6-2H     + level y
;        6-2L     -
;
; OUTPUT MESSAGE FORMAT
;       LOGGER #
;       PYGIN_4
;       PYGOUT_4
;       PYGOUT_1
;       PYGIN_4 Tcase
;       PYGIN_4 Tdome
;       PYGOUT_4 Tcase
;       PYGOUT_4 Tdome
;       PYGOUT_1 Tcase
;       PYGOUT_1 Tdome
;       LEVEL X
;       LEVEL Y
;
;
; EEROM  SPACE USED FOR CALIBRATION COEFFICIENTS
; the coefficients below are examples!

;location      value        Serial Number   description
;   6          255.10       xxxxxF3         Epply pyrgeometer, in_1 W/m^2/mV
;   7          268.82       xxxxxF3         Epply pyrgeometer, out_1 W/m^2/mV
;   8          255.10       xxxxxF3         Epply pyrgeometer, in_2 W/m^2/mV
;   9          268.82       xxxxxF3         Epply pyrgeometer, out_2 W/m^2/mV
;  10          2500                         PYGIN_4m Tcase,  Vref mV
;  11          .00335595                    A0
;  12          .277486                      A1
;  13          4.8993                       A2
;  14          10.4799                      A3
;  15          2500                         PYGIN_4m Tdome, Vref mV
;  16          .00303674                    A0
;  17          .289526                      A1
;  18          -3.0453                      A2
;  19          841.413                      A3
;  20          2500                         PYGOUT_4m Tcase, Vref mV
;  21          .00335405                    A0
;  22          .28145                       A1
;  23          3.04961                      A2
;  24          289.824                      A3
;  25          2500                         PYGOUT_4m Tdome, Vref mV
;  26          .0030332                     A0
;  27          .291681                      A1
;  28          -3.4779                      A2
;  29          870.025                      A3
;  30          2500                         PYGIN_1.5m Tcase,  Vref mV
;  31          .00335595                    A0
;  32          .277486                      A1
;  33          4.8993                       A2
;  34          10.4799                      A3
;  35          2500                         PYGIN_1.5m Tdome, Vref mV
;  36          .00303674                    A0
;  37          .289526                      A1
;  38          -3.0453                      A2
;  39          841.413                      A3
;  50          0.0000        xxx            LEVEL_X, A0
;  51          11.1289                      A1
;  52          0.0000                       A2
;  53          11.1299                      A3
;  54          0.0000                       A4
;  55          11.1289                      LEVEL_Y, A0
;  56          0.0000                       A1
;  57          11.1299                      A2
;  58          0.0000                       A3
;  59          11.1289                      A4
;
;
*Table 1 Program
  01: 1.0       Execution Interval (seconds)

1:  Z=F (P30)
 1: 1        F
 2: 00       Exponent of 10
 3: 10       Z Loc [ LoggerId  ]

;  Start collecting data from the multiplexer

2:  Do (P86)
 1: 41       Set Port 1 High

;  loop thru 416mux to Epply Pile

3:  Do (P86)
 1: 9        Call Subroutine 9

4:  Beginning of Loop (P87)
 1: 0        Delay
 2: 4        Loop Count

     5:  Do (P86)
      1: 9        Call Subroutine 9

;  get pyg and pyr radiation data
     6:  Volt (Diff) (P2)
      1: 1        Reps
      2: 3         25 mV Slow Range
      3: 1        DIFF Channel
      4: 1     -- Loc [ pygin_raw ]
      5: 1.0      Mult
      6: 0.0      Offset

;  get Epply temperature data
     7:  Volt (Diff) (P2)
      1: 1        Reps
      2: 5         2500 mV Slow Range
      3: 2        DIFF Channel
      4: 14    -- Loc [ PYGI4_Tc  ]
      5: 1.0      Mult
      6: 0.0      Offset

8:  End (P95)

; get the level sensor data
     9:  Do (P86)
      1: 9        Call Subroutine 9

     10:  Volt (Diff) (P2)
      1: 1        Reps
      2: 5        2500 mV Slow Range
      3: 1        DIFF Channel
      4: 20       Loc [ LEVEL_X   ]
      5: .001     Mult
      6: 0.0      Offset

;  get Epply temperature data
     11:  Volt (Diff) (P2)
      1: 1        Reps
      2: 5         2500 mV Slow Range
      3: 2        DIFF Channel
      4: 21       Loc [ LEVEL_Y   ]
      5: .001     Mult
      6: 0.0      Offset


12:  Do (P86)
 1: 51       Set Port 1 Low

;  now get PYGOUT at 1 meter

13:  Volt (Diff) (P2)
 1: 1        Reps
 2: 3        25 mV Slow Range
 3: 4        DIFF Channel
 4: 13       Loc [ PYGOUT_1  ]
 5: 1.0      Mult
 6: 0.0      Offset


14:  Volt (Diff) (P2)
 1: 1        Reps
 2: 5        2500 mV Slow Range
 3: 5        DIFF Channel
 4: 18       Loc [ PYGO1_Tc  ]
 5: 1.0      Mult
 6: 0.0      Offset


15:  Volt (Diff) (P2)
 1: 1        Reps
 2: 5        2500 mV Slow Range
 3: 6        DIFF Channel
 4: 19       Loc [ PYGO1_Td  ]
 5: 1.0      Mult
 6: 0.0      Offset


;  Compute some real values, piles first

16:  Z=X*F (P37)
 1: 1        X Loc [ pygin_raw ]
 2: 0.0      F @@0
 3: 11       Z Loc [ PYGIN_4   ]

17:  Z=X*F (P37)
 1: 2        X Loc [ pygot_raw ]
 2: 0.0      F @@1
 3: 12       Z Loc [ PYGOUT_4  ]

18:  Z=X*F (P37)
 1: 13       X Loc [ PYGOUT_1  ]
 2: 0.0      F @@2
 3: 13       Z Loc [ PYGOUT_1  ]

; now lets do the temperatures

;  convert raw PYG.IN.Tcase at 4m

19:  Z=F (P30)
 1: 0        F @@10
 2: 00       Exponent of 10
 3: 31       Z Loc [ V_Vx      ]

20:  Z=X/Y (P38)
 1: 31       X Loc [ V_Vx      ]
 2: 14       Y Loc [ PYGI4_Tc  ]
 3: 31       Z Loc [ V_Vx      ]

21:  Do (P86)
 1: 2        Call Subroutine 2

22:  Polynomial (P55)
 1: 1        Reps
 2: 31       X Loc [ V_Vx      ]
 3: 31       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@12
 6: 0        C2 @@13
 7: 0        C3 @@14
 8: 0.0      C4
 9: 0.0      C5

23:  Z=F (P30)
 1: 0.0      F @@11
 2: -3       Exponent of 10
 3: 33       Z Loc [ B0        ]

24:  Z=X+Y (P33)
 1: 31       X Loc [ V_Vx      ]
 2: 33       Y Loc [ B0        ]
 3: 31       Z Loc [ V_Vx      ]

25:  Do (P86)
 1: 3        Call Subroutine 3

26:  Z=X (P31)
 1: 31       X Loc [ V_Vx      ]
 2: 14       Z Loc [ PYGI4_Tc  ]

;  convert raw PYG.IN.Tdome at 4m

27:  Z=F (P30)
 1: 0        F @@15
 2: 00       Exponent of 10
 3: 31       Z Loc [ V_Vx      ]

28:  Z=X/Y (P38)
 1: 31       X Loc [ V_Vx      ]
 2: 15       Y Loc [ PYGI4_Td  ]
 3: 31       Z Loc [ V_Vx      ]

29:  Do (P86)
 1: 2        Call Subroutine 2

30:  Polynomial (P55)
 1: 1        Reps
 2: 31       X Loc [ V_Vx      ]
 3: 31       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@17
 6: 0        C2 @@18
 7: 0        C3 @@19
 8: 0.0      C4
 9: 0.0      C5

31:  Z=F (P30)
 1: 0.0      F @@16
 2: -3       Exponent of 10
 3: 33       Z Loc [ B0        ]

32:  Z=X+Y (P33)
 1: 31       X Loc [ V_Vx      ]
 2: 33       Y Loc [ B0        ]
 3: 31       Z Loc [ V_Vx      ]

33:  Do (P86)
 1: 3        Call Subroutine 3

34:  Z=X (P31)
 1: 31       X Loc [ V_Vx      ]
 2: 15       Z Loc [ PYGI4_Td  ]

;  convert raw PYG.OUT.Tcase at 4m

35:  Z=F (P30)
 1: 0        F @@20
 2: 00       Exponent of 10
 3: 31       Z Loc [ V_Vx      ]

36:  Z=X/Y (P38)
 1: 31       X Loc [ V_Vx      ]
 2: 16       Y Loc [ PYGO4_Tc  ]
 3: 31       Z Loc [ V_Vx      ]

37:  Do (P86)
 1: 2        Call Subroutine 2

38:  Polynomial (P55)
 1: 1        Reps
 2: 31       X Loc [ V_Vx      ]
 3: 31       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@22
 6: 0        C2 @@23
 7: 0        C3 @@24
 8: 0.0      C4
 9: 0.0      C5

39:  Z=F (P30)
 1: 0.0      F @@21
 2: -3       Exponent of 10
 3: 33       Z Loc [ B0        ]

40:  Z=X+Y (P33)
 1: 31       X Loc [ V_Vx      ]
 2: 33       Y Loc [ B0        ]
 3: 31       Z Loc [ V_Vx      ]

41:  Do (P86)
 1: 3        Call Subroutine 3

42:  Z=X (P31)
 1: 31       X Loc [ V_Vx      ]
 2: 16       Z Loc [ PYGO4_Tc  ]

;  convert raw PYG.OUT.Tdome at 4m

43:  Z=F (P30)
 1: 0        F @@25
 2: 00       Exponent of 10
 3: 31       Z Loc [ V_Vx      ]

44:  Z=X/Y (P38)
 1: 31       X Loc [ V_Vx      ]
 2: 17       Y Loc [ PYGO4_Td  ]
 3: 31       Z Loc [ V_Vx      ]

45:  Do (P86)
 1: 2        Call Subroutine 2

46:  Polynomial (P55)
 1: 1        Reps
 2: 31       X Loc [ V_Vx      ]
 3: 31       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@27
 6: 0        C2 @@28
 7: 0        C3 @@29
 8: 0.0      C4
 9: 0.0      C5

47:  Z=F (P30)
 1: 0.0      F @@26
 2: -3       Exponent of 10
 3: 33       Z Loc [ B0        ]

48:  Z=X+Y (P33)
 1: 31       X Loc [ V_Vx      ]
 2: 33       Y Loc [ B0        ]
 3: 31       Z Loc [ V_Vx      ]

49:  Do (P86)
 1: 3        Call Subroutine 3

50:  Z=X (P31)
 1: 31       X Loc [ V_Vx      ]
 2: 17       Z Loc [ PYGO4_Td  ]

;  convert raw PYG.OUT.Tcase at 1.5m

51:  Z=F (P30)
 1: 0        F @@30
 2: 00       Exponent of 10
 3: 31       Z Loc [ V_Vx      ]

52:  Z=X/Y (P38)
 1: 31       X Loc [ V_Vx      ]
 2: 18       Y Loc [ PYGO1_Tc  ]
 3: 31       Z Loc [ V_Vx      ]

53:  Do (P86)
 1: 2        Call Subroutine 2

54:  Polynomial (P55)
 1: 1        Reps
 2: 31       X Loc [ V_Vx      ]
 3: 31       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@32
 6: 0        C2 @@33
 7: 0        C3 @@34
 8: 0.0      C4
 9: 0.0      C5

55:  Z=F (P30)
 1: 0.0      F @@31
 2: -3       Exponent of 10
 3: 33       Z Loc [ B0        ]

56:  Z=X+Y (P33)
 1: 31       X Loc [ V_Vx      ]
 2: 33       Y Loc [ B0        ]
 3: 31       Z Loc [ V_Vx      ]

57:  Do (P86)
 1: 3        Call Subroutine 3

58:  Z=X (P31)
 1: 31       X Loc [ V_Vx      ]
 2: 18       Z Loc [ PYGO1_Tc  ]

;  convert raw PYG.OUT.Tdome at 1.5m

59:  Z=F (P30)
 1: 0        F @@35
 2: 00       Exponent of 10
 3: 31       Z Loc [ V_Vx      ]

60:  Z=X/Y (P38)
 1: 31       X Loc [ V_Vx      ]
 2: 19       Y Loc [ PYGO1_Td  ]
 3: 31       Z Loc [ V_Vx      ]

61:  Do (P86)
 1: 2        Call Subroutine 2

62:  Polynomial (P55)
 1: 1        Reps
 2: 31       X Loc [ V_Vx      ]
 3: 31       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@37
 6: 0        C2 @@38
 7: 0        C3 @@39
 8: 0.0      C4
 9: 0.0      C5

63:  Z=F (P30)
 1: 0.0      F @@36
 2: -3       Exponent of 10
 3: 33       Z Loc [ B0        ]

64:  Z=X+Y (P33)
 1: 31       X Loc [ V_Vx      ]
 2: 33       Y Loc [ B0        ]
 3: 31       Z Loc [ V_Vx      ]

65:  Do (P86)
 1: 3        Call Subroutine 3

66:  Z=X (P31)
 1: 31       X Loc [ V_Vx      ]
 2: 19       Z Loc [ PYGO1_Td  ]

67:  Polynomial (P55)
 1: 1        Reps
 2: 20       X Loc [ LEVEL_X   ]
 3: 20       F(X) Loc [ LEVEL_X   ]
 4: 0.0      C0 @@50
 5: 0.0      C1 @@51
 6: 0.0      C2 @@52
 7: 0.0      C3 @@53
 8: 0.0      C4 @@54
 9: 0.0      C5 

68:  Polynomial (P55)
 1: 1        Reps
 2: 21       X Loc [ LEVEL_Y   ]
 3: 21       F(X) Loc [ LEVEL_Y   ]
 4: 0.0      C0 @@55
 5: 0.0      C1 @@56
 6: 0.0      C2 @@57
 7: 0.0      C3 @@58
 8: 0.0      C4 @@59
 9: 0.0      C5 

; compute level values


;  Is it time to send data out?

69:  Time (P18)
 1: 0        Seconds into current minute (maximum 60)
 2: 5        Mod/By
 3: 32       Loc [ TimeHack  ]

70:  IF (X<=>F) (P89)
 1: 32       X Loc [ TimeHack  ]
 2: 4        <
 3: 1        F
 4: 10       Set Output Flag High

71:  Sample (P70)
 1: 1        Reps
 2: 10       Loc [ LoggerId  ]

72:  Resolution (P78)
 1: 1        High Resolution

73:  Average (P71)
 1: 11       Reps
 2: 11       Loc [ PYGIN_4   ]

74:  Serial Out (P96)
 1: 52       Printer Comma/9600 Baud


*Table 2 Program
  02: 0.0000    Execution Interval (seconds)

*Table 3 Subroutines

1:  Beginning of Subroutine (P85)
 1: 2        Subroutine 2

;  compute x = ln[(V/Vx) - 1] / 1000

     2:  Z=X+F (P34)
      1: 31       X Loc [ V_Vx      ]
      2: -1       F
      3: 31       Z Loc [ V_Vx      ]

     3:  Z=LN(X) (P40)
      1: 31       X Loc [ V_Vx      ]
      2: 31       Z Loc [ V_Vx      ]

     4:  Z=X*F (P37)
      1: 31       X Loc [ V_Vx      ]
      2: .001     F
      3: 31       Z Loc [ V_Vx      ]

5:  End (P95)

6:  Beginning of Subroutine (P85)
 1: 3        Subroutine 3

;  finish T (case/dome) computation

     7:  Z=1/X (P42)
      1: 31       X Loc [ V_Vx      ]
      2: 31       Z Loc [ V_Vx      ]

     8:  Z=X+F (P34)
      1: 31       X Loc [ V_Vx      ]
      2: -273.15  F
      3: 31       Z Loc [ V_Vx      ]

9:  End (P95)

10:  Beginning of Subroutine (P85)
 1: 9        Subroutine 9

; pulse the AM416 multiplexer to the next bank of channels
     11:  Do (P86)
      1: 72       Pulse Port 2

     12:  Excitation with Delay (P22)
      1: 1        Ex Channel
      2: 0        Delay W/Ex (units = 0.01 sec)
      3: 1        Delay After Ex (units = 0.01 sec)
      4: 0        mV Excitation

13:  End (P95)

End Program

-Input Locations-
1 pygin_raw 1 1 1
2 pygot_raw 1 1 0
3 _________ 0 0 0
4 _________ 0 0 0
5 _________ 0 0 0
6 _________ 0 0 0
7 _________ 0 0 0
8 _________ 0 0 0
9 _________ 1 0 0
10 LoggerId  1 1 1
11 PYGIN_4   5 1 1
12 PYGOUT_4  9 1 1
13 PYGOUT_1  9 2 2
14 PYGI4_Tc  9 2 2
15 PYGI4_Td  9 2 1
16 PYGO4_Tc  9 2 1
17 PYGO4_Td  9 2 1
18 PYGO1_Tc  9 2 2
19 PYGO1_Td  17 2 2
20 LEVEL_X   1 1 2
21 LEVEL_Y   1 1 2
22 _________ 1 0 0
23 _________ 1 0 0
24 _________ 1 0 0
25 _________ 1 0 0
26 _________ 0 0 0
27 _________ 0 0 0
28 _________ 0 0 0
29 _________ 0 0 0
30 _________ 0 0 0
31 V_Vx      1 29 29
32 TimeHack  1 1 1
33 B0        1 6 6
34 _________ 0 0 0
35 _________ 0 0 0
36 _________ 0 0 0
37 _________ 0 0 0
38 _________ 0 0 0
39 _________ 0 0 0
40 _________ 0 0 0
41 _________ 0 0 0
42 _________ 1 0 0
-Program Security-
0000
0000
0000
-Mode 4-
0
1
2
10
11
12
13
14
15
16
17
18
19
20
21
22
23
24
25
26
27
28
29
30
31
32
33
34
35
36
37
38
39
50
51
52
53
54
55
56
57
58
59
-Final Storage Area 2-
0
-CR10X ID-
0
-CR10X Power Up-
3



17: Radiometers, Site tower, Fri 15-Nov-2002 14:22:45 MST, 25 meter radiation CR10X code
NOTE: comments in this code may not be up to date

;{CR10X}
; CASES99 - 30 meter radiation system
; This is a special program for the FLOS-II project
; The logger handles 4 pyrgeometers and 1 level sensor
;
; CR10 CHANNEL ASSIGNMENTS
; ANALOG CHANNELS
;        1H - comm 1H of AM416
;        1L - comm 1L of AM416
;        2H - comm 2H of AM416
;        2L - comm 2L of AM416
;
;        7  - level_X
;        8  - level_Y
;
; CONTROL PORTS
;        C1 - reset signal to AM416
;        C2 - clock to AM416
;        C3 - control to soil moisture CS-615
;
; AM416 MULTIPLEXER CHANNELS
;        1-1H     +  PYGIN_1
;        1-1L     -
;
;        1-2H     +  PYGIN_2
;        1-2L     -
;
;        2-1H     + PYGOUT_1
;        2-1L     -
;
;        2-2H     + PYGOUT_2
;        2-2L     -
;
;        3-1H     + PYGIN_1 Tcase
;        3-1L     -
;
;        3-2H     + PYGIN_1 Tdome
;        3-2L     -
;
;        4-1H     + PYGOUT_1 Tcase
;        4-1L     -
;
;        4-2H     + PYGOUT_1 Tdome
;        4-2L     -
;
;        5-1H     + PYGIN_2 Tcase
;        5-1L     -
;
;        5-2H     + PYGIN_2 Tdome
;        5-2L     -
;
;        6-1H     + PYGOUT_2 Tcase
;        6-1L     -
;
;        6-2H     + PYGOUT_2 Tdome
;        6-2L     -
;
; OUTPUT MESSAGE FORMAT
;       LOGGER #
;       PYGIN_1
;       PYGOUT_1
;       PYGIN_2
;       PYGOUT_2
;       PYGIN_1 Tcase
;       PYGOUT_1 Tcase
;       PYGIN_2 Tcase
;       PYGOUT_2 Tcase
;       PYGIN_1 Tdome
;       PYGOUT_1 Tdome
;       PYGIN_2 Tdome
;       PYGOUT_2 Tdome
;       level_X
;       level_Y
;
;
; EEROM  SPACE USED FOR CALIBRATION COEFFICIENTS
; the coefficients below are examples!

;location      value        Serial Number   description
;   6          255.10       xxxxxF3         Epply pyrgeometer, in_1 W/m^2/mV
;   7          268.82       xxxxxF3         Epply pyrgeometer, out_1 W/m^2/mV
;   8          255.10       xxxxxF3         Epply pyrgeometer, in_2 W/m^2/mV
;   9          268.82       xxxxxF3         Epply pyrgeometer, out_2 W/m^2/mV
;  10          2500                         PYGIN_1 Tcase,  Vref mV
;  11          .00335595                    A0
;  12          .277486                      A1
;  13          4.8993                       A2
;  14          10.4799                      A3
;  15          2500                         PYGIN_1 Tdome, Vref mV
;  16          .00303674                    A0
;  17          .289526                      A1
;  18          -3.0453                      A2
;  19          841.413                      A3
;  20          2500                         PYGOUT_1 Tcase, Vref mV
;  21          .00335405                    A0
;  22          .28145                       A1
;  23          3.04961                      A2
;  24          289.824                      A3
;  25          2500                         PYGOUT_1 Tdome, Vref mV
;  26          .0030332                     A0
;  27          .291681                      A1
;  28          -3.4779                      A2
;  29          870.025                      A3
;  30          2500                         PYGIN_2 Tcase,  Vref mV
;  31          .00335595                    A0
;  32          .277486                      A1
;  33          4.8993                       A2
;  34          10.4799                      A3
;  35          2500                         PYGIN_2 Tdome, Vref mV
;  36          .00303674                    A0
;  37          .289526                      A1
;  38          -3.0453                      A2
;  39          841.413                      A3
;  40          2500                         PYGOUT_2 Tcase, Vref mV
;  41          .00335405                    A0
;  42          .28145                       A1
;  43          3.04961                      A2
;  44          289.824                      A3
;  45          2500                         PYGOUT_2 Tdome, Vref mV
;  46          .0030332                     A0
;  47          .291681                      A1
;  48          -3.4779                      A2
;  49          870.025                      A3
;  50          0.0000        xxx            level_X, A0
;  51          11.1289                      A1
;  52          0.0000                       A2
;  53          11.1299                      A3
;  54          0.0000                       A4
;  55          11.1289                      level_Y, A0
;  56          0.0000                       A1
;  57          11.1299                      A2
;  58          0.0000                       A3
;  59          11.1289                      A4
;
;
*Table 1 Program
  01: 1.0       Execution Interval (seconds)

1:  Z=F (P30)
 1: 1        F
 2: 00       Exponent of 10
 3: 1        Z Loc [ LoggerId  ]

;  Start collecting data from the multiplexer

2:  Do (P86)
 1: 41       Set Port 1 High

;  loop thru 416mux to Epply Pile

3:  Beginning of Loop (P87)
 1: 0000     Delay
 2: 2        Count(s)

     4:  Do (P86)
      1: 9        Call Subroutine 9

;  get PYGIN data

     5:  Volt (Diff) (P2)
      1: 1        Reps
      2: 3         25 mV Slow Range
      3: 1        DIFF Channel
      4: 2     -- Loc [ PYGIN_1   ]
      5: 1.0      Mult
      6: 0.0      Offset


     6:  Volt (Diff) (P2)
      1: 1        Reps
      2: 3         25 mV Slow Range
      3: 2        DIFF Channel
      4: 4     -- Loc [ PYGIN_2   ]
      5: 1.0      Mult
      6: 0.0      Offset

7:  End (P95)

;  now get PYG temperature data

8:  Beginning of Loop (P87)
 1: 0        Delay
 2: 4        Count(s)

     9:  Do (P86)
      1: 9        Call Subroutine 9

     10:  Volt (Diff) (P2)
      1: 1        Reps
      2: 5         2500 mV Slow Range
      3: 1        DIFF Channel
      4: 6     -- Loc [ PIN_1_TC  ]
      5: 1.0      Mult
      6: 0.0      Offset


     11:  Volt (Diff) (P2)
      1: 1        Reps
      2: 5         2500 mV Slow Range
      3: 2        DIFF Channel
      4: 10    -- Loc [ PIN_1_TD  ]
      5: 1.0      Mult
      6: 0.0      Offset

12:  End (P95)

13:  Do (P86)
 1: 51       Set Port 1 Low


;  Now the level sensors

14:  Volts (SE) (P1)
 1: 2        Reps
 2: 5         2500 mV Slow Range
 3: 7        SE Channel
 4: 14    -- Loc [ LEVEL_X   ]
 5: .001     Mult
 6: 0.0      Offset

;  Compute some real values, piles first


15:  Scaling Array (A*Loc+B) (P53)
 1: 2        Start Loc [ PYGIN_1   ]
 2: 0.0      A1 @@6
 3: 0.0      B1
 4: 0.0      A2 @@7
 5: 0.0      B2
 6: 0.0      A3 @@8
 7: 0.0      B3
 8: 0.0      A4 @@9
 9: 0.0      B4

;  now the temperatures

;  convert raw PYGIN_1 Tcase to degrees C

16:  Z=F (P30)
 1: 0        F @@10
 2: 00       Exponent of 10
 3: 21       Z Loc [ V_Vx      ]

17:  Z=X/Y (P38)
 1: 21       X Loc [ V_Vx      ]
 2: 6        Y Loc [ PIN_1_TC  ]
 3: 21       Z Loc [ V_Vx      ]

18:  Do (P86)
 1: 2        Call Subroutine 2

19:  Polynomial (P55)
 1: 1        Reps
 2: 21       X Loc [ V_Vx      ]
 3: 21       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@12
 6: 0        C2 @@13
 7: 0        C3 @@14
 8: 0.0      C4
 9: 0.0      C5

20:  Z=F (P30)
 1: 0.0      F @@11
 2: -3       Exponent of 10
 3: 23       Z Loc [ B0        ]

21:  Z=X+Y (P33)
 1: 21       X Loc [ V_Vx      ]
 2: 23       Y Loc [ B0        ]
 3: 21       Z Loc [ V_Vx      ]

22:  Do (P86)
 1: 3        Call Subroutine 3

23:  Z=X (P31)
 1: 21       X Loc [ V_Vx      ]
 2: 6        Z Loc [ PIN_1_TC  ]

;  convert raw PYGOUT_1 Tcase to degrees C

24:  Z=F (P30)
 1: 0        F @@15
 2: 00       Exponent of 10
 3: 21       Z Loc [ V_Vx      ]

25:  Z=X/Y (P38)
 1: 21       X Loc [ V_Vx      ]
 2: 7        Y Loc [ POT_1_TC  ]
 3: 21       Z Loc [ V_Vx      ]

26:  Do (P86)
 1: 2        Call Subroutine 2

27:  Polynomial (P55)
 1: 1        Reps
 2: 21       X Loc [ V_Vx      ]
 3: 21       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@17
 6: 0        C2 @@18
 7: 0        C3 @@19
 8: 0.0      C4
 9: 0.0      C5

28:  Z=F (P30)
 1: 0.0      F @@16
 2: -3       Exponent of 10
 3: 23       Z Loc [ B0        ]

29:  Z=X+Y (P33)
 1: 21       X Loc [ V_Vx      ]
 2: 23       Y Loc [ B0        ]
 3: 21       Z Loc [ V_Vx      ]

30:  Do (P86)
 1: 3        Call Subroutine 3

31:  Z=X (P31)
 1: 21       X Loc [ V_Vx      ]
 2: 7        Z Loc [ POT_1_TC  ]


;  convert raw PYGIN_2 Tcase to degrees C

32:  Z=F (P30)
 1: 0        F @@20
 2: 00       Exponent of 10
 3: 21       Z Loc [ V_Vx      ]

33:  Z=X/Y (P38)
 1: 21       X Loc [ V_Vx      ]
 2: 8        Y Loc [ PIN_2_TC  ]
 3: 21       Z Loc [ V_Vx      ]

34:  Do (P86)
 1: 2        Call Subroutine 2

35:  Polynomial (P55)
 1: 1        Reps
 2: 21       X Loc [ V_Vx      ]
 3: 21       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@22
 6: 0        C2 @@23
 7: 0        C3 @@24
 8: 0.0      C4
 9: 0.0      C5

36:  Z=F (P30)
 1: 0.0      F @@21
 2: -3       Exponent of 10
 3: 23       Z Loc [ B0        ]

37:  Z=X+Y (P33)
 1: 21       X Loc [ V_Vx      ]
 2: 23       Y Loc [ B0        ]
 3: 21       Z Loc [ V_Vx      ]

38:  Do (P86)
 1: 3        Call Subroutine 3

39:  Z=X (P31)
 1: 21       X Loc [ V_Vx      ]
 2: 8        Z Loc [ PIN_2_TC  ]

;  convert raw PYGOUT_2 Tcase to degrees C

40:  Z=F (P30)
 1: 0        F @@25
 2: 00       Exponent of 10
 3: 21       Z Loc [ V_Vx      ]

41:  Z=X/Y (P38)
 1: 21       X Loc [ V_Vx      ]
 2: 9        Y Loc [ POT_2_TC  ]
 3: 21       Z Loc [ V_Vx      ]

42:  Do (P86)
 1: 2        Call Subroutine 2

43:  Polynomial (P55)
 1: 1        Reps
 2: 21       X Loc [ V_Vx      ]
 3: 21       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@27
 6: 0        C2 @@28
 7: 0        C3 @@29
 8: 0.0      C4
 9: 0.0      C5

44:  Z=F (P30)
 1: 0.0      F @@26
 2: -3       Exponent of 10
 3: 23       Z Loc [ B0        ]

45:  Z=X+Y (P33)
 1: 21       X Loc [ V_Vx      ]
 2: 23       Y Loc [ B0        ]
 3: 21       Z Loc [ V_Vx      ]

46:  Do (P86)
 1: 3        Call Subroutine 3

47:  Z=X (P31)
 1: 21       X Loc [ V_Vx      ]
 2: 9        Z Loc [ POT_2_TC  ]

;  convert raw PYGIN_1 Tdome to degrees C

48:  Z=F (P30)
 1: 0        F @@30
 2: 00       Exponent of 10
 3: 21       Z Loc [ V_Vx      ]

49:  Z=X/Y (P38)
 1: 21       X Loc [ V_Vx      ]
 2: 10       Y Loc [ PIN_1_TD  ]
 3: 21       Z Loc [ V_Vx      ]

50:  Do (P86)
 1: 2        Call Subroutine 2

51:  Polynomial (P55)
 1: 1        Reps
 2: 21       X Loc [ V_Vx      ]
 3: 21       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@32
 6: 0        C2 @@33
 7: 0        C3 @@34
 8: 0.0      C4
 9: 0.0      C5

52:  Z=F (P30)
 1: 0.0      F @@31
 2: -3       Exponent of 10
 3: 23       Z Loc [ B0        ]

53:  Z=X+Y (P33)
 1: 21       X Loc [ V_Vx      ]
 2: 23       Y Loc [ B0        ]
 3: 21       Z Loc [ V_Vx      ]

54:  Do (P86)
 1: 3        Call Subroutine 3

55:  Z=X (P31)
 1: 21       X Loc [ V_Vx      ]
 2: 10       Z Loc [ PIN_1_TD  ]


;  convert raw PYGOUT_1 Tdome to degrees C

56:  Z=F (P30)
 1: 0        F @@35
 2: 00       Exponent of 10
 3: 21       Z Loc [ V_Vx      ]

57:  Z=X/Y (P38)
 1: 21       X Loc [ V_Vx      ]
 2: 11       Y Loc [ POT_1_TD  ]
 3: 21       Z Loc [ V_Vx      ]

58:  Do (P86)
 1: 2        Call Subroutine 2

59:  Polynomial (P55)
 1: 1        Reps
 2: 21       X Loc [ V_Vx      ]
 3: 21       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@37
 6: 0        C2 @@38
 7: 0        C3 @@39
 8: 0.0      C4
 9: 0.0      C5

60:  Z=F (P30)
 1: 0.0      F @@36
 2: -3       Exponent of 10
 3: 23       Z Loc [ B0        ]

61:  Z=X+Y (P33)
 1: 21       X Loc [ V_Vx      ]
 2: 23       Y Loc [ B0        ]
 3: 21       Z Loc [ V_Vx      ]

62:  Do (P86)
 1: 3        Call Subroutine 3

63:  Z=X (P31)
 1: 21       X Loc [ V_Vx      ]
 2: 11       Z Loc [ POT_1_TD  ]


;  convert raw PYGIN_2 Tdome to degrees C

64:  Z=F (P30)
 1: 0        F @@40
 2: 00       Exponent of 10
 3: 21       Z Loc [ V_Vx      ]

65:  Z=X/Y (P38)
 1: 21       X Loc [ V_Vx      ]
 2: 12       Y Loc [ PIN_2_TD  ]
 3: 21       Z Loc [ V_Vx      ]

66:  Do (P86)
 1: 2        Call Subroutine 2

67:  Polynomial (P55)
 1: 1        Reps
 2: 21       X Loc [ V_Vx      ]
 3: 21       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@42
 6: 0        C2 @@43
 7: 0        C3 @@44
 8: 0.0      C4
 9: 0.0      C5

68:  Z=F (P30)
 1: 0.0      F @@41
 2: -3       Exponent of 10
 3: 23       Z Loc [ B0        ]

69:  Z=X+Y (P33)
 1: 21       X Loc [ V_Vx      ]
 2: 23       Y Loc [ B0        ]
 3: 21       Z Loc [ V_Vx      ]

70:  Do (P86)
 1: 3        Call Subroutine 3

71:  Z=X (P31)
 1: 21       X Loc [ V_Vx      ]
 2: 12       Z Loc [ PIN_2_TD  ]

;  convert raw PYGOUT_2 Tdome to degrees C

72:  Z=F (P30)
 1: 0        F @@45
 2: 00       Exponent of 10
 3: 21       Z Loc [ V_Vx      ]

73:  Z=X/Y (P38)
 1: 21       X Loc [ V_Vx      ]
 2: 13       Y Loc [ POT_2_TD  ]
 3: 21       Z Loc [ V_Vx      ]

74:  Do (P86)
 1: 2        Call Subroutine 2

75:  Polynomial (P55)
 1: 1        Reps
 2: 21       X Loc [ V_Vx      ]
 3: 21       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@47
 6: 0        C2 @@48
 7: 0        C3 @@49
 8: 0.0      C4
 9: 0.0      C5

76:  Z=F (P30)
 1: 0.0      F @@46
 2: -3       Exponent of 10
 3: 23       Z Loc [ B0        ]

77:  Z=X+Y (P33)
 1: 21       X Loc [ V_Vx      ]
 2: 23       Y Loc [ B0        ]
 3: 21       Z Loc [ V_Vx      ]

78:  Do (P86)
 1: 3        Call Subroutine 3

79:  Z=X (P31)
 1: 21       X Loc [ V_Vx      ]
 2: 13       Z Loc [ POT_2_TD  ]


;  Now the level sensors

80:  Polynomial (P55)
 1: 1        Reps
 2: 14       X Loc [ LEVEL_X   ]
 3: 14       F(X) Loc [ LEVEL_X   ]
 4: 0.0      C0 @@50
 5: 0.0      C1 @@51
 6: 0.0      C2 @@52
 7: 0.0      C3 @@53
 8: 0.0      C4 @@54
 9: 0.0      C5

81:  Polynomial (P55)
 1: 1        Reps
 2: 15       X Loc [ LEVEL_Y   ]
 3: 15       F(X) Loc [ LEVEL_Y   ]
 4: 0.0      C0 @@55
 5: 0.0      C1 @@56
 6: 0.0      C2 @@57
 7: 0.0      C3 @@58
 8: 0.0      C4 @@59
 9: 0.0      C5

;  Is it time to send data out?

82:  Time (P18)
 1: 0        Seconds into current minute (maximum 60)
 2: 5        Mod/By
 3: 22       Loc [ TimeHack  ]

83:  IF (X<=>F) (P89)
 1: 22       X Loc [ TimeHack  ]
 2: 4        <
 3: 1        F
 4: 10       Set Output Flag High

84:  Sample (P70)
 1: 1        Reps
 2: 1        Loc [ LoggerId  ]

85:  Resolution (P78)
 1: 1        High Resolution

86:  Average (P71)
 1: 14       Reps
 2: 2        Loc [ PYGIN_1   ]

87:  Serial Out (P96)
 1: 52       Printer Comma/9600 Baud


*Table 2 Program
  02: 0.0000    Execution Interval (seconds)

*Table 3 Subroutines
1:  Beginning of Subroutine (P85)
 1: 2        Subroutine 2

;  compute x = ln[(V/Vx) - 1] / 1000

     2:  Z=X+F (P34)
      1: 21       X Loc [ V_Vx      ]
      2: -1       F
      3: 21       Z Loc [ V_Vx      ]

     3:  Z=LN(X) (P40)
      1: 21       X Loc [ V_Vx      ]
      2: 21       Z Loc [ V_Vx      ]

     4:  Z=X*F (P37)
      1: 21       X Loc [ V_Vx      ]
      2: .001     F
      3: 21       Z Loc [ V_Vx      ]

5:  End (P95)

6:  Beginning of Subroutine (P85)
 1: 3        Subroutine 3

;  finish T (case/dome) computation

     7:  Z=1/X (P42)
      1: 21       X Loc [ V_Vx      ]
      2: 21       Z Loc [ V_Vx      ]

     8:  Z=X+F (P34)
      1: 21       X Loc [ V_Vx      ]
      2: -273.15  F
      3: 21       Z Loc [ V_Vx      ]

9:  End (P95)

10:  Beginning of Subroutine (P85)
 1: 9        Subroutine 9

; pulse the AM416 multiplexer to the next bank of channels
     11:  Do (P86)
      1: 72       Pulse Port 2

     12:  Excitation with Delay (P22)
      1: 1        Ex Channel
      2: 0        Delay W/Ex (units = 0.01 sec)
      3: 1        Delay After Ex (units = 0.01 sec)
      4: 0        mV Excitation

13:  End (P95)



End Program

-Input Locations-
1 LoggerId  1 1 1
2 PYGIN_1   5 2 2
3 PYGOUT_1  9 2 1
4 PYGIN_2   9 2 2
5 PYGOUT_2  25 2 1
6 PIN_1_TC  9 2 2
7 POT_1_TC  9 2 1
8 PIN_2_TC  9 2 1
9 POT_2_TC  9 2 1
10 PIN_1_TD  9 2 2
11 POT_1_TD  9 2 1
12 PIN_2_TD  9 2 1
13 POT_2_TD  9 2 1
14 LEVEL_X   13 2 2
15 LEVEL_Y   17 2 2
16 _________ 0 0 0
17 _________ 0 0 0
18 _________ 0 0 0
19 _________ 0 0 0
20 _________ 1 0 0
21 V_Vx      1 37 37
22 TimeHack  1 1 1
23 B0        1 8 8
24 _________ 0 0 0
25 _________ 0 0 0
26 _________ 0 0 0
27 _________ 0 0 0
28 _________ 0 0 0
-Program Security-
0000
0000
0000
-Mode 4-
6
7
8
9
10
11
12
13
14
15
16
17
18
19
20
21
22
23
24
25
26
27
28
29
30
31
32
33
34
35
36
37
38
39
40
41
42
43
44
45
46
47
48
49
50
51
52
53
54
55
56
57
58
59
-Final Storage Area 2-
0
-CR10X ID-
0
-CR10X Power Up-
3



16: Radiometers, Site all stns, Fri 15-Nov-2002 13:49:33 MST, Grass, Lake, and Sage CR10X code
NOTE: comments in this code may not be up to date

;{CR10}
; PROGRAM:  FLOSS02 PLATFORM
; This program is specific to the FLOSS02 platforms. Additional
; sensors have been added to the data collection. A soil profile
; consisting of 3 sets of 2 levels with 2 sensors at each level has been added.
; These new data are multiplexed thru a AM16/32 unit.
; Refer to the information below on wiring.
; NOTE: An additional tsoil has been added to each soil profile

; CR10 CHANNEL ASSIGNMENTS
; ANALOG CHANNELS
;        1H - comm 1H of AM416
;        1L - comm 1L of AM416
;        2H - comm 2H of AM416
;        2L - comm 2L of AM416
;
;        3H - comm 1L of AM16/32 profile 1
;        3L - comm 2L of AM16/32 profile 1
;
;        4H + TP01, Hukseflux sensor
;        4L -
;
;        5H + TP01 heater voltage
;        5L -
; 
; PULSE CHANNELS
;        P1 - tipping bucket rain input
;
; EXCITATION CHANNELS
;        E1 - excitation for T107 probes (H1 comm)
;        E2 - excitation for ECHO probes (H2 comm)
;        E3 - excitation for T soil probe

; NOTE: E1 and E2 get wired to H1 comm and H2 comm of the AM16/32 units

; CONTROL PORTS
;        C1 - reset signal to AM416
;        C2 - clock to AM416
;        C3 - control to soil moisture CS-615
;        C4 - reset signal to all AM16/32
;
; AM416 MULTIPLEXER CHANNELS
;        1-1H     +  Q7 net radiation
;        1-1L     -
;
;        1-2H     +  LICOR 200-SA
;        1-2L     -
;
;        2-1H     + PYG_IN thermopile
;        2-1L     -
;
;        3-1H     + PYG_OUT thermopile
;        3-1L     -
;
;        4-1H     + PYR_IN (KIP)
;        4-1L     -
;
;        5-1H     + PYR_OUT (KIP)
;        5-1L     -
;
;        2-2H     + PYG_IN T_case
;        2-2L     -
;
;        3-2H     + PYG_IN T_dome
;        3-2L     -
;
;        4-2H     + PYG_OUT T_case
;        4-2L     -
;
;        5-2H     + PYG_OUT T_dome
;        5-2L     -
;
;        6-1H     + level sensor, X
;        6-1L     -
;
;        6-2H     + level sensor, Y
;        6-2L     -
;
;        7-1H     + HFT
;        7-1L     -
;
;        8-1H     Tsoil, refer to CR10X manual
;        8-1L
;        8-2H
;        8-2L

; AM16/32 MULTIPLEXER CHANNELS

;        1-1H     T107-1 excitation
;        1-1L     T107-1 signal
;        1-2H     ECHO-1 excitation
;        1-2L     ECHO-1 signal
;
;        2-1H     T107-2 excitation
;        2-1L     T107-2 signal
;        2-2H     ECHO-2 excitation
;        2-2L     ECHO-2 signal
;
;        3-1H     T107-3 excitation
;        3-1L     T107-3 signal
;        3-2H     ECHO-3 excitation
;        3-2L     ECHO-3 signal
;
;        4-1H     T107-4 excitation
;        4-1L     T107-4 signal
;        4-2H     ECHO-4 excitation
;        4-2L     ECHO-4 signal
;
;        5-1H     T107-5 excitation
;        5-1L     T107-5 signal
;        5-2H     ECHO-5 excitation
;        5-2L     ECHO-5 signal
;
;        6-1H     T107-6 excitation
;        6-1L     T107-6 signal
;        6-2H     ECHO-6 excitation
;        6-2L     ECHO-6 signal
;
;        7-1H     T107-Tref excitation
;        7-1L     T107-Tref signal
;
;        8-1l     + HFT 1
;        8-2L     -
;
;        9-1L     + HFT 2
;        9-2L     -
;
;        10-1L    + HFT 3
;        10-2L    -
;
;        11-1H    T107-7 excitation
;        11-1L    T107-7 signal
;
;        12-1H    T107-8 excitation
;        12-1L    T107-8 signal
;
;        13-1H    T107-9 excitation
;        13-1L    T107-9 signal

; FLOSS-II OUTPUT MESSAGE FORMAT
;       LOGGER ID
;       rain (tips per 10 second period)
;       Q7
;       PYG_IN
;       PYG_OUT
;       PYR_IN
;       PYR_OUT
;       PYG_IN Tcase
;       PYG_IN Tdome
;       PYG_OUT Tcase
;       PYG_OUT Tdome
;       T107 - 1
;       T107 - 2
;       T107 - 3
;       T107 - 4
;       T107 - 5
;       T107 - 6
;       T107 - 7
;       T107 - 8
;       T107 - 9
;       ECHO - 1
;       ECHO - 2
;       ECHO - 3
;       ECHO - 4
;       EHCO - 5
;       ECHO - 6
;       HFT1
;       HFT2
;       HFT3
;       TP01 senosr
;       TP01 heater
;       TP01 time 63%
;       

; EEROM  SPACE USED FOR CALIBRATION COEFFICIENTS

;location      value        Serial Number   description
;   0          0.00         xxxxxx          Q7 positive output  W/m^2/mV
;   1          0.00                         negative output
;   2                       xxxxxx          PYG in
;   3                       xxxxxx          PYG out
;   4                       xxxxxx          PSP in
;   5                       xxxxxx          PSP out
;   6          35.6         Hxxxxx          REBS Heat Flux Transducer, W/m^2/mV
;   7          35.6         Hxxxxx          REBS Heat Flux Transducer, W/m^2/mV
;   8          35.6         Hxxxxx          REBS Heat Flux Transducer, W/m^2/mV
;   9          xxxx         xxxxxx          LICOR quantum sensor
;  10                                       A0 pygin_Tc
;  11                                       A1
;  12                                       A2
;  13                                       A3
;  14                                       A4
;  15                                       A0 pygin_Td
;  16                                       A1
;  17                                       A2
;  18                                       A3
;  19                                       A4
;  20                                       A0 pygout_Tc
;  21                                       A1
;  22                                       A2
;  23                                       A3
;  24                                       A4
;  25                                       A0 pygout_Td
;  26                                       A1
;  27                                       A2
;  28                                       A3
;  29                                       A4

; SUBROUTINES
;
; sub 1 : Q7 calibration
;
; sub 2 : stienhart equation
;
; sub 3 : stienhart equation
;
; sub 8 : step mulitplexer for soil profile array

; NOTE: the calibation for the ECHO probe is embedded in the
;       data collection code, instruction # 12

*Table 1 Program
  01: 10        Execution Interval (seconds)


1:  Z=F (P30)
 1: 1        F
 2: 00       Exponent of 10
 3: 10       Z Loc [ LoggerID  ]


2:  Pulse (P3)
 1: 1        Reps
 2: 1        Pulse Input Channel
 3: 2        Switch Closure, All Counts
 4: 11       Loc [ rain      ]
 5: 1.0      Mult
 6: 0.0      Offset

;  loop thru 416mux to get radiation data
3:  Do (P86)
 1: 41       Set Port 1 High

4:  Do (P86)
 1: 9        Call Subroutine 9

5:  Volt (Diff) (P2)
 1: 1        Reps
 2: 4         250 mV Slow Range
 3: 1        DIFF Channel
 4: 12       Loc [ Q7        ]
 5: 1.0      Mult
 6: 0.0      Offset

6:  Beginning of Loop (P87)
 1: 0        Delay
 2: 4        Loop Count

     7:  Do (P86)
      1: 9        Call Subroutine 9

;  get pyg and pyr radiation data
     8:  Volt (Diff) (P2)
      1: 1        Reps
      2: 3         25 mV Slow Range
      3: 1        DIFF Channel
      4: 5     -- Loc [ pygi_raw  ]
      5: 1.0      Mult
      6: 0.0      Offset

;  get Epply temperature data
     9:  Volt (Diff) (P2)
      1: 1        Reps
      2: 5         2500 mV Slow Range
      3: 2        DIFF Channel
      4: 17    -- Loc [ pygin_Tc  ]
      5: 1.0      Mult
      6: 0.0      Offset

10:  End (P95)

; end mutliplexer data collection
11:  Do (P86)
 1: 51       Set Port 1 Low


;  convert raw data to calibrated information

;  Q7 convert
12:  Do (P86)
 1: 1        Call Subroutine 1

;  convert PYG and PYR data
13:  Block Move (P54)
 1: 4        No. of Values
 2: 5        First Source Loc [ pygi_raw  ]
 3: 1        Source Step
 4: 13       First Destination Loc [ pygin     ]
 5: 1        Destination Step

14:  Scaling Array (A*Loc+B) (P53)
 1: 13       Start Loc [ pygin     ]
 2: 0.0      A1 @@2
 3: 0.0      B1
 4: 0.0      A2 @@3
 5: 0.0      B2
 6: 0.0      A3 @@4
 7: 0.0      B3
 8: 0.0      A4 @@5
 9: 0.0      B4

; now convert pyg temps

;  convert raw PYGIN Tcase

15:  Z=F (P30)
 1: 0        F @@10
 2: 00       Exponent of 10
 3: 50       Z Loc [ V_Vx      ]

16:  Z=X/Y (P38)
 1: 50       X Loc [ V_Vx      ]
 2: 17       Y Loc [ pygin_Tc  ]
 3: 50       Z Loc [ V_Vx      ]

17:  Do (P86)
 1: 2        Call Subroutine 2

18:  Polynomial (P55)
 1: 1        Reps
 2: 50       X Loc [ V_Vx      ]
 3: 50       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@12
 6: 0        C2 @@13
 7: 0        C3 @@14
 8: 0.0      C4
 9: 0.0      C5

19:  Z=F (P30)
 1: 0.0      F @@11
 2: -3       Exponent of 10
 3: 51       Z Loc [ B0        ]

20:  Z=X+Y (P33)
 1: 50       X Loc [ V_Vx      ]
 2: 51       Y Loc [ B0        ]
 3: 50       Z Loc [ V_Vx      ]

21:  Do (P86)
 1: 3        Call Subroutine 3

22:  Z=X (P31)
 1: 50       X Loc [ V_Vx      ]
 2: 17       Z Loc [ pygin_Tc  ]

;  convert raw PYGIN Tdome

23:  Z=F (P30)
 1: 0        F @@15
 2: 00       Exponent of 10
 3: 50       Z Loc [ V_Vx      ]

24:  Z=X/Y (P38)
 1: 50       X Loc [ V_Vx      ]
 2: 18       Y Loc [ pygin_Td  ]
 3: 50       Z Loc [ V_Vx      ]

25:  Do (P86)
 1: 2        Call Subroutine 2

26:  Polynomial (P55)
 1: 1        Reps
 2: 50       X Loc [ V_Vx      ]
 3: 50       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@17
 6: 0        C2 @@18
 7: 0        C3 @@19
 8: 0.0      C4
 9: 0.0      C5

27:  Z=F (P30)
 1: 0.0      F @@16
 2: -3       Exponent of 10
 3: 51       Z Loc [ B0        ]

28:  Z=X+Y (P33)
 1: 50       X Loc [ V_Vx      ]
 2: 51       Y Loc [ B0        ]
 3: 50       Z Loc [ V_Vx      ]

29:  Do (P86)
 1: 3        Call Subroutine 3

30:  Z=X (P31)
 1: 50       X Loc [ V_Vx      ]
 2: 18       Z Loc [ pygin_Td  ]

;  convert raw PYGOUT Tcase

31:  Z=F (P30)
 1: 0        F @@20
 2: 00       Exponent of 10
 3: 50       Z Loc [ V_Vx      ]

32:  Z=X/Y (P38)
 1: 50       X Loc [ V_Vx      ]
 2: 19       Y Loc [ pygout_Tc ]
 3: 50       Z Loc [ V_Vx      ]

33:  Do (P86)
 1: 2        Call Subroutine 2

34:  Polynomial (P55)
 1: 1        Reps
 2: 50       X Loc [ V_Vx      ]
 3: 50       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@22
 6: 0        C2 @@23
 7: 0        C3 @@24
 8: 0.0      C4
 9: 0.0      C5

35:  Z=F (P30)
 1: 0.0      F @@21
 2: -3       Exponent of 10
 3: 51       Z Loc [ B0        ]

36:  Z=X+Y (P33)
 1: 50       X Loc [ V_Vx      ]
 2: 51       Y Loc [ B0        ]
 3: 50       Z Loc [ V_Vx      ]

37:  Do (P86)
 1: 3        Call Subroutine 3

38:  Z=X (P31)
 1: 50       X Loc [ V_Vx      ]
 2: 19       Z Loc [ pygout_Tc ]

;  convert raw PYGOUT Tdome

39:  Z=F (P30)
 1: 0        F @@25
 2: 00       Exponent of 10
 3: 50       Z Loc [ V_Vx      ]

40:  Z=X/Y (P38)
 1: 50       X Loc [ V_Vx      ]
 2: 20       Y Loc [ pygout_Td ]
 3: 50       Z Loc [ V_Vx      ]

41:  Do (P86)
 1: 2        Call Subroutine 2

42:  Polynomial (P55)
 1: 1        Reps
 2: 50       X Loc [ V_Vx      ]
 3: 50       F(X) Loc [ V_Vx      ]
 4: 0.0      C0
 5: 0        C1 @@27
 6: 0        C2 @@28
 7: 0        C3 @@29
 8: 0.0      C4
 9: 0.0      C5

43:  Z=F (P30)
 1: 0.0      F @@26
 2: -3       Exponent of 10
 3: 51       Z Loc [ B0        ]

44:  Z=X+Y (P33)
 1: 50       X Loc [ V_Vx      ]
 2: 51       Y Loc [ B0        ]
 3: 50       Z Loc [ V_Vx      ]

45:  Do (P86)
 1: 3        Call Subroutine 3

46:  Z=X (P31)
 1: 50       X Loc [ V_Vx      ]
 2: 20       Z Loc [ pygout_Td ]

; here is the new code for the profile stuff

47:  Do (P86)
 1: 44       Set Port 4 High

;  Loop to collect T107 and ECH2O
48:  Beginning of Loop (P87)
 1: 0000     Delay
 2: 6        Loop Count

     49:  Do (P86)
      1: 8        Call Subroutine 8

     50:  Temp (107) (P11)
      1: 1        Reps
      2: 5        SE Channel
      3: 1        Excite all reps w/E1
      4: 21    -- Loc [ T107_1    ]
      5: 1.0      Mult
      6: 0.0      Offset

; Echo calibration is applied at this instruction
     51:  Excite-Delay (SE) (P4)
      1: 1        Reps
      2: 5        2500 mV Slow Range
      3: 6        SE Channel
      4: 2        Excite all reps w/Exchan 2
      5: 1        Delay (units 0.01 sec)
      6: 2500     mV Excitation
      7: 30    -- Loc [ ECHO_1    ]
      8: .0695    Mult
      9: -29.0    Offset

52:  End (P95)

; pulse AM16/32 forward to HFT stuff, skip Tref
53:  Do (P86)
 1: 8        Call Subroutine 8

;  Loop to collect T107 and ECH2O
54:  Beginning of Loop (P87)
 1: 0000     Delay
 2: 3        Loop Count

     55:  Do (P86)
      1: 8        Call Subroutine 8

     56:  Volt (Diff) (P2)
      1: 1        Reps
      2: 3        25 mV Slow Range
      3: 3        DIFF Channel
      4: 36    -- Loc [ HFT1      ]
      5: 1.0      Mult
      6: 0.0      Offset

57:  End (P95)

;  Loop to collect 3 more T107 probes
58:  Beginning of Loop (P87)
 1: 0000     Delay
 2: 3        Loop Count

     59:  Do (P86)
      1: 8        Call Subroutine 8

     60:  Temp (107) (P11)
      1: 1        Reps
      2: 5        SE Channel
      3: 1        Excite all reps w/E1
      4: 27    -- Loc [ T107_7    ]
      5: 1.0      Mult
      6: 0.0      Offset

61:  End (P95)

; turn off AM16/32
62:  Do (P86)
 1: 54       Set Port 4 Low

; convert HFT data to engineering units

;  HFT1
63:  Z=X*F (P37)
 1: 36       X Loc [ HFT1      ]
 2: 0.0      F @@6
 3: 36       Z Loc [ HFT1      ]

;  HFT2
64:  Z=X*F (P37)
 1: 37       X Loc [ HFT2      ]
 2: 0.0      F @@7
 3: 37       Z Loc [ HFT2      ]

;  HFT3
65:  Z=X*F (P37)
 1: 38       X Loc [ HFT3      ]
 2: 0.0      F @@8
 3: 38       Z Loc [ HFT3      ]


;xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx

; now add the code for the Hukseflux TP01 sensor

; Measures the TP01 sensor output, parameter U in manual.
66:  Volt (Diff) (P2)
 1: 1        Reps
 2: 3        25 mV Slow Range
 3: 4        DIFF Channel
 4: 39       Loc [ U_sen     ]
 5: 1.0      Mult
 6: 0.0      Offset

; Measures the heater voltage, called U_heat in the manual.
; Multiplication by 0.001 to convert to Volts
67:  Volt (Diff) (P2)
 1: 1        Reps
 2: 5        2500 mV Slow Range
 3: 5        DIFF Channel 
 4: 40       Loc [ U_heat    ]
 5: 0.001    Mult
 6: 0.0      Offset

; xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx

; ENTER the calibration interval
; Recommended: every 180 minutes (= 3 hours) start a measurement
; of thermal conductivity and diffusivity
68:  If time is (P92)
 1: 0        Minutes (Seconds --) into a
 2: 180      Interval (same units as above)
 3: 30       Then Do

; Copy the reading of the TP01 differential temperature sensor
     69:  Z=X (P31)
      1: 39       X Loc [ U_sen     ]
      2: 55       Z Loc [ U_sen0    ]

; Put the heater on, control via port 7 and CR10X 12V switch
     70:  Do (P86)
      1: 47       Set Port 7 High

71:  End (P95)

; xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx

; 3 minutes (180 seconds) after the heating started,
; the sensor output will be stable, and another measurement is taken.
72:  If time is (P92)
 1: 3        Minutes (Seconds --) into a
 2: 180      Interval (same units as above)
 3: 30       Then Do

; Take the sensor output.
     73:  Z=X (P31)
      1: 39       X Loc [ U_sen     ]
      2: 56       Z Loc [ U_senmax  ]

     74:  Z=X (P31)
      1: 40       X Loc [ U_heat    ]
      2: 57       Z Loc [ U_heatmax ]

; Take the voltage over the heater.
     75:  Z=X*Y (P36)
      1: 57       X Loc [ U_heatmax ]
      2: 57       Y Loc [ U_heatmax ]
      3: 58       Z Loc [ U_heat_sq ]

; Calculate the mV difference caused due to the heater
     76:  Z=X-Y (P35)
      1: 56       X Loc [ U_senmax  ]
      2: 55       Y Loc [ U_sen0    ]
      3: 59       Z Loc [ U_senamp  ]

; Determination of 63% of the signal amplitude.
     77:  Z=X*F (P37)
      1: 59       X Loc [ U_senamp  ] ;
      2: 0.63     F
      3: 60       Z Loc [ U_sen90   ]

; The value to compare with is U_comp.
     78:  Z=X-Y (P35)
      1: 56       X Loc [ U_senmax  ]
      2: 60       Y Loc [ U_sen90   ]
      3: 61       Z Loc [ U_comp    ]

; The absolute value to avoid problems with polarity
     79:  Z=ABS(X) (P43)
      1: 61       X Loc [ U_comp    ]
      2: 61       Z Loc [ U_comp    ]

; Turn off the heater, the signal now starts settling down.
     80:  Do (P86)
      1: 57       Set Port 7 Low

; Set the timer for the thermal diffusivity measurement to zero
     81:  Z=F (P30)
      1: 1        F
      2: 00       Exponent of 10
      3: 62       Z Loc [ timer     ]

; This activates the table 2 measurement of thermal diffusivity.
     82:  Do (P86)
      1: 12       Set Flag 2 High

83:  End (P95)

; check to see if 63% timer data is available
84:  If Flag/Port (P91)
 1: 23       Do if Flag 3 is Low
 2: 30       Then Do

; set timer data to -9999
     85:  Z=F (P30)
      1: -9999    F
      2: 00       Exponent of 10
      3: 41       Z Loc [ td_time   ]

86:  Else (P94)

 ; clear flag 3 and leave timer data alone
     87:  Do (P86)
      1: 23       Set Flag 3 Low

88:  End (P95)

; end of Hukseflux code
;xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx


     89:  Do (P86)
      1: 10       Set Output Flag High

     90:  Sample (P70)
      1: 1        Reps
      2: 10       Loc [ LoggerID  ]

     91:  Totalize (P72)
      1: 1        Reps
      2: 11       Loc [ rain      ]

     92:  Resolution (P78)
      1: 1        High Resolution

     93:  Average (P71)
      1: 30       Reps
      2: 12       Loc [ Q7        ]

     94:  Z=F (P30)
      1: 0.0      F @@99
      2: 00       Exponent of 10
      3: 52       Z Loc [ BAUD      ]

     95:  If (X<=>F) (P89)
      1: 52       X Loc [ BAUD      ]
      2: 1        =
      3: 1        F
      4: 30       Then Do

;  PAM at 1200 baud
          96:  Serial Out (P96)
           1: 51       Printer Comma/1200 Baud

     97:  Else (P94)

;  ASTER at 9600 baud
          98:  Serial Out (P96)
           1: 52       Printer Comma/9600 Baud

     99:  End (P95)


;xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx

*Table 2 Program
  02: 1         Execution Interval (seconds)
; Every 1 second is needed to get sufficient accuracy for the thermal diffusivity
; measurement.

; Only executed when the heater is has just been switched off.
1:  If Flag/Port (P91)
 1: 12       Do if Flag 2 is High
 2: 30       Then Do

; timer one second incremented
     2:  Z=Z+1 (P32)
      1: 62       Z Loc [ timer     ]

; Measures the TP01 sensor output .
     3:  Volt (Diff) (P2)
      1: 1        Reps
      2: 3        25 mV Slow Range
      3: 4        DIFF Channel
      4: 63       Loc [ U_sen_ab  ]
      5: 1.0      Mult
      6: 0.0      Offset

     4:  Z=ABS(X) (P43)
      1: 63       X Loc [ U_sen_ab  ]
      2: 63       Z Loc [ U_sen_ab  ]

; Compare the sensor output to the 37% value
     5:  If (X<=>Y) (P88)
      1: 61       X Loc [ U_comp    ]
      2: 3        >=
      3: 63       Y Loc [ U_sen_ab  ]
      4: 30       Then Do

; The timer information is stored in td_time
; Now we know how much time it took to resettle.
          6:  Z=X (P31)
           1: 62       X Loc [ timer     ]
           2: 41       Z Loc [ td_time   ]

; ready again for new measurement instruction.
          7:  Do (P86)
           1: 22       Set Flag 2 Low

; set flag to trigger output signal in program 1
          8:  Do (P86)
           1: 13       Set Flag 3 High

     9:  End (P95)

     10:  If (X<=>F) (P89)
      1: 62       X Loc [ timer     ]
      2: 3        >=
      3: 0.0      F @@50
      4: 30       Then Do

; The thermal diffusivity is set to zero, because too much time has passed
; before return to the original value.
          11:  Z=F (P30)
           1: 0        F
           2: 00       Exponent of 10
           3: 41       Z Loc [ td_time   ]

; Ready again for new measurement instruction
          12:  Do (P86)
           1: 22       Set Flag 2 Low

; set flag to trigger output signal in program 1
          13:  Do (P86)
           1: 13       Set Flag 3 High

     14:  End (P95)

15:  End (P95)



;xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx

*Table 3 Subroutines

1:  Beginning of Subroutine (P85)
 1: 1        Subroutine 1

;  calibrate value of net rad based on sign
     2:  IF (X<=>F) (P89)
      1: 12       X Loc [ Q7        ]
      2: 4        <
      3: 0        F
      4: 30       Then Do

          3:  Z=X*F (P37)
           1: 12       X Loc [ Q7        ]
           2: 0        F @@1
           3: 12       Z Loc [ Q7        ]

     4:  Else (P94)

          5:  Z=X*F (P37)
           1: 12       X Loc [ Q7        ]
           2: 0        F @@0
           3: 12       Z Loc [ Q7        ]

     6:  End (P95)

7:  End (P95)

8:  Beginning of Subroutine (P85)
 1: 2        Subroutine 2

;  compute x = ln[(V/Vx) - 1] / 1000

     9:  Z=X+F (P34)
      1: 50       X Loc [ V_Vx      ]
      2: -1       F
      3: 50       Z Loc [ V_Vx      ]

     10:  Z=LN(X) (P40)
      1: 50       X Loc [ V_Vx      ]
      2: 50       Z Loc [ V_Vx      ]

     11:  Z=X*F (P37)
      1: 50       X Loc [ V_Vx      ]
      2: .001     F
      3: 50       Z Loc [ V_Vx      ]

12:  End (P95)

13:  Beginning of Subroutine (P85)
 1: 3        Subroutine 3

;  finish T (case/dome) computation

     14:  Z=1/X (P42)
      1: 50       X Loc [ V_Vx      ]
      2: 50       Z Loc [ V_Vx      ]

     15:  Z=X+F (P34)
      1: 50       X Loc [ V_Vx      ]
      2: -273.15  F
      3: 50       Z Loc [ V_Vx      ]

16:  End (P95)


17:  Beginning of Subroutine (P85)
 1: 8        Subroutine 8

     18:  Do (P86)
      1: 75       Pulse Port 5

     19:  Excitation with Delay (P22)
      1: 1        Ex Channel
      2: 0000     Delay W/Ex (units = 0.01 sec)
      3: 1        Delay After Ex (units = 0.01 sec)
      4: 0000     mV Excitation

20:  End (P95)

21:  Beginning of Subroutine (P85)
 1: 9        Subroutine 9

; pulse the AM416 multiplexer to the next bank of channels
     22:  Do (P86)
      1: 72       Pulse Port 2

     23:  Excitation with Delay (P22)
      1: 1        Ex Channel
      2: 0        Delay W/Ex (units = 0.01 sec)
      3: 1        Delay After Ex (units = 0.01 sec)
      4: 0        mV Excitation

24:  End (P95)

End Program

-Input Locations-
1 ACTIVE    1 0 0
2 TimeHack  1 0 0
3 Tref      1 0 0
4 _________ 1 0 0
5 pygi_raw  1 6 1
6 pygo_raw  1 5 0
7 pyri_raw  1 5 0
8 pyro_raw  1 5 0
9 _________ 1 0 0
10 LoggerID  1 1 1
11 rain      1 1 1
12 Q7        5 9 6
13 pygin     15 1 2
14 pygout    11 1 2
15 pyrin     11 1 2
16 pyrout    27 1 2
17 pygin_Tc  9 1 2
18 pygin_Td  9 1 1
19 pygout_Tc 9 1 1
20 pygout_Td 9 1 1
21 T107_1    9 2 1
22 T107_2    9 2 0
23 T107_3    9 2 0
24 T107_4    9 2 0
25 T107_5    9 2 0
26 T107_6    9 2 0
27 T107_7    15 1 1
28 T107_8    11 1 0
29 T107_9    27 1 0
30 ECHO_1    25 2 1
31 ECHO_2    9 1 0
32 ECHO_3    9 1 0
33 ECHO_4    9 1 0
34 ECHO_5    9 1 0
35 ECHO_6    31 1 0
36 HFT1      9 1 2
37 HFT2      25 1 1
38 HFT3      9 0 1
39 U_sen     1 2 1
40 U_heat    1 1 1
41 td_time   1 0 3
42 _________ 0 0 0
43 _________ 0 0 0
44 _________ 0 0 0
45 _________ 0 0 0
46 _________ 0 0 0
47 _________ 0 0 0
48 _________ 0 0 0
49 _________ 0 0 0
50 V_Vx      1 21 21
51 B0        1 4 4
52 BAUD      1 1 1
53 _________ 0 0 0
54 _________ 0 0 0
55 U_sen0    1 1 1
56 U_senmax  1 2 1
57 U_heatmax 1 2 1
58 U_heat_sq 1 0 1
59 U_senamp  1 1 1
60 U_sen90   1 1 1
61 U_comp    1 2 2
62 timer     1 2 2
63 U_sen_ab  1 2 2
-Program Security-
0000
0000
0000
-Mode 4-
0
1
2
3
4
5
6
7
8
10
11
12
13
14
15
16
17
18
19
20
21
22
23
24
25
26
27
28
29
50
99
-Final Storage Area 2-
0




15: Radiometers, Site sage, Fri 15-Nov-2002 13:43:53 MST, PAM sage CR10X coffs
PAM sage CR10X coefficients

Q7-94916	NET(A+)	00	9.21
x	A-	01	11.54
P26417	x	02	212.68
P27907	x	03	286.53
K940187	x	04	101.4198783
K970378 x       05      50.02501251
H983212	x	06	38.4
H993560	x	07	37.7
H943147	x	08	42.1
Tc-A0	x	10	2500
A1	x	11	3.355204
A2	x	12	.2795166
A3	x	13	3.950246
A4	x	14	209.3092
Tdin-A0	x	15	2500
A1	x	16	3.354969
A2	x	17	.2793838
A3	x	18	3.85506
A4	x	19	212.913
Tcot-A0	x	20	2500
A1	x	21	3.353713
A2	x	22	.2787871
A3	x	23	3.778048
A4	x	24	177.421
Tdot-A0	x	25	2500
A1	x	26	3.354732
A2	x	27	.2787514
A3	x	28	3.651471
A4	x	29	190.8095
HUX	x	50	180
baud	x	99	1








14: Radiometers, Site lake, Fri 15-Nov-2002 13:42:02 MST, PAM Lake CR10X coffs
PAM Lake Site CR10X coefficients

Q7-94197	NET(A+)	00	9.0
x	A-	01	11.22
P29137	x	02	240.211386
P29136	x	03	234.741784
K970377	x	04	43.55400697
K940186	x	05	82.71298594
H983213	x	06	38.2
H983211 x	07	36.4
H943149	x	08	39.7
Tcin-A0	x	10	2500
A1      x	11	3.354167
A2	x	12	.278817
A3	x	13	3.653292
A4	x	14	222.2646
Tdin-A0	x	15	2500
A1	x	16	3.350057
A2	x	17	.279229
A3	x	18	3.147866
A4	x	19	-193.854
Tcot-A0	x	20	2500
A1	x	21	3.354344
A2	x	22	.278532
A3	x	23	3.845153
A4	x	24	185.129
Tdot-A0	x	25	2500
A1	x	26	3.354233
A2	x	27	.278537
A3	x	28	3.802989
A4	x	29	184.2122
HUX	x	50	180
baud	x	99	1






13: Radiometers, Site tower, Fri 15-Nov-2002 13:01:12 MST, tower-grass CR10X coffs
Tower Grass CR10X coefficients

Q7-97069	NET(A+)	00	9.27
x	A-	01	11.48
P26416	x	02	290.866783
P29260	x	03	308.261406
K940181	x	04	82.3723229
K970379	x	05	45.51661356
H993563	x	06	38.9
H993562	x	07	37.3
H87034	x	08	36.8
Tc-A0	x	10	2500
A1	x	11	3.354472
A2	x	12	.2797812
A3	x	13	3.912974
A4	x	14	215.5208
Tdin-A0	x	15	2500
A1	x	16	3.052399
A2	x	17	.270981
A3	x	18	3.408667
A4	x	19	188.8687
Tcot-A0	x	20	2500
A1	x	21	3.353832
A2	x	22	.278595
A3	x	23	4.02801
A4	x	24	180.4266
Tdot-A0	x	25	2500
A1	x	26	3.053253
A2	x	27	.2706527
A3	x	28	3.530291
A4	x	29	166.282
HUX	x	50	180
baud	x	99	0









12: Camera, Site tower, Tue 12-Nov-2002 19:10:20 MST, Axis camera settings
Uploaded the latest firmware, version 2.32 to the 2420 axis camera.
(Had to re-try it several times.  Disable the operational program.
Also do a power cycle just before doing the upload, it seems to help).

These were the settings used during FLOSS01

MAC address: 00:40:8C:56:E0:E0
Image - General
  Appearance
    Images: single
    Resolution: 352x240 ~13Kb
  Tuning:
    Compression: Low
    Brightness: 10  (0..15)
    White Balance: Outdoor
    Color level: 6  (0..15)

  Heading:
    Text:  FLOSS
    Date: yes
    Time: 24h
  Default viewer:
    Java Applet

Image - Advanced
  Appearance:
    Sharpness: 166  0..255
  DC Iris:
    Exposure Control: DC-Iris
    Shutter Speed: 1/60
    Gain: 22dB   (10,16,22,28)
    ALC Level: 0

Image - Focus
  Assistant disabled

Network - TCP/IP
  Enable DHCP

  Internet addr: 192.168.10.40  (done via DHCP)
  Subnet mask: 255.255.255.0    (ditto)
  Default router: 192.168.10.1  (ditto

  HostName: webcam

  Domain Name: atd.ucar.edu
  Primary DNS: blank
  Secondary: blank


  Select Media: auto-negotiate
  Max Bandwidth: 1.0 MBit/s
  HTTP Port Number: 80


Network - SMTP
  Primary Mail server: blank
  Secondary Mail server: blank

Network Nofification:
  Http: disabled

  Smtp: disabled

  Ftp: disabled

COMM ports:
  RS232:
    Modem
    Baud rate: 9600
  RS485:
    none


Modem General:
  defaults

System - Date and Time

  New Camera time:
    Synchronize with NTP
      IP: 192.168.10.1
      TimeZone: GMT-07 Arizona, Mountain Time

System - Users
  root: ADV
  aster: ADV


Applications
  Operations - Scheduler
    Sequential Mode
    Primary Time Enabled: yes
      Restricted between
        Start: 05 00
        Stop:  20 00
        Mon-Sun
      Primary Image Freq:
        Every 15 mintues
        Regardless of input

  Sequential Operation - Upload
    Upload via: FTP, using: Network
    Remote Host: 192.168.10.1
    User name: aster
    password:

    Use Passive Mode: no
    FTP Port number: 21

    Image file:
      size: 352x240
      upload path: /data/aster/projects/FLOSSII/webcam
      base file name: flossii

      Overwrite: no
      Date/time suffix: yes
      Sequence # up to def max: no
      Sequence # up to spec max: no blank
  Application: enabled


Motion detection:
  not used

Layout - General:

  Enable personal settings: no

 





11: Kryptons, Site all stns, Mon 11-Nov-2002 09:36:00 MST, Krypton Cals
krypton		path		V0		Kw

1101		2.6		-0.1270443	666.8774800

1389		2.6		-0.1457689	753.3314535

1391		2.6		-0.1414481	1078.4729713

1392		2.6		-0.1370274	502.1903044

1394		2.6		-0.1395649 	1036.5063415

1395		2.6		-0.1418386	1262.7934824
		
1397		2.6		-0.134113	590.245198



10: Adams, Site tower, Fri 08-Nov-2002 21:22:39 MST, Cosmos
Cosmos
VME backplane #3
Matrix 9013
Ironics 422016
Datel 320689
Peltier cooler #2
VME Power supply: Tracer 298
Breezecom C

Boots ok, looks good

9: Adams, Site tower, Fri 08-Nov-2002 20:42:17 MST, Daisy
Daisy  
VME chassis #2
Matrix 9006
Ironics 821003
Datel 320589
Power supply questionable (rocker switch feels weird): Tracer 238
Teca: new 0033316
Breezecom B
Netgear ethernet hub getting power from backplane

Booted ok, but after about an hour, started getting errors:
interrupt: med0: receive buffer overflow, all buffers cleared.                 

Errors happened at bootup, before VxWorks logo is displayed, and
afterwards.  pings were erratic.

Network was through breezecom, on separate 192.168.10 subnet
DS11 statistics looked good.  Did not help to bypass breezecom
and go straight into hub.

Swapped in Matrix 9009.  No errors, looking good.

ingest_control -i daisy shows 310 samples/sec

Data_stats look good: 4 sonics @ 60hz, 5 TRHs, one CR10:

data_stats < dai021109.035234
Opened: 
chan #points     start time       end time        delta time    freq 
  0       23    03:52:34.470    03:56:15.030        220.560    0.100
100     4406    03:52:35.480    03:56:15.720        220.240   20.001
101     4406    03:52:35.480    03:56:15.720        220.240   20.001
102     4406    03:52:35.480    03:56:15.720        220.240   20.001
107      221    03:52:35.480    03:56:15.470        219.990    1.000
200    11327    03:53:06.970    03:56:15.740        188.770   59.999
201    11328    03:53:06.960    03:56:15.760        188.800   59.995
202    11329    03:53:06.940    03:56:15.750        188.810   59.997
203    11303    03:53:07.350    03:56:15.750        188.400   59.989
204      209    03:52:47.130    03:56:15.020        207.890    1.001
205      209    03:52:47.090    03:56:15.050        207.960    1.000
206      210    03:52:46.230    03:56:15.190        208.960    1.000
207      210    03:52:46.240    03:56:15.210        208.970    1.000
208      210    03:52:46.240    03:56:15.210        208.970    1.000
209       42    03:52:47.680    03:56:12.680        205.000    0.200











8: Logistics, Site base, Mon 04-Nov-2002 09:05:11 MST, PHONE #'s
Base Phones:
	Per Century Tel Active 11/02/02 through 04/30/03.
	Same phone numbers as last year:
970-723-3251  voice
970-723-3253  data
	Includes direct distance dialing.


Other:
Mark Lanier @ Reserve:
970-723-8202  ext 5

KK Cell	303-579-1130

7: Soil, Site all stns, Mon 14-Oct-2002 10:02:32 MDT, FLOSS02 soil sensors
On Thurs, 10 Oct''02 Tony and John visited the FLOSS site
The Echo soil sensors at Sage, Lake and Grass were reinstalled.
They had been installed horizontally the previous week and had been watered 
to help them meld into the soil. This melding was apparently sucessful at 
Sage and Lake, where the soil was suitable, but not so sucessful at Grass,
where the soil was actually peat.

The Echo probes were reinstalled vertically, with their midlines at the 
designated depths of 5 cm and 10 cm. They were positions at right angles
to each other with ~12cm between their cabled ends.


Sage  A 1 @  5cm
      A 2 @ 10cm
Sage  B 3 @  5cm
      B 4 @ 10cm
Sage  C 5 @  5cm
      C 6 @ 10cm
Each set of Echo probes waterd with 1 gallon


Lake  A 1 @  5cm
      A 2 @ 10cm
Lake  B 3 @  5cm
      B 4 @ 10cm
Lake  C 5 @  5cm
      C 6 @ 10cm
Each set of Echo probes watered with 1 gallon
 

Grass A 1 @  5cm (It is possible that A1 is actually at 10cm)
      A 2 @ 10cm (It is possible that A2 is actually at  5cm)
Grass B 3 @  5cm
      B 4 @ 10cm      
Grass C 5 @  5cm
      C 6 @ 10cm
Each set of Echo probes watered with 2 gallons
     
6: Logistics, Site lake, Fri 11-Oct-2002 15:12:18 MDT, Lake Site Visit : + Boom Angles
10/10/02

Site Visit:
Krypton Added, s/n 1393, using IHOP config.
Baro hose Added
Echo probes rotated.  Echo 'c' fixed
Rnet installed: Not working due to logger program problem.
Logger Cable installed

Photos Taken.
Angles Taken.

Angles NOTE: the assumed declination is 11.5-east.  The readings were
	adjusted up to account for the settings of the compass (8.5E,
	so added 3 / rounded) and the data-scope(6E so added 5.5 / rounded)

		     Data-Scope	      Compass
	CSAT		90		91	(Looking at sensor from west)
	RAD		90		90	(Looking at tripod from west)

PowerTower to PAM	230		229	(approx. mast-mast angle)
PowerTower to Rad	180		180	(approx. centers of structures)
Sonic to Rad		110			(Approx. from sensor to near
						 side of sawhorse)





5: Logistics, Site sage, Fri 11-Oct-2002 14:53:21 MDT, Sage Site Visit : + Boom Angles
10/10/02

Site Visit:
Krypton Added, s/n 1133, using IHOP config.
Serializer Added: sonic running at 20hz
Baro hose Added
Echo probes rotated.
GOES antenna (yaggi) angle adjusted slightly to correct setting
Rnet installed: Not working due to logger program problem.

Photos Taken.
Angles Taken.

Angles NOTE: the assumed declination is 11.5-east.  The readings were
	adjusted up to account for the settings of the compass (8.5E,
	so added 3 / rounded) and the data-scope(6E so added 5.5 / rounded)

		     Data-Scope	      Compass
	CSAT		95		92	(Looking at sensor from west)
	RAD		90		90	(Looking at tripod from west)

PowerTower to PAM	257		253	(approx. mast-mast angle)
PowerTower to Rad	204		199	(approx. centers of structures)
Sonic to Rad		150			(approx. from sensor to near
						 side of sawhorse)


4: Soil, Site lake, Fri 11-Oct-2002 14:51:44 MDT, Hft Probes 1,2 upside down
10/10/02

The sensors apparently didn't have the white dot and were installed
up-side down.  This is very apparent in the data.

Fix: can either be swap wires in logger, or change sign of logger coef.

3: Soil, Site all stns, Fri 11-Oct-2002 14:48:08 MDT, Soil Echo Probes Rotated to Vertical
10/10/02

Soil Echo Probes Rotated:
from Horizontal to Vertical orientation at Sage/Lake/Grass.
Water was dumped on to help bed the sensors in.  Tony said they already
were looking good in that respect.

The logger cable was added to the Lake site.  Both Sage/Lake are now reporting
data through goes.

2: Logistics, Site tower, Fri 11-Oct-2002 10:29:39 MDT, GUY WIRE TENSIONS
Guy wire tensions on the tall (114'/35m) tower are as follows

Note: Guy Wire Numbers are BOTTOM to TOP

Notes:
	- 10/03/02 Readings taken before adjustments and tower raised.
	- 10/10/02 Readings taken in a stiff S-SW wind @ >6m/s at sfc.
	  I could see the tensions changing (30-40) along with the velocity.
	- 11/20/02 Readings taken in calm air.
	- 1/6/03 Readings taken in calm air.
        - 2/18/03 dead calm, Tony Delany

				2002				 2003....
Date		10/3  10/4/02     10/10/02	11/20?	 12/16	1/6 1/27  2/18
Tensiometer	     (#1) (#2)	   (#2) 	(#1)(#2)
Array  Guy#
-----  ---	---	----	----	----	----	  ---- ---- ---- -----
NE	1 (bot)	390	620	630	540	640 650	  575   600  450  540
	2	390	580	590	550	600 600	  500   660  425  480
	3	390	590	590	570	600 610	  510	640  430  480
	4		500	510	450	450 460	  400	420  510  470
	5 (top)		560	550	500	480 49	  430 	430  550  510

SE	1	380	610	650	550	600 610	  575	620  520  500
	2	350	725	750	690	700 700	  660	610  580  560
	3	460	680	710	690	630 640	  610	590  520  490
	4		550	550	470	450 450	  450	410  600  510
	5		510	510	470	400 420	  415	360  500  500

SW	1	360	630	630	550	600 610	  575	600  540  510
	2	380	590	610	560	510 530	  505	470  470  450
	3	410	630	650	630	600 590	  575	490  500  440
	4		610	630	575	490 500	  625	420  580  510
	5		590	610	610	475 470	  450	360  600  490

NW	1	400	600	610	510	560 580	  490	550  440  440
	2	350	690	710	680	660 650	  600	590  500  530
	3	480	730	770	710	690 700	  600	560  570  580
	4		680	680	630	560 550	  475	490  610  640
	5		550	550	520	440 450	  390	340  520  490



















1: Soil, Site all stns, Tue 08-Oct-2002 08:48:09 MDT, Installation of soil sensors
On 2 Oct''02 the locations for the soil sensors for Grass, Sage and Lake
 soil sensor arrays were chosen by Tony, Steve O and Jielun.
Tony installed the sensors at Grass and Sge on 2 Oct and installed the 
sensors at Lake on the 3 Oct. After installation the sensors were wetted down
with approximately 6 liters of water to "bed them in". It is anticipated that 
given another month of warm weather the local soil moisture will come to 
equilibrium before the start of the program in December.
For the schematic of the placement see